### v11

# Minimal Algorithmic Information Loss Methods for Dimension Reduction, Feature Selection and Network Sparsification

Abstract
We present a novel, domain-agnostic, model-independent, unsupervised, and universally applicable approach for data summarization. Specifically, we focus on addressing the challenge of reducing certain dimensionality aspects, such as the number of edges in a network, while retaining essential features of interest. These features include preserving crucial network properties like degree distribution, clustering coefficient, edge betweenness, and degree and eigenvector centralities. Our approach outperforms state-of-the-art network reduction techniques by achieving an average improvement in feature preservation. Previous methods grounded in statistics or classical information theory have been limited in their ability to capture more intricate patterns and features, particularly nonlinear patterns stemming from deterministic computable processes. Moreover, these approaches heavily rely on a priori feature selection, demanding constant supervision. Our findings demonstrate the effectiveness of the algorithms proposed in this study in overcoming these limitations, all while maintaining a time-efficient computational profile. In many instances, our approach not only matches but also surpasses the performance of established network reduction algorithms. Furthermore, we extend the applicability of our method to lossy compression tasks involving images or any bi-dimensional data. This highlights the versatility and broad utility of our approach in various domains.
Keywords: data summarisation; lossy compression; data dimensionality reduction; network complexity; image segmentation; algorithmic machine learning
1 Introduction
The study of large and complex datasets, or Big Data, organised as networks, has emerged as one of the central challenges in most areas of science and technology, cellular and molecular networks in biology being one of the prime examples. The analysis of large-scale networks poses significant computational and analytical challenges. Hence, network sparsification methods have emerged as indispensable tools to address these challenges by reducing the complexity of networks while preserving essential structural or functional properties [14, 19, 21, 27, 31, 35, 46]. Moreover, it’s worth noting that data reduction techniques extend beyond network data. Methods for image sparsification also play pivotal roles in managing and analysing large datasets. In essence, data reduction, in a broader sense, involves transforming various types of numerical, textual, or visual digital information into more compact representations while retaining specific properties of significance.
This overarching concept of data reduction encompasses a wide range of methods, including network sparsification, image compression, and simplifying cellular automata rules. The common thread among these methods is the pursuit of more efficient data representation, enabling streamlined analysis and storage across diverse fields. The primary objective of data dimensionality reduction is twofold, much like two sides of the same coin. One facet emphasises minimising information loss, ensuring that valuable insights and nuances are preserved as data is transformed into a more manageable form. The other aspect underscores the importance of preserving the most ’meaningful’ features that uniquely characterise an object, a concept commonly referred to as feature selection. This dual objective serves as a fundamental principle in data reduction, often constituting a trade-off to be tackled in the data analysis phase. Traditionally, identifying these meaningful features relies on an user-centric, subjective criterion. Specific analytical goals and human expertise guide selection processes. For example, linear algebraic (e.g., matrix analysis) and statistically-based dimensionality reduction techniques attempt to minimise statistical information loss under certain algebraic (interpreted as signal and noise) conditions, consequently maximizing the statistical mutual information between the desired information and the dimensionally-reduced output.
The primary objective of dimensionality reduction in network analysis is to approximate a complex network with a sparser representation. Numerous methods for graph sparsification, a form of graph summarization, have been documented in the literature [46]. Some earlier approaches, such as Chew’s [27], relied on criteria like the shortest path distance between pairs of vertices for network sparsification. Others, like Benczur and Karger [20], utilized cut problems as a basis for sparsification. More recently, techniques such as spectral similarity of graph Laplacians have been employed by Spielman et al. [58]. When it comes to network dimensionality reduction, one must select criteria for preserving graph-theoretic properties like graph distance, clustering coefficient, or degree distribution, or combinations thereof. However, it’s important to acknowledge that no finitely computable approach can comprehensively identify all possible features of interest in a dataset. This includes recursively enumerable features that the set of all Turing machines can characterise concurrently [65]. Consequently, data analysts often face the daunting task of making arbitrary choices regarding which features to prioritize, as exemplified in previous research [66, 67].
This paper’s main goal is to develop an unsupervised method that remains theoretically optimal and is independent of computational models, programming languages, physical implementations, and available resources. Furthermore, this method should be agnostic to selecting features of interest, encompassing all computably enumerable parts within a given dataset earmarked for reduction. We introduce a family of semi-computable algorithms designed to preserve computable properties, encompassing both statistical and algorithmic aspects. This approach is a generalization of various dimension reduction techniques since statistically definable, generable, or recognizable patterns constitute a subset of those that algorithmic approaches can theoretically handle.
Our method represents a novel approach to designing theoretically optimal lossy compression techniques, drawing inspiration from principles and estimations related to theoretical optimal lossless compression [63, 18]. By employing efficient (polynomial) approximations to algorithmic complexity values from recent numerical methods of algorithmic probability [32, 56, 68, 55, 62, 69], we demonstrate how these algorithms can preserve structure properties, performing similarly to, if not consistently outperforming, state-of-the-art algorithms in e.g. the area of network dimension reduction. We test our algorithms on non-trivial cases against transitive and non-linear (spectral) methods involving simple graphs where statistical regularities are even easier to conceal more intricate computable processes or embedded encodings, and thus may easily fool weaker, linear and computable measures [66, 65, 11, 12].
We will demonstrate how our method preserves critical network topological properties, including degree distribution, clustering coefficient, edge betweenness, degree, and eigenvector centralities compared to other methods. These fundamental properties provide essential insights into network structure and functionality. Degree distribution, for instance, reveals how nodes are interconnected, helping us understand the overall network architecture and the distribution of connections. It is indispensable for identifying hubs and assessing network robustness. The clustering coefficient quantifies the tendency of nodes to form tightly connected clusters. This metric is invaluable for detecting network communities and assessing resilience against failures or attacks. Edge betweenness measures the significance of edges in facilitating efficient communication between nodes. Identifying bottleneck edges is vital for understanding network flow and vulnerability. Degree and eigenvector centralities help pinpoint influential nodes within a network. High-degree nodes are hubs, while high-eigenvector centrality nodes are well-connected to other influential nodes. These centralities are vital for identifying key players in various network types.
By preserving these critical properties, our approach strives to provide more meaningful and accurate representations of complex networks [67, 11, 12] while addressing the challenges inherent in data reduction techniques. Central to our approach are perturbations in the form of edge deletions aimed at minimizing the loss of algorithmic information content [13]. We have aptly named our algorithm MILS, signifying ’Minimal Information Loss Selection,’ as it encapsulates our mission of preserving essential information while reducing data dimensionality. Notably, the algorithms presented here remain agnostic to the specific method for approximating algorithmic complexity. They can be implemented using various approaches, including entropy or lossless compression. However, our description employs a technique based on [68, 72, 63], which not only encompasses but also enhances other statistics- and entropic-based methods due to its intrinsic structural characteristics.
In Section 2.2, where we detail the algorithm, it’s essential to recognise that substituting the underlying methods for approximating algorithmic information content with alternatives like Shannon entropy or lossless compression algorithms represents special cases within the broader algorithm based on algorithmic complexity. This comprehensive approach effectively covers all these less powerful cases because the block decomposition method (BDM) calculation we employ subsumes entropy-based approximations [68]. BDM presents a novel approach to understanding and quantifying the complexity of biological networks and systems. BDM extends the power of the Coding Theorem Method (CTM) [62, 55], offering a closer connection to algorithmic complexity than previous methods based on statistical regularities. This method effectively decomposes complex objects into smaller components using a set of short computer programs. These components are then sequenced to reconstruct the original object, allowing for efficient estimations of algorithmic complexity. BDM as has been shown by our team and other researcher addresses a fundamental challenge in science: capturing the complexity of objects for classification and profiling. It bridges the gap between computable measures like Shannon entropy [55], which lacks invariance to object descriptions and uncomputable universal estimates of complexity. The method lies between these two, dividing data into smaller pieces to partially circumvent the uncomputability problem. This results in efficient estimations, albeit with some loss of accuracy.
BDM has been applied in various fields and applications outperforming other complexity measures [28, 67], for example to evolutionary systems [38], artificial intelligence [41, 17], and the detection of bio- and technosignature [26, 29]. Significantly, BDM’s application to graphs and causal calculus has enabled the deconvolution of complex network behaviors, identifying distinct topological generating mechanisms [33, 63, 18, 61]. This approach uses a perturbation-based causal calculus to infer model representations [13, 60, 40], effectively tackling the challenges of causation in complex systems. It has been demonstrated by us and others how by analyzing data from discrete dynamical systems like cellular automata and complex networks, BDM-based algorithms provides a deeper understanding of the intricate interplay of elements within these systems, revealing hidden patterns and relationships, and contributing significantly to the field of algorithmic complexity in biological networks[33, 18, 64, 70, 72, 63]. Moreover, algorithmic complexity, being an asymptotically optimal measure of irreducible information content [26, 34, 43, 24], holds the unique advantage of capturing algorithmically determined patterns that defy reduction to statistical patterns [28, 66].
Our implementation of the MILS algorithm, rooted in algorithmic complexity, offers the added benefit of avoiding certain low-complexity deceptive phenomena. Traditional methods based solely on entropy and statistics are prone to such issues [28, 66]. Significantly, this avoidance does not compromise running time efficiency or data size reduction performance, as evidenced by our results in Section 4. These results demonstrate that we either match or, more often than not, surpass the performance of the best current algorithms, both for local and global graph properties.
The approach presented in this article lies in its potential to assess the effectiveness of all reduction techniques while striving for optimal reduction by minimising algorithmic information loss. This represents a nonlinear generalisation encompassing a broader range of techniques beyond merely preserving statistical or domain-specific algebraic properties. While our focus in this article centers on edge deletions in graphs to describe the MILS algorithm, it’s essential to emphasise that MILS can be applied to a wide array of objects and various types of constructive or destructive perturbations [13, 60, 72, 70] on their constituent elements. For instance, MILS can seamlessly extend its applicability to node deletion, offering versatility in addressing diverse algorithmic perturbations [13] aimed at minimising information differences between objects.
2 Method
2.1 Calculating information difference values
All methods employed in this article are founded on the information difference among an object’s elements. Specifically, we focus on the contribution of each element to the system’s overall information content, such as the impact of nodes or links in a network context. Formally, this involves assessing the impact of local perturbations (changes made to individual elements) on the algorithmic information content of the entire object, compared to its prior state. These methods are grounded in algorithmic/causal perturbation analysis, a key component of algorithmic information dynamics (AID) [13, 60, 40]. In the specific context of this article, our approach entails the perturbation of system elements through the removal of components, followed by measuring and ranking the effects on the system’s algorithmic information content. This dynamic approach to studying and calculating the potential changes in an object for evaluating the contribution of its components was initially introduced in [61, 64, 72, 71].
We extend these concepts to data/network dimension reduction and lossy compression of bi-dimensional data, such as images. For a graph $G$ with edges $e_{1},\ldots,e_{|E(G)|}\in E(G)$, we consider the subset $F\subseteq E(G)$ representing single-edge perturbations. The graph $G\backslash F$ results from removing edges in $F$. The algorithmic complexity $C(G)$ (as detailed in the Supplementary Material) allows us to define the information difference for edge deletion:
$I(G,F)=C(G)-C(G\setminus F)$ (1)
which quantifies the information difference of each edge $e_{i}$ in the context of edge deletion. Notice that this difference could be defined for perturbations on the graph’s vertices, but here, we will restrict our attention to perturbations on edges. In the present article, as introduced in the next Section 2.2, we focus on the empirical investigation of destructive perturbations (i.e., the edge deletion problem) to perform network reduction while minimising information loss. However, in the Sup. Mat. we also theoretically investigate the edge-insertion case as an analogous inverse case of edge deletion. Future research will be fruitful for empirically correlating the present results in Section 4 (which hold for destructive perturbations) with those for constructive perturbations.
2.2 An algorithm for minimal information loss selection
MILS is an unsupervised and mostly parameter-free algorithm, i.e., asymptotically independent of model or domain, as it does not need to be instructed or designed to preserve any particular property, and maximises the preservation of all computable elements that contribute to the algorithmic information content of the data.
Algorithm 1 Minimal Information Loss Selection (MILS)
1:function InfoRank($G$)
2:     $\textit{informationLoss}\leftarrow\varnothing$
3:      // for each edge $e_{i}$
4:     for $e_{i}\in E(G)$ do
5:          // store information contribution into informationLoss
6:         $\textit{informationLoss}\leftarrow\textit{informationLoss}\cup\{I(G,e_{i})\}$
7:     end for
8:     sort informationLoss in increasing order
9:      // return information rank
10:     return informationLoss
11:end function
12:function MILS($G,N$), $1\leq N\leq|E(G)|$
13:     while $|E(G)|>N$ do
14:          // calculate minimal loss across all edges
[…]
16:          // remove all candidate edges from $G$
17:         $\textit{G}\leftarrow G\backslash\{e_{i}\in E(G):I(G,e_{i})=\textit{minLoss}\}$
18:     end while
19:     return $G$
20:end function
Let $G$ be a graph, and $\left|E\left(G\right)\right|$ denote its number of edges. MILS seeks to reduce $G$ to a graph on $N<\left|E\left(G\right)\right|$ edges so that the loss of information is minimized. Ideally, one would do so by calculating the information differences between $G$ and $G\backslash S$, where $\left|S\right|\geq 1$, and then finding the subset $F\subset E(G)$ such that $I(G,F)\leq I(G,S)$ for all non-empty proper subsets of edges $S\subset E(G)$, repeating this task with $G\coloneqq G\backslash F$ until the target size is reached (i.e., when $|E(G)|=N$). This algorithmic procedure’s time complexity is clearly in $O(exp)$ because of how it performs searches and deletions across all subsets of edges (see also Fig. 2), but significant improvements to this bound are still possible as shown by Algorithm 1 (see also Algorithms $2$ and $3$ in the Sup. Mat.).
A more efficient but suboptimal version of such an exponential-time MILS iterates over single elements (in the case, edges) or singletons. Algorithm $2$ in the Sup. Mat. performs such a sequential deletions of edges, and then removes the edge that contributes less to the information content of the graph before moving to the next step. When $e$ is such that $\left|I(G,e)\right|$ is minimal, where $I(G,e)$ denotes $I(G,\{e\})$ (i.e., the information difference from deleting only a single edge), we call it a neutral information edge [72] because it is an edge that contributes the least to the information content of $G$ (in particular, it minimizes information loss or the introduction of spurious information into the network according to the information difference when removed from the original network). The existence of such edges whose information is maximally neutral (or redundant) is demonstrated in Theorem $3.8$ in the Sup. Mat., where it is shown that there are infinite families of graphs for which each member has at least one maximally neutral information edge $e_{i}$ (in the edge deletion case), where $\left|I(G,e_{i})\right|=\left|C(G)-C(G\backslash e_{i})\right|=\mathbf{O}(1)$. This is the opposite case of a maximal information edge, i.e., a single edge whose deletion changes the most the information content of the entire $G$, case which refers to incompressible graphs (see Section 3 and Fig. 1 discussed below). Thus, although suboptimal in general, our application of more efficient versions of MILS based on single edges is capable of grasping the full range between minimal and maximal information contribution to the whole network. See Section 3 for the information difference error of our efficient algorithms in comparison to the optimal exponential-time MILS.
The pseudocode in Algorithm $2$ in the Sup. Mat. assumes that there is a unique such $e$, which may not necessarily be the case in general. Algorithm 1 solves this problem by performing simultaneous deletions on all edges with an information contribution given by those in minLoss. To this end, Algorithm 1 introduces InfoRank, a method that produces a ranking of $e_{1},\dots,e_{|E(G)|}$ from least informative to most informative edge, i.e., a list of edges sorted in increasing order by their information contribution to $G$. This ranking facilitates the search for the (most) neutral elements of the system (see Sections $2.8$ and $3.3$ in the Sup. Mat.), which in turn helps MILS preserve the components that maximise the information content of the resulting object with a speed-up in running time compared to the ideal exponential-time version of MILS—a speed-up achieve even in case neutral elements are not unique. See also Section 3 for a theoretical discussion.
Fig. 1 presents three illustrative examples of Algorithm 1 returning the same graph (with $N\leq 2$ present edges) as output for three distinct inputs displayed in Fig. 1A, Fig. 1B, and Fig. 1C, respectively.
Figure 1: Three examples of edge deletions occurring in each iteration $j$ of [Algorithm 1, Step 13], indicated by the downward blue arrows (except for the bottommost blue arrows), while the subsequent red squares indicate where the deletions occurred in the previous step. The bottommost blue arrows in each Fig. 1A, 1B, and 1C refer to [Algorithm 1, Step 19] after which the final output of the algorithm is returned, given an adjacency matrix of a graph $G$ (such as anyone of the three topmost $4\times 4$ matrices) as input. The input adjacency matrix in Fig. 1A corresponds exactly to a substring of the halting probability as in [23]. Mixing both redundant and incompressible components, the second row of the input matrix in Fig. 1B is a repetition of the first row, and the remaining $8$ bits equals to those of the input matrix in Fig. 1A. Fig. 1C presents one of the most redundant cases in which all oriented edges are present, that is, the input is an adjacency matrix of a complete (directed) graph.
The topmost matrices in columns A, B, and C display the adjacency matrices of distinct directed graphs (without self-loops), all of them with four vertices, that are given as inputs. Each of these graphs are subjected to the edge deletions that Algorithm 1 performs so as to generate its final output. As the adjacency matrix given as input becomes less random, one should expect that the number of iterations can be reduced. Indeed, this is what happens in Fig. 1A,  1B and 1C, respectively: in example A, the input matrix is (algorithmically) random, i.e., incompressible, hence no edge in the (ordered) list given by InfoRank will be as neutral as the other ones, and minLoss will contain only one edge at the time; in example B, the edge in the first row is a neutral information edge as much as the edge in the second row, and therefore both edges will be deleted in only one iteration step; in example C, except for the last two edges in the last row, every edge will be considered neutral in InfoRank. Further discussion on Fig. 1 can be found in Section $2.8$ in the Sup. Mat..
The tree diagram in Fig. 2 illustrates the decision process involved in each iteration of Algorithm 1 that results in the next size-reduced graph, speeding up this process both locally and globally across the tree. Each child vertex represents a graph resulting from the deletion of one or more edges in the parent vertex (i.e., a graph before the destructive perturbations on the edges). In each descending step starting from the root at the top, as the height of the tree downward increases, a subset of (present) edges is deleted from the parent graph at the lower level. Notice that from each parent vertex, there can be as many child vertices as the size of the power set of the edges (excluding the empty set) in the graph represented by the parent vertex. The child vertex/graph chosen by Algorithm 1 is the one whose neutral information edges (according to InfoRank) were deleted from the parent vertex/graph. In this manner, the iteration process in [Algorithm 1, Step 13] is represented by the rooted path highlighted in the color blue in the Fig. 2. A subset of edges that has neutral information with respect to the parent graph is deleted from the parent graph so as to generate the child graph in this path. Thus, InfoRank allows Algorithm 1 to select a child graph that minimises the information difference with respect to the parent graph, not having to search over all the possible subsets of present edges. This single-thread process continues until the leaf vertex, i.e., the graph with (at most) $N$ edges is achieved.
MILS Tree Diagram
Figure 2: Tree Diagram for Algorithm 1 with the sequence of graphs with their respective subsets of edges reduced (as selected in the iterations $j$ of [Algorithm 1, Step 13]). So, one has that $F_{i_{1}}=\textit{minLoss}_{1}=\min\left(\textsc{InfoRank}\left(G\right)\right)$, $F_{i_{2}}=\textit{minLoss}_{2}=\min\left(\textsc{InfoRank}\left(G\setminus F_{i_{1}}\right)\right)$, and so on, which continues up until the final graph $G\setminus F_{i_{h}}$ is found by the algorithm. This process is represented by the blue path $\left(G,G\setminus F_{i_{1}},G\setminus F_{i_{2}},\dots,G\setminus F_{i_{h}}\right)$ from the root $G$ to the leaf $G\setminus F_{i_{h}}$ in $h$ iterations. Notice that from each parent vertex/graph, there are as many children as (non-empty) subsets of (present) edges in the parent. Fig. 2 illustrates a speed-up both globally (as the number of iterations $j\leq h$ is always upper bounded by $\mathbf{O}\left(\left|E\left(G\right)\right|-N\right)$) and locally (as from each parent, Algorithm 1 does not have to search for the best candidate across all possible subsets of edges of the parent) achieved by Algorithm 1 from the decision process that this tree diagram represents.
This allows a local speed-up between each parent graph and its children in Fig. 2, in particular by replacing a search over an exponential-size set of subsets with a search over a list (of edges), therefore quadratic with the number of vertices/nodes in the root graph. It also allows a global speed-up between the root vertex and the final output of Algorithm 1, particularly by replacing a search over all possible rooted paths with a recursive iteration in [Algorithm 1, Step 13] whose maximum number of iterations is always upper bounded by $h\leq\left|E\left(G\right)\right|-N+1$, therefore quadratic with the number of vertices in the root graph. This makes the height of the tree always much shorter in comparison to its width.
It is also important to notice that the decision tree diagram (i.e., the one for the minimisation of the information difference between the parent and the child vertex) in Fig. 2 can be subsumed into a probability tree diagram. As shown in Section $2.8.1$ in the Sup. Mat., each path from the root to a leaf has an associate joint probability of a sequence of independent events, and each event corresponds to a destructive perturbation on a subset of edges. The rooted path highlighted in the color blue in the Fig. 2 illustrates the path to the final leaf vertex (i.e., the vertex representing the outcome of Algorithm 1) that maximises the (algorithmic) probability of edge deletions, while each edge in this path corresponds to an iteration in [Algorithm 1, Step 13]. As demonstrated in Section $2.8.1$ in the Sup. Mat., which in turn is a result that evinces the generality of our methods, the MILS tree diagram in Fig. 2 shows that a decision process of minimisation of information differences implies a maximization of (algorithmic) probabilities both locally for each arrow and globally for each path from the root to a leaf. This holds theoretically independent of the choice of the computation model, programming language, or any arbitrary (computable) probability (semi-)measure that one aims at maximising (whether or not one assumes independence between parent vertices and subsequent child events in such a measure). See Section $2.8.1$ in the Sup. Mat. for a thorough explanation of the tree diagram and its mathematical properties.
As described in [72] and discussed in the Sup. Mat., one can easily modify Algorithm 1 to deal with multiple edge deletions at once, as in $I(G,F)$, but possibly at the expense of much more computational resources. In addition, note that these algorithms may be applied, mutatis mutandis, to nodes or to any data element of an object or dataset, e.g. a pixel of an image, or a row or column in a spreadsheet, which constitutes an exciting application of our current methods as shown in Section 4.2.
3 Theory
In this section, we introduce the theoretical results that support the methods employed in Section 4.
As presented in Section 2.1 and formalized in [13, 60, 72, 64], we have that edge deletions can be seen as transformations/perturbations that a particular network is being subjected to. This way, it is immediate to ask about the maximum overall impact of these transformations on the algorithmic information content of the network. Indeed, in Section $3.1$ in the Sup. Mat. we show that the theoretically optimal information difference (i.e., the information difference calculated by the exact theoretical algorithmic complexity values) produced by edge deletions are bounded by terms that depend on the number of single-edge perturbations, the network size and/or graph sparsity (or edge density). In the general case, Theorem $3.1$ in the Sup. Mat. demonstrates that algorithms for approximating $I(G,F)$ ideally calculate values that fall into the range of values upper- and lower-bounded by terms in a logarithmic order of the network size $N=\left|V(G)\right|$ and in a linear order of the number $\left|F\right|$ of single-edge perturbations, except for other additive terms that are strongly dominated by the former two terms, respectively. As shown in Corollary $3.2$ in the Sup. Mat., this behavior remains the same even in case the graph is (algorithmically) random, i.e. incompressible—as it is the case of most sufficiently large Erdős-Rényi (ER) random graphs with edge probability $1/2$—, so that even in the case each edge theoretically carries the maximum information it can, or carries the least amount of redundancies (that is, each edge contributes as much as the others to the overall irreducible information content of the graph), those upper and lower bounds hold.
As also shown in Sup. Mat., by replacing network size with edge density (i.e., the size of $\left|E(G)\right|$ in comparison to $\frac{N^{2}-N}{2}$), we demonstrate even tighter lower and upper bounds for multiple edge deletions that are in a logarithmic order of the edge density and, improving on the general case to the above, in direct proportion to the number $\left|F\right|$ of single-edge perturbations. Thus, our results demonstrate that these limits for the algorithmic information loss (or gain) from edge perturbations have a linearly dominant dependence on the number of single-edge perturbations, which is denoted by $\left|F\right|$, except for a multiplicative logarithmic term depending either on the network size or on the edge density. In turn, these strongly dominate all the other reminiscent additive terms.
In addition, as expected in an alternative approach for estimating those lower and upper bounds but based on a straightforward and direct encoding of the subsets of edges to be perturbed—therefore, more simplistic than the proof techniques employed in this work—, we demonstrate that one can always retrieve other lower and upper bounds which (although not as tight as the ones we obtained) do only depended on the network size, but now in a quadratic order. See Corollary $3.3$ in the Sup. Mat..
Moreover, as should be expected for only one single-edge perturbation (i.e., when $\left|F\right|=1$), we demonstrate that one can also derive that these upper and lower bounds are still tight in general (see Corollary $3.4$ in the Sup. Mat.), but in this case only depending on a logarithmic order of the network size.
These theoretical results pave the way for understanding why MILS algorithms are optimal with respect to their chosen edge-deletion scheme, respectively, once one fixes the arbitrarily chosen method capable of approximating the algorithmic complexity values in the asymptotic limit. That is, the algorithms return smaller networks whose respective edge count reductions are done so that the information difference between the original network and the smaller one is minimised according to those optimal upper and lower bounds described in the above paragraphs. This holds not only for the ideal EXPTIME version of MILS mentioned in Section 2.2, whose iterations cover the power set of the set $E(G)$ of present edges, but our results show that switching to a summarization procedure based on successive single-edge deletions (like in Algorithm 1) is sound. In particular, as discussed in this section, the latter case is grounded on the fact we proved that any maximum error does not dependent on the particular ideal subset of edges to be deleted, but only on the sheer size of this subset (once the worst-case logarithmic contribution of the network size becomes fixed). This means that in the theoretical limit with unbounded resources, Algorithm 1 can never misestimate the minimisation of information difference by an error larger than a linear factor of the number of single-edge perturbations that should otherwise be ideally deleted. Even if one compares the Algorithm 1’s results with the ones obtained from the best deletion scheme (i.e., the one in EXPTIME), the maximum information difference error (when choosing a scheme distinct from the ideal exponential-time MILS) increases linearly with a logarithmic term of the network size as the number of single-edge perturbations increases. Thus, this error is far from the quadratic dependence on the network size (see Corollary $3.3$ in the Sup. Mat.) that one might expect from a more simplistic method of encoding subsets of edges, as discussed to the above.
Theorem $3.11$ in the Sup. Mat. demonstrates that Algorithm 1 not only returns a unique output for each input (which holds even in the case of more than one neutral element at each iteration, as discussed in Section 2.2) but also has: a worst-case time complexity with a cubic dependence on input $E(G)$, at the same time that it scales linearly with the worst-case time complexity of the chosen method for approximating the algorithmic complexity values of any edge.
Thus, our methods employed in Section 4 take advantage of a trade-off between the information difference error and the running time speed-up in which: the information difference can only diverge from the ideal one in the EXPTIME MILS algorithm by a linear order of the number of single-edge perturbations, divergence which does not depend on the ideal subset of edges that could have been otherwise deleted in the EXPTIME version of MILS; while the speed-up improves from an exponential to a cubic worst-case time complexity. This speed-up is illustrated by the tree diagram in Fig. 2. Instead of searching along all possible paths, which would demand exponential running time, each iteration corresponds to an one-step increase in the height of the tree from the root, and in each step the algorithm runs the InfoRank procedure in order to select the edge deletions to be made.
4 Results
4.1 A comparison between MILS and other data dimensionality reduction methods
Figs. 34567, and 8 illustrate the effectiveness of the Minimal Information Loss Selection (MILS) method in retaining crucial local and global characteristics of synthetic and natural networks, encompassing varied types and topologies. These figures demonstrate MILS’s performance, typically surpassing leading graph sparsification algorithms. The study involved well-known networks from [49], such as genetic regulatory, protein, power grid, and social networks. MILS was applied to these networks for comparative analysis with transitive reduction [14] and spectral sparsification [58], two powerful sparsification methods. The former seeks to minimise edges while maintaining the same reachability within a directed graph, while the latter, thoroughly introduced in [19], aims to reduce network dimensions based on spectral similarity of graph Laplacians, thus preserving significant properties through the adjacency matrix Laplacian spectrum. Specifically, Figs. 3 and 4 showcase the performance of MILS in preserving the degree distribution and edge betweenness distribution in a typical synthetically (recursively) generated Erdős-Rényi (ER) random graph with low edge density, comparing its results against both random edge deletion and spectral sparsification methods.
While MILS is not significantly better at preserving the clustering coefficient of random networks, Fig. 5 shows that MILS does significantly better at preserving the clustering coefficient of real-world (biological, social, and electric grid) networks taken from [49], outperforming both transitive and spectral reduction/sparsification methods.
Figs. 6, 7 and 8 illustrate how MILS outperforms spectral sparsification at preserving edge betweenness, and degree and eigenvector centralities, respectively.
4.2 An application of MILS to image lossy compression
Fig. 9(A) and Fig. 9(B) illustrate the application of MILS in the reduction and coarse-graining of two simple cases of the space-time evolution of Elementary Cellular Automata (ECA) rules 22 and 158. The reduction is by minimisation of algorithmic information loss. MILS effectively extracts the salient elements that characterise each of these systems. The examples in Fig. 9(A) and Fig. 9(B) depicts the way in which regions with high or low algorithmic content can be ranked, selected, or preserved for dimensional reduction purposes. The method performs an unsupervised lossy compression able to preserve both cellular automata’s main features (not covered in grey), with no intervention and no parameter choice. MILS proceeds by deleting the regions with the lowest algorithmic information content contribution and maximizing the preservation of the features that contribute the most to the algorithmic description of the objects. The extracted features will not be as clear as in these examples as they may pick more complicated patterns, even not statistical ones based on algorithmic probability. Unlike statistical approaches, the algorithm can approximate (and thus preserve/extract) features that are of an algorithmic nature and which are not statistically apparent as it was in this case (see [68, 66]) and next examples.
Fig. 10 shows the MILS algorithm applied to images. Fig. 10(A) shows how vertical and horizontal compression preserves features even if distorting the image hence showing different mechanisms and goals than those from popular image compression algorithms. The purpose of image compression algorithms such as JPEG is to maximise storage compressibility and require a decoder, here the application is directly to the image itself both at storage and visualisation stages. Is also important to mention that our purpose is not to maximise storage compression but to minimise the loss of algorithmic content in the reduction process. Fig. 10(C) shows how the algorithm preserves the main features of the image leaving almost intact the formulae.
MILS is highly context-sensitive and adapts to different kind of information such as images. For example, it handles in different ways images as seen in Fig. 10. This is because information in Fig. 10(A) is distributed more uniformly than Fig. 10(C) but, according to the algorithm, some features in certain regions are more important than others and thus some distortions are allowed if such features have high algorithmic content according to the underlying computer programs that reconstruct the compressed images. However, in Fig. 10(C), blank space between formulae can be sacrificed first as they do not contain any information. While this behaviour for handling different cases can be replicated in ad-hoc algorithms, MILS handles it naturally and determines when distorting can be allowed or proportion can be preserved, something that image compression algorithms would not do as their goal is to recover the same image dimensions and maximise storage compression and not to preserve algorithmic information content even sacrificing otherwise fundamental properties of the image for those algorithms such as image dimension.
5 Conclusion
In this paper, we have introduced Minimal Information Loss Selection (MILS), a novel agnostic approach for data summarisation that is particularly effective in reducing data dimensionality while preserving essential features such as degree distribution, clustering coefficient, edge betweenness, and centralities. As studied in Section 4.1, MILS demonstrates comparable or superior performance to existing network reduction algorithms on different classes and types of networks. Moreover, it extends to lossy compression of images and bi-dimensional data as shown in Section 4.2. It operates on the principle of local perturbation analysis and algorithmic information dynamics, offering robustness against varying methods of approximating algorithmic complexity. The method presented in this article optimally preserves the algorithmic information content (which might be affected by the local perturbations) whenever an optimal method is used to approximate algorithmic complexity. This guarantees that its performance can only increase as more computational resources are made available, which differs from other methods based on statistics for which theoretical limitations are known to hold. Nevertheless, the empirical application of MILS only scales linearly with the worst-case time complexity of the chosen method for approximating algorithmic complexity.
Funding
FSA acknowledges support from the São Paulo Research Foundation (FAPESP), grants $2023$/$05593$-$1$ and $2021$/$14501$-$8$. NAK is supported by Horizon $2020$ EU grant (MultipleMS, $733161$).
Appendix A Figures
Figure 3: MILS or neutral edge deletion (blue) outperforms random edge deletion (red) at preserving both edge degree distribution (top, showing removed edges) and edge betweenness distribution (bottom) on an Erdős-Rényi random graph of vertex size 100 and low edge density ($\sim 4\%$) after up to 60 edges were removed (degree distribution comparison) and 150 edges were removed (edge betweenness) out of a total of 200 edges (notice also the scale differences on the $x$-axis).
Figure 4: Histograms showing preservation of degree distribution from 20% to 80% edge removal. Green highlights the overlapping and the preserved area of the distributions after random deletion (top), MILS and spectral removal (bottom pairs).
Figure 5: MILS mean clustering coefficient preservation against two other sophisticated graph sparsification methods based on graph spectral and transitive reduction techniques on biological, electric and social networks taken from [49]. The transitive method does not allow selection of edges to be deleted, and in some cases it either fails to significantly reduce the network edge density if no cycles are present (such as, generally, in electric and genetic networks) and/or takes the clustering coefficient to 0 (e.g. for protein networks) if cycles are only local. Comparisons with other methods are unnecessary because they destroy local or global properties by design, such as clustering coefficients for the spanning tree algorithm.
Figure 6: Stacked histograms showing edge betweenness preservation of MILS versus spectral sparsification across different families of networks. The similarity in height of each segment is an indication of the preservation of such properties. Blue bars (MILS) approximate yellow (original) bars better than spectral sparsification. On average MILS was 1.5 times the edge betweenness distribution of these representative graphs measured by the area similarity of the respective bars.
Figure 7: Stacked histograms showing the preservation of degree centrality after application of MILS versus spectral sparsification across different families of networks: bars with height closest to the original graph signify better preservation. Blue bars (MILS) approximate yellow (original) bars compared with spectral sparsification. MILS only slightly outperformed spectral sparsification in this test but never did worse.
Figure 8: Stacked histograms showing eigenvector centrality preservation of MILS versus spectral sparsification across the different families of networks: bars with height closest to the graph’s original bar signify better preservation. Blue bars (MILS) approximate yellow (original) bars better than spectral sparsification both in distribution shape and individual bar height. On average MILS preserved the eigenvector centrality distribution of these representative networks 1.5 times better.
Figure 9: In (A) and (B) we show a feature selection and image reduction by application of MILS, starting from the original and second step, highlighting the regions that are earmarked to be omitted (in grey) versus the features that are kept along the way, thereby optimally preserving the main properties of these objects, properties whose persistence enables a ranking of such features. Here can be seen how boundaries are favoured as key features. In (C) we see the image reconstruction by preservation of extracted features from the image with the highest algorithmic probability (indicated by red signatures in the first two image processes). The MILS algorithm selects features which are considered most important in the compressed image reconstruction.
Figure 10: Minimal Information Loss Reduction, a MILS-based lossy compression algorithm. A: Row compression preserves features on even highly coarse-grained versions (B starts from A at 0% but sampled to make it a 100 $\times$ 100 image from the original 600 $\times$ 600) of the same image producing a cartoonish representation. The image in (B) reduced by 23% is 10% the size of the image at 0% in (A). C: Compressed formula by MILS vs random row deletion (D) with resulting images of the same size but MILS preserving text proportion and minimizing information loss while random row deletion distorts the text.
Appendix B Preliminary definitions and background results
B.1 Cellular automata
To illustrate how our algorithm works and can be applied we will use a very well studied dynamical system called a cellular automaton. A cellular automaton is a computer program that applies in parallel a global rule composed of local rules on a tape of cells with symbols (e.g. binary). Broadly speaking, it consists of a set of states, a neighbourhood template and a local rule $f$. Here we will use space-time diagrams of cellular automata to illustrate the way in which the MILS algorithm operates.
Thoroughly studied in [1], Elementary Cellular Automata (or ECA) are defined as one-dimensional cellular automata that take into consideration in their local rules the cells next to the centre cell and the centre cell. In the case of 1-dimensional CA it is common to introduce the radius of the neighbourhood template, which can be written as $\langle-r,-r+1,\dots,r-1,r\rangle$ and has length $12r+$ 1 cells. Given a radius $r$, the local rule is a function $f:\mathbb{Z}_{|S|}^{{|S|}^{(2r+1)}}\rightarrow\mathbb{Z}_{|S|}$, where $S$ is the set of single-cell states.
Elementary Cellular Automata (ECA) have a radius $r=1$ (closest neighbours), having the neighbourhood template $\langle-1,0,1\rangle$, meaning that the three-cell neighbourhood comprises a central cell. From this it follows that the rule space for ECA contains $2562^{2^{3}}=$ 256 rules.
Enumeration of ECA rules: It is common to follow the lexicographic ordering scheme introduced by Wolfram [1]. According to this encoding, the 256 ECA rules can be encoded by 8-bits.
A space-time diagram captures the evolution of a cellular automaton for a given initial condition and is read from the top starting from time step $t=0$ (the initial condition) and evolving towards the bottom of the diagram (see Figs. 1.(A) and 1.(B) in the main article).
B.2 Classical Information Theory and Shannon Entropy
Central to information theory is the concept of Shannon’s information entropy, which quantifies the average number of bits needed to store or communicate the statistical description of an object.
Definition B.1.
Let $X=(R,p)$ be an ensemble, where $R$ is the set of all possible outcomes (the random variable), $n=|R|$ and $p(x_{i})$ is the probability of an outcome $x_{i}\in R$. The Shannon entropy of $X$ is given by
$H(X)=-\sum_{i=1}^{n}p(x_{i})\log_{2}p(x_{i}).$ (2)
It is clear from this definition that in order to calculate $H(X)$ one has to know (or assume the existence of) a probability mass function of the ensemble $X$.
While entropy can be used as a measure of the information contained in a graph (or another object), we will restrict our analyses to the Block Decomposition Method, which we describe in Section B.6.
B.3 Graph definitions and notation
Graphs (or networks) are mathematical structures that allow a simple representation of systems consisting of a set of agents and a set of interactions between them. Formally, we may define a graph as follows.
Definition B.2.
A (labeled)111 In this article, we will only deal with labeled graphs. graph is an ordered pair $G=(V,E)$ comprising a finite set $V(G)=\{1,\dots,n\}$ of labeled vertices with $n\in\mathbb{N}$ and a set $E(G)\subseteq\mathbb{E}(G)=V\times V$ of edges. For simplicity of notation, we will often let $V=V(G)$ and $E=E(G)$.
A graph $H$ is contained in $G$, and we write $H\subseteq G$, if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. In this case, we say $H$ is a subgraph of $G$. If either $V(H)\neq V(G)$ or $E(H)\neq E(G)$, we say $H$ is a proper subgraph of $G$, and we normally use $H\subset G$ to denote this situation. If on the other hand $V(H)=V(G)$ and $E(H)=E(G)$, we say the graphs $H$ and $G$ are equal.
A directed graph is a graph where the edges have a direction associated with them. In a directed graph (possibly with self-loops), we represent edges by ordered pairs
$(v,w)\in\mathbb{E}(G)\coloneqq\{(x,y)\mid x,y\in V\}=V\times V$
instead of unordered pairs as in Definition B.3, such that $v$ is the source vertex and $w$ is the target vertex of the edge. We call a traditional directed graph if $G$ is a directed graph without self-loops, so that
$(v,w)\in\mathbb{E}_{d}(G)\coloneqq\mathbb{E}(G)\setminus\{(x,x)\mid x\in V\}$
Definition B.3.
A labeled undirected graph $G=(V,E)$ without self-loops is a labeled graph with a restriction $\mathbb{E}_{c}$ in the edge set $E$ such that each edge is an unordered pair with
$E\subseteq\mathbb{E}_{c}\left(G\right)\coloneqq\{\{x,y\}\mid x,y\in V\}$
We also refer to these graphs as simple labeled (or classical) graphs. In this paper, we will normally refer to simple graphs as just graphs.
We say that two vertices $v,w\in V$ are adjacent if there is an edge $e=\{v,w\}$ in $E$ that connects them, and call $v$ and $w$ the endpoints of such undirected edge. The adjacency matrix $A=(a_{ij})$ of $G$ is a matrix that encodes all the adjacency relationships of the graph. The entries of this matrix are $a_{vw}=1$ if vertices $v$ and $w$ are adjacent and $a_{vw}=0$ otherwise. Note that while the adjacency matrix of an undirected graph is always symmetric, this is not necessarily true for a directed graph where adjacency can be asymmetric. The list of eigenvalues of the graph’s adjacency matrix sorted from largest to smallest is known as the spectrum of the graph.
The degree $d(v)$ of $v$ is the number of vertices to which $v$ is connected, that is $d(v)=\sum_{w\in V}a_{vw}$. The degree matrix $D=(d_{ij})$ of $G$ is a diagonal matrix that contains information about the degree of each vertex, and is defined by $d_{ij}=d(i)$ if $i=j$ and $d_{ij}=0$ otherwise. We are now in the position to introduce the notion of graph Laplacian, which will be of interest for us in forthcoming sections. Consider the following definition.
Definition B.4.
Let $G$ be a simple graph on $n$ vertices. The Laplacian matrix (of order $n$) of the graph $G$ is defined as $L=D-A$, where $D$ is the degree matrix and $A$ is the adjacency matrix of $G$.
Throughout the text, and trusting that it does not lead to confusion, we will use the terms nodes and vertices, and links and edges, interchangeably.
B.4 Spectral sparsification
Generally speaking, the goal of network sparsification is to approximate a given graph $G$ by a sparse graph $H$ on the same set of vertices. If $H$ is close to $G$ in some appropriate metric, then $H$ can be used as a signature preserving important properties of $G$ for faster computation after reducing the size of $G$ and without introducing too much error. Obvious trivial sparsification methods include edge deletion by some criterion, such as the outermost ones (called the $k$-shell method [22, 39], often used to identify the core and the periphery of the network), but most of these methods (such as this shell one) are rather arbitrary or ad-hoc, rather than general methods aimed at preserving important algebraic, topological or dynamical properties of the original graph, all of which constitute and contribute to the information content of the graph, that is, the information necessary to fully describe a network and reconstruct the network from that description.
A popular sparsification algorithm is the spanning tree [36] designed to preserve node distance but clearly destroy all other local node properties, such as the clustering coefficient. Not many non-trivial methods for network sparsification exist today. Some clearly destroy local properties, such as the spanning tree algorithm, which destroys the clustering coefficient. It is acknowledged [58, 59, 19], however, that spectral graph sparsification is among the most efficient, both at preserving important algebraic and dynamical properties of a network and in terms of fast calculation. In part the dearth of methods is due to a lack of assessment tools to decide whether one method is better than another in general terms, rather than being designed to preserve one or another specific graph theoretic property (e.g. the transitive edge deletion method destroys the clustering coefficient of the original graph [14]). The spectral method considered in this paper is a high-quality algorithm described in [19, 58].
Transitive reduction was introduced in [14]. A graph $G$ is said to be transitive if, for every pair of vertices $u$ and $v$, not necessarily distinct, $(u,v)\in E(G)$ whenever there is a directed path in $G$ from $u$ to $v$. That is, if there is a path from a vertex $x$ to a vertex $y$ in graph $G$, there must also be a path from $x$ to $y$ in the transitive reduction of $G$, and vice versa. If a given graph is a finite directed acyclic graph, its transitive reduction is unique, and is a subgraph of the given graph.
Graph sparsification is the approximation of an arbitrary graph by a sparse graph. We compare MILS against random, simple (e.g. spanning tree) and two powerful graph sparsification and reduction methods (spectral and transitive). Spectral graph sparsification is based on the spectral similarity of graph Laplacians. A spectral sparsifier is a subgraph of the original whose Laplacian quadratic form is approximately the same as that of the original graph on all real vector inputs. Spectral graph sparsification is a stronger notion than cut sparsifiers [59] and is considered one of the most, if not the most, sophisticated sparsification or network reduction method, as it is believed to preserve some of the most important algebraic, topological and potentially dynamical properties of a network.
B.5 Graph-theoretic measures
The global clustering coefficient of $G$ is the fraction of paths of length 2 in $G$ that are closed over all paths of length two in $G$. The mean or average clustering coefficient is the mean over all local clustering coefficients of vertices of $G$.
The betweenness centrality for a vertex $i$ in a connected graph is given by $\sum_{s,t\in v\land s\neq i\land t\neq i}\frac{n_{s,t}^{i}}{n_{s,t}}$, where $n_{s,t}$ is the number of shortest paths from s to t and $n^{i}_{s,t}$ is the number of shortest paths from $s$ to $t$ passing through $i$. The ratio $\frac{n_{s,t}^{i}}{n_{s,t}}$ is taken to be zero when there is no path from $s$ to $t$.
Degree centrality is a measure of the centrality of a node in a network and is defined as the number of edges (including self-loops) that lead into or out of the node. The degree centrality of $G$ is the list of nonnegative integers (“degree centralities") lying between 00 and $n-1$ inclusive, where $n$ is the number of vertices of $G$, and identifies nodes in the network by their influence on other nodes in their immediate neighbourhood.
Eigenvector centrality is a list of normalized nonnegative numbers (“eigenvector centralities", also known as Gould indices) that are particular centrality measures of the vertices of a graph. Eigenvector centrality is a measure of the centrality of a node in a network based on the weighted sum of centralities of its neighbours. It therefore identifies nodes in the network that are connected to many other well-connected nodes. For undirected graphs, the vector of eigenvector centralities $c$ satisfies the equation $c=1/\lambda_{1}a.c$, where $\lambda_{1}$ is the largest eigenvalue of the graph’s adjacency matrix $a$.
B.6 Algorithmic information theory and approximations to graph information content
Here, we present some basics concepts regarding algorithmic information theory. For a complete introduction to these concepts and theorems, see [34, 24, 26, 43, 18].
First, for some preliminary notation, let $\{0,1\}^{*}$ be the set of all finite binary strings. Let $|X|\in\mathbb{N}$ denote the number of elements (i.e., the cardinality) in a set, if $X$ is a finite set, or the number of elements in a data structure, if $X$ is computational object. In the particular case of strings, let $l(x)$ denote the length of a finite bit string $x\in\{0,1\}^{*}$. Let $(x)_{2}$ denote the binary representation of the number $x\in\mathbb{R}$. Let $x\upharpoonright_{n}$ denote the ordered sequence of the first $n$ bits of the fractional part in the binary representation of $x\in\mathbb{R}$. That is, $x\upharpoonright_{n}=x_{1}x_{2}\dots x_{n}\equiv(x_{1},x_{2},\dots,x_{n})$, where ${(x)}_{2}=y.x_{1}x_{2}\dots x_{n}x_{n+1}\dots$ with $y\in\{0,1\}^{*}$ and $x_{1},x_{2},\dots,x_{n}\,\in\{0,1\}$.
With respect to weak asymptotic dominance of function $f$ by a function $g$, we employ the usual $\mathbf{O}(g(x))$ for the big O notation, $\mathbf{\Omega}(g(x))$ for the big Omega notation, and $\mathbf{\Theta}(g(x))$ for the big Theta notation; and with respect to strong asymptotic dominance by a function $g$, we employ the usual $\mathbf{o}(g(x))$ when $g$ dominates $f$, ${\bm{\omega}}(g(x))$ when $f$ dominates $g$, and $\sim g(x)$ when both $f$ and $g$ dominates each other. The reader should pay attention not to confuse the big Omega notation (in boldface) $\mathbf{\Omega}(g(x))$ with the halting probability $\Omega$ as in Definition B.10.
A Turing machine $\mathbf{M}$ is a general abstraction of a computer program, similar to cellular automata but sequential, that given an input, may either produce an output (and halt) or never halt. As one of the most fundamental mathematical objects in theoretical computer science, the Turing machine is thus an algorithmically mechanistic causal explanation of an deterministic process’ outcome, and underpins the central ideas of the algorithms introduced here.
With respect to the chosen computational model, let $\mathbf{L^{\prime}_{U}}$ denote a binary prefix-free [^2] The group of valid programs constitutes a prefix-free set (no element is a prefix of any other, which is a property necessary to keep $10<\mathbf{m}(G)<$ 1). (or self-delimiting) universal programming language for a prefix universal Turing machine $\mathbf{U}$. For example, a program of $\mathbf{U}$ may represent a 2-dimensional (prefix-free) Turing machine (a typical deterministic Turing machine whose single head can move up and down as well as left and right) starting from an empty 2-dimensional grid (instead of the typical 1-dimensional tape Turing machine) as defined in [61, 68].
As in [43, 34, 18]:
Notation B.5.
[…]
$l\left(\left<x,y\right>\right)\leq l(x)+l(y)+\mathbf{O}\left(\log_{2}\left(l(x)+l(y)\right)\right)$
holds and the bit string $x,y\right>$ is computably recognizable as an ordered pair in polynomial time. For example, with an encoding in the form
$x,y\right>\coloneqq$
$\underbrace{0\cdots 0}_{\log_{2}\left(l\left(\left(2\right)_{2}\right)\right)\text{ times}}1\,\left(2\right)_{2}\underbrace{0\cdots 0}_{\log_{2}\left(\log_{2}\left(l(x)\right)\right)\text{ times}}1\,\left(l(x)\right)_{2}\,x\,\underbrace{0\cdots 0}_{\log_{2}\left(\log_{2}\left(l(y)\right)\right)\text{ times}}1\,\left(l(y)\right)_{2}\,y\text{ .}$
[…]
Definition B.6.
The (unconditional) prefix333 In this paper, we will normally refer to prefix algorithmic complexity as just algorithmic complexity. algorithmic complexity (also known as K-complexity, prefix Kolmogorov complexity, prefix program-size complexity or prefix Solomonoff-Komogorov-Chaitin complexity) of a finite binary string $w$, denoted here by $C(w)$, is the length of the shortest program $w^{*}\in\mathbf{L^{\prime}_{U}}$ such that $\mathbf{U}(w^{*})=w$444 $w^{*}$ denotes the lexicographically first $\mathrm{p}\in\mathbf{L^{\prime}_{U}}$ such that $l(\mathrm{p})$ is minimum and $\mathbf{U}(p)=w$.
In addition, when an input is informed to the Turing machine, we also have that the conditional prefix algorithmic complexity of a binary finite string $y$ given a binary finite string $x$, denoted by $C(y\,|x)$, is the length of the shortest program $w\in\mathbf{L^{\prime}_{U}}$ such that $\mathbf{U}(\left<x,w\right>)=y$. Note that $C(y)=C(y\,|\epsilon)$, where $\epsilon$ is the empty string. Moreover, the joint prefix algorithmic complexity of strings $x$ and $y$ is denoted by $C(x,y)\coloneqq C(\left<x,y\right>)$.
From algorithmic information theory (AIT) [43, 26, 34, 24], one knows the following results:
Lemma B.1.
For every $x,y\in\{0,1\}^{*}$ and $n\in\mathbb{N}$,555 For the sake of simplicity, we choose to omit the ceiling operator on the logarithm when calculating the length of encoding natural numbers, so that in such cases $\log_{2}\left(x\right)$ actually stands for $\lceil\log_{2}\left(x\right)\rceil$ in the present paper. Note that in terms of asymptotic dominance and bounds it is indeed irrelevant.
$C(x)$ $\leq l(x)+\mathbf{O}(\log_{2}(l(x)))$ (3)
$C(y\,|x)$ $\leq C(y)+\mathbf{O}(1)$ (4)
$C(y)$ $\leq C(x\,y)+C(l(x))+\mathbf{O}(1)$ (5)
$C(x\,y)$ $\leq l(y)+C(l(y))+C(x)+\mathbf{O}(1)$ (6)
$C(x)\leq C(x,y)+\mathbf{O}(1)$ $\leq\;C(y)+C(x\,|y)+\mathbf{O}(1)$ (7)
$C(n)$ $=\mathbf{O}(\log_{2}(n))$ (8)
Lemma B.2.
[…]
$C(f_{c}(n))\leq C(n)+\mathbf{O}(1)$
For graphs (as those defined in Section B.3) we use the notation:
Notation B.7.
Let $\left(e_{1},\dots,e_{\left|\mathbb{E}(G)\right|}\right)$ be a previously fixed ordering (or indexing) of the set $\mathbb{E}(G)$. For an edge set $E(G)$, let $C(E(G))\coloneqq C(\left<E(G)\right>)$, where the edge set string $E(G)\right>$ denotes
[…]
such that666 See also Definition B.8.
$z_{i}=1\iff e_{i}\in E(G)\text{ ,}$
where $z_{i}\in\{0,1\}$ with $1\leq i\leq n=|\mathbb{E}(G)|$. The same applies analogously to the conditional and joint cases of the algorithmic complexity.
From [9, 11, 12], we remember that the family of all possible simple graphs is always recursively labeled. That is, there is a unified recursive way to encode a sequence of all possible undirected edges given any unordered pair $\{x,y\}$ of natural numbers $x,y\in\mathbb{N}$. Thus, we know that, for any arbitrarily chosen method of recursively encoding a matrix into a string,
$\left|C(G)-C(E(G))\right|=\mathbf{O}(1)$
holds, where $C(G)$ denotes the algorithmic complexity of generating the adjacency matrix of any arbitrarily chosen $G$. Hence, one can interchangeably employ $C(G)$ or $C(E(G))$ to measure the algorithmic information content of a graph when dealing with the theoretical optimal form of the algorithmic complexity as in Definition B.6. Moreover, from [9, 11, 12], we will have that
$C(G\,|\,x)$ $=\mathbf{O}(1)$ (9)
$C(x\,|\,G)$ $=\mathbf{O}(1)$ (10)
$C(x)$ $=C(G)\pm\mathbf{O}\left(1\right)\text{ ,}$ (11)
where $x\in\{0,1\}^{*}$ is the characteristic string of $G$ [9]. The characteristic string of a graph is any bit string that determines the presence or absence of edges in $G$:
Definition B.8.
Let $\left(e_{1},\dots,e_{\left|\mathbb{E}(G)\right|}\right)$ be any arbitrarily fixed ordering of all possible edges of a graph $G$ as in Notation B.7. We say that a string $x\in\{0,1\}^{*}$ with $l(x)=\left|\mathbb{E}(G)\right|$ is a characteristic string of a graph $G$ iff, for every $e_{j}\in\mathbb{E}(G)$,
[…]
where $1\leq j\leq l(x)$.777 And the same definition applies analogously to the simple graph case.
Thus, we have that the algorithmic information content of a graph is equivalently captured (except for a constant that only depends on the chosen universal programming language) by the algorithmic complexity of the adjacency matrix, the edge set string, or of the characteristic string. Note that the same cannot always be said in the case of multilayered ou multidimensional networks [11, 12]. See also [9, 10] for more details on characteristic strings and algorithmically characteristic strings.
We know from the algorithmic coding theorem [26, 24, 43, 34] that
$C\left(x\right)=$ (12)
$-\log\left(\sum\limits_{\mathbf{U}\left(p\right)=x}\frac{1}{2^{\left|p\right|}}\right)\pm\mathbf{O}(1)=$
$-\log\left(\mathbf{m}\left(x\right)\right)\pm\mathbf{O}(1)\text{ ,}$
holds, where: $\mathbf{m}\left(\cdot\right)$ is a maximal computably enumerable semimeasure; and
$\mathbf{P}_{\mathbf{U}}\left[\text{``event }x\text{ occurs''}\right]=\sum\limits_{\mathbf{U}\left(p\right)=x}2^{-\left|p\right|}$
denotes the universal a priori probability of an arbitrary event. $\mathbf{P}_{\mathbf{U}}$ can be understood as the probability of randomly generating (by an i.i.d. stochastic process) a prefix-free (or self-delimiting) program that outputs $x$. A computably enumerable semimeasure $\mathbf{m}\left(\cdot\right)$ is said to be maximal if, for any other computably enumerable semimeasure $\mu\left(\cdot\right)$ with domain defined for all possible encoded objects, where $\sum\limits_{x\in\left\{0,1\right\}^{*}}\mu\left(x\right)\leq 1$, there is a constant $C>0$ (which does not depend on $x$) such that, for every encoded object $x$, $\mathbf{m}\left(x\right)\geq C\,\mu\left(x\right)\text{ .}$ We also have it that $2^{-\mathbf{K}\left(x\right)}$ is called the algorithmic probability of $x$, which therefore converges to the universal a priori probability (except of an object-independent constant). Also note that the algorithmic coding theorem applies analogously to the conditional algorithmic complexity $\mathbf{K}(z\,|w)$.
The concept of algorithmic probability (also known as Levin’s semi-measure) yields a method for approximating algorithmic complexity related to the frequency of patterns in the adjacency matrix of a network, including therefore the number of subgraphs in a network. Following from Equation (12), the algorithmic probability [57, 42, 25] of a graph $G$ is a measure that describes the probability that a random computer program $p$ will produce $G$ when run on a 2-dimensional tape universal (prefix-free) Turing machine $\mathbf{U}$. That is,
$\mathbf{m}(G)=\sum_{\mathbf{U}(p)=A_{G}}1/2^{|p|}\text{ .}$
An example of a popular 2-dimensional tape Turing machine is Langton’s ant [41], commonly referred to as a Turmite.
The Algorithmic Coding Theorem [26, 43, 34, 24] therefore establishes the connection between $\mathbf{m}(G)$ and $C(G)$ as:
$|-\log_{2}\mathbf{m}(G)-C(G)|<c\text{ ,}$ (13)
where $c$ is some fixed constant, independent of $G$. This way, the size $C(G)$ of the algorithmic information content of a graph $G$ (i.e., its prefix algorithmic complexity) is given by the so-called algorithmic coding theorem [42, 30, 56] by simply rewriting Eq. 13 as:
$C(G)=-\log_{2}\mathbf{m}(G)\pm\mathbf{O}(1)\text{ .}$ (14)
In [32] a technique called Coding Theorem Method (CTM) was advanced for approximating $\mathbf{m}(G)$ by means of a function that considers all Turing machines of increasing size (by number of states). Indeed, for Turing machines $\mathbf{M}$ with $n$ states, $k$ alphabet symbols (usually 2 alphabet symbols only), and empty input $\lambda$ in which $n$ is small enough, we have that
$\textit{D}(n,k)(s)=\frac{\left|\left\{(\mathbf{M}\in\left(n,k\right)\colon\mathbf{M}\left(\lambda\right)=s\right\}\right|}{\left|\left\{(\mathbf{M}\in\left(n,k\right)\colon\exists w^{\prime}\left(\mathbf{M}\left(\lambda\right)=w^{\prime}\right)\right\}\right|}$ (15)
is computable for values of the Busy Beaver problem [54] that are known, providing a means to numerically approximate the algorithmic complexity of small graphs, such as network motifs. In Equation (15), $\mathbf{M}\left(x\right)=y$ denotes machine $\mathbf{M}$ returning $y$ with input $x$ and, therefore, $\textit{D}(n,k)(s)$ gives the frequency of halting Turing machines with $n$ states and $k$ alphabet symbols that ended up returning $s$ as output. The algorithmic coding theorem then establishes that graphs produced with lower frequency by random computer programs have higher algorithmic complexity, and vice versa.
In this article, we will use the Block Decomposition Method (BDM) [68, 61, 72] as an estimator of algorithmic complexity C(G) and, hence, of the information difference $I(G,e_{i})$, but the algorithm and methods introduced are independent of the particular method used to approximate algorithmic complexity. The BDM consists in decomposing the adjacency matrix of a graph into subgraphs of sizes for which complexity values have been estimated, then reconstructing an approximation of the algorithmic complexity of the graph by adding the complexity of the individual pieces according to the rules of information theory, as follows:
$K_{log}(G)=\sum_{(r_{u},n_{u})\in Adj(G)_{d\times d}}\log_{2}(n_{u})+K_{m}(r_{u})\text{ ,}$ (16)
where $Adj(G)_{d\times d}$ represents the set with elements $(r_{u},n_{u})$ obtained when decomposing the adjacency matrix of $G$ into all subgraphs of size $d$ contained in $G$ and $K_{m}(r_{u})$ is an approximation to $C(r_{u})$ using the CTM [32, 56, 55, 62] from Equations (14) and (15). In each $(r_{u},n_{u})$ pair, $r_{u}$ is one such submatrix of the adjacency matrix and $n_{u}$ its multiplicity (number of occurrences). As can be seen from the formula, repeated subgraphs only contribute to the complexity value with the subgraph BDM complexity value once plus a logarithmic term as a function of the number of occurrences. This is because the information content of subgraphs is only sub-additive, as one would expect from the growth of their description lengths. Applications of CTM and BDM have been explored in [32, 56, 55, 62], and include applications to graph theory and complex networks [61] and [62], where the technique was first introduced.
The only parameters used in the application of BDM are the use of strings up to 12 bits for strings and 4 bits for arrays given the current best CTM approximations [56] and the suggestions in [68] based on an empirical distribution based on all Turing machines with up to 5 states, and with no string/array overlapping in the decomposition for maximum efficiency (as it runs in linear time), and for which the error (due to boundary conditions) has been shown to be bounded [68].
Definition B.9.
We say a real number $x\in\left[0,1\right]\subset\mathbb{R}$ is 1-random (K-incompressible up to a constant, K-random or prefix algorithmically random) if and only if it satisfies
$C(x\upharpoonright_{n})\geq n-\mathbf{O}(1)\text{ ,}$
where $n\in\mathbb{N}$ is arbitrary.
In order to avoid ambiguities between plain and prefix algorithmic complexity and ambiguities in relation to randomness deficiencies, we choose to say that an algorithmically random real number with respect to prefix algorithmic complexity in Definition B.9 is $\mathbf{O}(1)$-K-random.
One of the most important results in algorithmic information theory is the investigation and proper formalisation of a mathematical theory for randomness [43, 34, 24]. We will employ one of these important mathematical objects: the halting probability (see Definition B.10). This is a widely known example of infinite binary sequence, or real number, that is algorithmically random with respect to prefix algorithmic complexity, i.e., $\mathbf{O}(1)$-K-random.
Definition B.10.
Let $\Omega\in\left[0,1\right]\subset\mathbb{R}$ denote the halting probability (also known as Chaitin’s constant or Omega number). The halting probability is defined by
$\Omega=\sum\limits_{\begin{subarray}{c}\exists y\left(\mathbf{U}(\mathrm{p})=y\right)\\ \mathrm{p}\in\mathbf{L^{\prime}_{U}}\end{subarray}}\quad\frac{1}{2^{l(\mathrm{p})}}$
Theorem B.3.
Let $n\in\mathbb{N}$. Then,
$C(\Omega\upharpoonright_{n})\geq n-\mathbf{O}(1)$
That is, $\Omega$ is $\mathbf{O}(1)$-K-random.
The same analogous notion of incompressibility can be assigned to classical graphs (see Definition B.3) as in [9, 10]:
Definition B.11.
We say a simple graph $G$ is $\mathbf{O}(1)$-K-random iff it satisfies
$C(E(G))\geq\binom{\left|V(G)\right|}{2}-\mathbf{O}(1)$
This way, we have from [9] that:
Corollary B.4.
There is an infinite number of classical graphs $G$ (as in Definition B.3) that are $\mathbf{O}(1)$-K-random. In particular, those in which their characteristic strings are long enough initial segments of a $\mathbf{O}(1)$-K-random real number.
B.7 Information difference
The information difference is based on a concept of algorithmic/causal perturbation analysis as introduced in [61, 64, 72]. The procedure consists in the perturbation of all elements of a system by the removal of elements whose effects on its algorithmic information content are measured and ranked accordingly. Formally:
Definition B.12.
Let $G$ be a graph with edges $e_{1},\ldots,e_{|E(G)|}\in E(G)\subseteq\mathbb{E}(G)$. Now, let $G\backslash e_{i}$ denote the graph obtained by removing $e_{i}$ from $G$, and $I(G,e_{i})$ denote the information difference or information value/contribution of $e_{i}$ to $G$ given by
$I(G,e_{i})\coloneqq C(G)-C(G\backslash e_{i})\text{ ,}$
where $C(G)$ is the algorithmic complexity of graph $G$. If one allows multiple edge deletions of a subset $F\subseteq E(G)$, we analogously define
$I(G,F)\coloneqq C(G)-C(G\setminus F)\text{ .}$
Also, for the analogously inverse operation of edge insertion $G\cup e_{i}$, we denote the information difference by
$I^{-1}(G,e_{i})\coloneqq C(G)-C(G\cup e_{i})$
and
$I^{-1}(G,F)\coloneqq C(G)-C(G\cup F)\text{ ,}$
respectively.
These differences can also be referred to the graph’s nodes, but here we will restrict our attention to the edges. In addition, we focus our results on destructive perturbations, i.e., the edge deletion problem. Nevertheless, we will investigate in the next Section C.1 the edge-insertion case (i.e., edge insertion) as an analogous inverse case of edge deletion. This will become most relevant in the proof of Theorem C.1.
When taking the difference $C(G)-C(G\backslash e_{i})$ by itself we will refer to it as the graph (dis)similarity between graph $G$ and $G\backslash e_{i}$. $I(G,e_{i})$ applied to graphs suggests a similarity “distance” between graphs based on algorithmic information content (in [63], we show that this similarity measure can classify networks by the family they belong to, differentiating variant synthetic and natural network topologies similar to graph motifs, as shown in [49]).
In the description of the algorithm to be presented in Section B.8, replacing the underlying methods to approximate the algorithmic information content (which is the theoretically optimal measure of information content) by, e.g., Shannon entropy or lossless compression algorithms represents special cases of the more general algorithm based on algorithmic complexity, and thus it covers all these less powerful cases. The idea of a dynamic study/calculus of the (possible) changes that can be wrought upon an object to evaluate the contribution of each of its components for different purposes was introduced in [72, 71], and here we extend these ideas to the area of data/network dimension reduction.
Note that MILS can actually be applied to any general object and only to graphs or networks. Let $\mathbb{U}$ be an object (possibly multidimensional, like those in [11, 12]) and $I(\mathbb{U},u_{i})$ be the information difference of an element $u_{i}\in\mathbb{U}$ to $\mathbb{U}$, defined by $I(\mathbb{U},u_{i})=C(\mathbb{U})-C(\mathbb{U}\backslash u_{i})$ [60]. If $u_{i}$ is a neutral information element, then by definition it does not contribute to the algorithmic information content of $\mathbb{U}$, and so $I(\mathbb{U},u_{i})\simeq\log|\mathbb{U}|$, which indicates that $u_{i}$ is part of the dynamical causal path of $\mathbb{U}$ and $\mathbb{U}$ can regenerate $u_{i}$ by its normal dynamical course. (See Theorem C.8 for an investigation of neutral elements in graphs in general, and the existence of maximally neutral elements). In general, however, if there is no element $u_{i}$ such that $I(\mathbb{U},u_{i})\simeq\log|\mathbb{U}|$—which may be the case found in practice due to the semi-computability of algorithmic complexity—, there is always the most neutral element (possibly not unique, in case which heuristics such as InfoRank in Algorithm 2 can deal with) in $\mathbb{U}$ to be removed in the application of MILS. Moreover, for the most neutral element $u_{i}$, we will have it that
$\left|I(\mathbb{U},u_{i})\right|=\min\{|{C(\mathbb{U})-C(\mathbb{U}\backslash u_{i})}|\sim\log|\mathbb{U}|)\}\text{ ,}$
that is, the element with an information contribution closest to $\log|\mathbb{U}|$. This result can be directly obtained by an analogous argument to those in Section C.1.
B.8 An algorithm for minimal information loss selection (MILS) and its universal decision-probability tree diagram
MILS is an unsupervised and mostly parameter-free algorithm, i.e., asymptotically independent of model or domain, as it does not need to be instructed or designed to preserve any particular property, and maximises the preservation of all computable elements that contribute to the algorithmic information content of the data.
Let $G$ be a graph and $\left|E\left(G\right)\right|$ denote its number of edges. Ideally, MILS seeks to reduce $G$ to a graph on $N<\left|E\left(G\right)\right|$ edges so that the loss of information is minimized. It does this by calculating the information differences between $G$ and $G\backslash S$, where $\left|S\right|\geq 1$, and then finding the subset $F\subset E(G)$ such that $I(G,F)\leq I(G,S)$ for all non-empty proper subsets of edges $S\subset E(G)$, repeating this task with $G\coloneqq G\backslash F$ until the target size is reached. The algorithm’s time complexity is clearly in $O(exp)$ because of how MILS performs searches and deletions across all subsets of edges, but significant improvements to this bound are still possible, as we will show in this article. One can easily modify the following algorithms in order to deal with multiple edge deletions at once, as in $I(G,F)$ to the above, but possibly at the expense of much more computational resources.
A more efficient but suboptimal version of such an exponential-time MILS iterates over single elements (in the case, edges) or singletons. Algorithm 3 performs such a sequential deletions of edges, and then removes the edge that contributes less to the information content of the graph before moving to the next step.
When $e$ is such that $\left|I(G,e)\right|$ is minimal, we call it a neutral information edge [72] because it is the edge that contributes less to the information content of $G$ (in particular, it minimizes information loss or the introduction of spurious information into the network according to the information difference when removed from the original network). The pseudocode in Algorithm 3 assumes that there is a unique such $e$, which may not necessarily be the case. Algorithm 2 solves this problem by performing simultaneous perturbations on all edges with an information contribution given by those in minLoss, which contains the neutral information edges. To this end, Algorithm 2 also introduces InfoRank, a method that produces a ranking of $e_{1},\dots,e_{|E(G)|}$ from least informative to most informative edge, i.e., a list of edges sorted in increasing order by their information contribution to $G$. Therefore, note that unlike minLoss in Algorithm 3, the set minLoss in Algorithm 2 can have more than one edge. This ranking facilitates the search for the (most) neutral elements of the system (see Section C.3), which in turn helps MILS preserve the components that maximise the information content of the resulting object. (As described in [72], note that Algorithms 3 and 2 may be applied, mutatis mutandis, to nodes or to any data element of an object or dataset, e.g. a pixel of an image, or a row or column in a spreadsheet).
Algorithm 2 Minimal Information Loss Selection (MILS)
1:function InfoRank($G$)
2:     $\textit{informationLoss}\leftarrow\varnothing$
3:      // for each edge $e_{i}$
4:     for $e_{i}\in E(G)$ do
5:          // store information contribution into informationLoss
6:         $\textit{informationLoss}\leftarrow\textit{informationLoss}\cup\{I(G,e_{i})\}$
7:     end for
8:     sort informationLoss in increasing order
9:      // return information rank
10:     return informationLoss
11:end function
12:function MILS($G,N$), $1\leq N\leq|E(G)|$
13:     while $|E(G)|>N$ do
14:          // calculate minimal loss across all edges
[…]
16:          // remove all candidate edges from $G$
17:         $\textit{G}\leftarrow G\backslash\{e_{i}\in E(G):I(G,e_{i})=\textit{minLoss}\}$
18:     end while
19:     return $G$
20:end function
Algorithm 3 MILS with sequential perturbation by deletion on edges
1:function Perturb($G$)
2:     $\textit{informationLoss}\leftarrow\varnothing$
3:      // for each edge $e$
4:     for $e\in E(G)$ do
5:          // store information contribution into informationLoss
6:         $\textit{informationLoss}\leftarrow\textit{informationLoss}\cup\{I(G,e)\}$
7:     end for
8:      // calculate minimal information loss across all edges
9:     $\textit{minLoss}\leftarrow\min({informationLoss})$
10:     for $e\in E(G)$ do
11:          // if $e$ is a neutral edge
12:         if $I(G,e)=\textit{minLoss}$ then
13:               // remove $e$
14:              return $G\backslash e$
15:         end if
16:     end for
17:end function
18:function Sparsify($G,N$), $1\leq N\leq|E(G)|$
19:      // until reaching target size
20:     while $|E(G)|>N$ do
21:          // perturb edges in $G$
22:         $G\leftarrow\textsc{Perturb}(G)$
23:     end while
24:     return $G$
25:end function
Fig. 11 presents three illustrative examples of Algorithm 2 returning the same graph (with $N\leq 2$ present edges) as output. The topmost matrices in columns A, B, and C display the adjacency matrices of distinct directed graphs (without self-loops), all of them with four vertices. These graphs are subjected to the edge deletions that Algorithm 2 performs so as to generate its final output. The edge deletions occur in each iteration $j$ of [Algorithm 2, Step 13] which is indicated by the downward blue arrows (except for the bottommost blue arrows), while the subsequent red squares (i.e., the matrix entries) indicate where the deletions occurred in the previous step. The bottommost blue arrows in each Fig. 11A, 11B, and 11C refer to [Algorithm 2, Step 19] after which the final output of the algorithm is returned, given an adjacency matrix of a graph $G$, such as those three topmost $4\times 4$ matrices, as input.
The input adjacency matrix in Fig. 11A corresponds exactly (by concatenating its rows) to the $16$-bit-length substring that starts at the $17$-th bit of the first $64$ bits of the halting probability calculated for a particular Turing machine in [23]. Notice that this substring was chosen for illustrative purposes so that: its string is proven to be algorithmically random (i.e., incompressible); it starts at the next bit after the first $4^{2}$ bits of its original string, where $16$ is equal to the length of the substring itself and, therefore, its starting position at the original string neither adds nor subtracts information from the algorithmic complexity of the substring; the resulting diagonal is empty, which corresponds to an adjacency matrix of a graph without self-loops; and the number of present edges is almost the same as absent edges, as it is expected to occur in sufficiently long segments of an incompressible number. Since the input graph in Fig. 11A is (algorithmically) random, we can assume for illustrative purposes that a $16$-bit-length random string is sufficiently long in comparison to a reference universal machine. In this scenario, every edge becomes non-redundant. Then, no edge in the (hierarchical) list given by InfoRank will be as neutral as the other ones. Therefore, minLoss will contain only one edge at the time in each iteration $j$ of [Algorithm 2, Step 13], resulting in $3$ iterations of deletions necessary to reach the desired target of $N\leq 2$ edges presented in Fig. 11A. As the adjacency matrix given as input becomes less random, one should expect that the number of iterations can be reduced. Indeed, this is what happens in Fig. 11B and 11C.
Fig. 11B presents a slight variation of the input matrix in Fig. 11A. Unlike in the latter case, the second row is a repetition of the first row, that is, the matrix corresponds to a sequence in which the first $4$ bits are repeated and the remaining $8$ bits equals to those of the input matrix in Fig. 11A. For illustrative purposes, we can assume that the algorithmic complexity of the program that repeats a given $4$-bit-length string is sufficiently small in comparison to the algorithmic complexity of the $16$-bit-length string that corresponds to the random input matrix of Fig. 11A. In this case, we will have that InfoRank will consider the information difference resulting from the deletion of the edge (i.e., the deletion of the second black square) in the second row within the same ballpark of the information difference resulting from the deletion of the edge in the first row. In other words, the first edge is a neutral information edge as much as the second edge is. Therefore, both edges will be deleted in only one iteration step, namely $j=1$. Notice that the subsequent remaining $8$ bits are exactly those of Fig. 11A, subsequence which is random. Then, one iteration will be required for each edge deletion. Since $N\leq 2$ was the input parameter of Algorithm 2, just one extra black square will have to be deleted in step $j=2$ before the algorithm returns the output.
Fig. 11C presents one of the most redundant cases in which all oriented edges are present, that is, the input is an adjacency matrix of a complete (directed) graph. In accordance with the illustrative assumptions for Fig. 11A and Fig. 11B, we can also assume that the algorithmic complexity of the program that repeats every subsequent bit other than the ones in the diagonal of the matrix is sufficiently small. Notice that the last $3$ bits in the last row corresponds to a subsequence of a random sequence of Fig. 11A. Therefore, except for the last two edges in the last row—two edges which belong to the last $3$ bits in the last row corresponding to a subsequence of a random sequence used in cases A and B—, every edge will be considered neutral in InfoRank; and at the first iteration step ($j=1$), all of these will belong to minLoss.
Figure 11: Three examples of edge deletions occurring in each iteration $j$ of [Algorithm 2, Step 13], indicated by the downward blue arrows (except for the bottommost blue arrows), while the subsequent red squares indicate where the deletions occurred in the previous step. The bottommost blue arrows in each Fig. 11A, 11B, and 11C refer to [Algorithm 2, Step 19] after which the final output of the algorithm is returned, given an adjacency matrix of a graph $G$ (such as anyone of the three topmost $4\times 4$ matrices) as input. The input adjacency matrix in Fig. 11A corresponds exactly to a substring of the halting probability as in [23]. Mixing both redundant and incompressible components, the second row of the input matrix in Fig. 11B is a repetition of the first row, and the remaining $8$ bits equals to those of the input matrix in Fig. 11A. Fig. 11C presents one of the most redundant cases in which all oriented edges are present, that is, the input is an adjacency matrix of a complete (directed) graph.
In Section C.3 we will show that Algorithm 2 is deterministic, polynomially time bounded, and describes a criterion to select and remove the neutral elements of an object. We employ this more efficient version in all our experiments, and even in this limited form the procedure excels at preserving important characteristics of the networks under study. See Theorem C.11.
B.8.1 A decision-probability tree diagram for MILS
The tree diagram in Fig. 12 illustrates the decision process involved in each iteration of Algorithm 2 that results in the next size-reduced graph.
MILS Tree Diagram
Figure 12: Tree Diagram for Algorithm 2 with the sequence of graphs with their respective subsets of edges reduced (as selected in the iterations $j$ of [Algorithm 2, Step 13]) highlighted in the color blue. So, one has that $F_{i_{1}}=\textit{minLoss}=\min\left(\textsc{InfoRank}\left(G\right)\right)$, $F_{i_{2}}=\textit{minLoss}=\min\left(\textsc{InfoRank}\left(G\setminus F_{i_{1}}\right)\right)$, and so on, which continues up until the final graph $G\setminus F_{i_{h}}$ is found by the algorithm. This process is represented by the blue path $\left(G,G\setminus F_{i_{1}},G\setminus F_{i_{2}},\dots,G\setminus F_{i_{h}}\right)$ from the root $G$ to the leaf $G\setminus F_{i_{h}}$ in $h$ iterations. Notice that from each parent vertex/graph, there are as many children as non-empty subsets of (present) edges in the parent. Fig. 12 illustrates a speed-up both globally (since the number of iterations $j\leq h$ is always upper bounded by $\mathbf{O}\left(\left|E\left(G\right)\right|-N\right)$) and locally (since from each parent, Algorithm 2 does not have to search for the best candidate across all possible subsets of edges of the parent) achieved by Algorithm 2 from the decision process that this tree diagram represents.
Each child vertex represents a graph resulting from the deletion of one or more edges in the parent vertex (i.e., a graph before the destructive perturbations on the edges). In each descending step starting from the root at the top, as the height of the tree downward increases, a subset of (present) edges is deleted from the graph at the lower level, that is, a deletion from the parent vertex/graph closer to the root vertex. Notice that from each parent vertex, there can be as many child vertices/graphs as the size of the power set $\mathcal{P}\left(E(G)\right)$ of the edges in the graph represented by the parent vertex. The child vertex chosen by Algorithm 2 is the one whose neutral information edges (according to InfoRank) were deleted. In this manner, the iteration process in [Algorithm 2, Step 13] is represented by the rooted path highlighted in the color blue in the Fig. 12. A subset of edges that has neutral information with respect to the parent graph is deleted from the parent graph in order to generate the child graph in this path.
InfoRank allows Algorithm 2 to speed up the selection of the child vertex that (in general, suboptimally) minimises the information difference with respect to the parent vertex, not having to search over all the possible subsets of present edges in the parent vertex. This constitutes a local speed-up between each parent vertex and its children in Fig. 12, in particular replacing a search over an exponential-size set of subsets with a search over a list (of edges), therefore quadratic with the number of vertices/nodes in the root graph. For the illustrative purposes of Fig. 12, one can take the arrow labelings in the form $i_{j}$ so as to correspond to the exact sequence of subsets of edges that Algorithm 2 chose to delete from the output of InfoRank. (Notice that in practice one would in general start with a fixed labeling, and thereafter track the rooted path in the tree that corresponds to the exact sequence of subsets of edges that Algorithm 2 ended up choosing to delete). Thus, since Algorithm 2 starts by deleting the first element in the list given by InfoRank, and possibly continuing the deletions until subsequently exhausting every neutral information edge in the ordered list, one will have only $\left|E\left(G\right)\right|$ possible choices of subsets of edges to delete from the output of InfoRank. This is displayed in Fig. 12 by the search interval comprising the set $\left\{x\middle|i_{j}-k\leq x\leq i_{j}+k\right\}$, for each iteration $j$ with $2k+1\coloneqq\left|E\left(G^{\prime}\right)\right|$,888 Or $12k+1\coloneqq\left|E\left(G^{\prime}\right)\right|+$ 1, if $\left|E\left(G^{\prime}\right)\right|$ is an even number in which case one can assume $i_{j}-k$ to correspond to the empty subset just to keep the symmetry of the interval around $i_{j}$ for illustrative purposes. where $G^{\prime}$ is any graph. By illustrating the speed-up aforementioned in this paragraph, no matter how many subsets of edges a parent graph has, the width of the interval between $i_{j}-k$ and $i_{j}+k$ remains999 Except for including or not the empty set to account for the eventual even value of $\left|E\left(G\right)\right|$. $\left|E\left(G^{\prime}\right)\right|\leq\left|E\left(G\right)\right|$ across all the iterations $j$ for every child graph $G^{\prime}$ until the total height $h$ of the tree rooted in graph $G$ is reached by Algorithm 2.
The single-thread process (between $j=0$ and $j=h$) highlighted in blue in Fig. 12 continues until the leaf vertex, i.e., the graph with $\leq N$ edges is achieved by Algorithm 2. This constitutes a global speed-up between the root vertex and the final output of Algorithm 2, particularly replacing a search over all possible rooted paths with a recursive iteration in [Algorithm 2, Step 13] whose maximum number of iterations is always upper bounded by $h\leq\left|E\left(G\right)\right|-N+1$, therefore quadratic with the number of vertices in the root graph. This makes the height of the tree always much shorter in comparison to its width.
It is also important to notice that the decision tree diagram (i.e., the one for the minimisation of the information difference between the parent and the child vertex) in Fig. 12 can be subsumed into a probability tree diagram. Each path from the root to a leaf has an associate joint probability of a sequence of independent events as defined in Equation (23), and each event corresponds to a destructive perturbation on a subset of edges. The rooted path highlighted in the color blue in the Fig. 12 illustrates the path to the final leaf vertex (i.e., the vertex representing the outcome of Algorithm 2) that also maximises the (algorithmic) probability of edge deletions, while each edge in this path corresponds to an iteration in [Algorithm 2, Step 13].
Now, one can demonstrate that the MILS tree diagram in Fig. 12 shows that a decision process of minimisation of information differences is subsumed into a maximization of (universal algorithmic) probabilities. In turn, it evinces the generality of our methods (based on the foundational contributions of algorithmic information theory). This follows from the algorithmic coding theorem [26, 24, 43, 34] as in Equation (12).
To this end, from Definition B.12, first note that from each parent graph $G$ (which is the graph before the destructive perturbations take place) and each child graph $G\setminus F$ (the one after a destructive perturbation on the subset $F$ of present edges occurs) there is a unique value
$\left|I(G,F)\right|=\left|C(G)-C(G\setminus F)\right|\text{ .}$ (17)
Let $\mathfrak{T}$ be a tree in which each arrow (labeled $i_{j}$) from $G$ to $G\setminus F$ has a probability measured by $\mathbf{P}$ as defined by
$\mathbf{P}\left[G\to G\setminus F\right]=c_{G}\,\left(\frac{1}{2^{\left|C(G)-C(G\setminus F)\right|}}\right)\text{ ,}$ (18)
where $c_{G}$ is a normalising constant that only depends on $G$ so that
$c_{G}\coloneqq\sum\limits_{F\in\mathcal{P}\left(E(G)\right)}\frac{1}{2^{\left|C(G)-C(G\setminus F)\right|}}\text{ .}$ (19)
Let $\mathfrak{p}_{F}^{-}$ denote the destructive algorithmic perturbation [13] that generates $G\setminus F$ from $G$, and $\mathfrak{p}_{F}^{+}$ the constructive perturbation that realises the inverse operation. Therefore, from the results in algorithmic information theory (AIT) discussed in Section B.6, one can straightforwardly demonstrate the inequality
$\frac{c_{G}}{2^{\max\left\{C(\mathfrak{p}_{F}^{-}),C(\mathfrak{p}_{F}^{+})\right\}}}\leq\mathbf{P}\left[G\to G\setminus F\right]\leq 1\text{ .}$ (20)
Hence, one also has that
$c_{G}\,\left(\sum\limits_{F\in\mathcal{P}\left(E(G)\right)}\frac{1}{2^{\max\left\{C(\mathfrak{p}_{F}^{-}),C(\mathfrak{p}_{F}^{+})\right\}}}\right)\leq\sum\limits_{F\in\mathcal{P}\left(E(G)\right)}\mathbf{P}\left[G\to G\setminus F\right]=1$ (21)
holds. Thus, each rooted path $G\to^{*}G_{\left(h,i_{h}\right)}$ (from the root $G$ to a leaf $G_{\left(h,i_{h}\right)}$) is a sequence
$\left(G,G_{\left(1,i_{1}\right)},G_{\left(2,i_{2}\right)},\cdots,G_{\left(h,i_{h}\right)}\right)$ (22)
of graphs with an associate joint probability (of independent events, i.e., independent perturbations $\mathfrak{p}_{F_{i_{j}}}^{-}$) given by
$\mathbf{P}\left[G\to^{*}G_{\left(h,i_{h}\right)}\right]=\prod\limits_{\scriptsize\begin{array}[]{c}1\leq j\leq h\text{,}\\ \left|\left\{i_{k}\middle|k=j\right\}\right|=\sum\limits_{i_{j-1}}\left|\mathcal{P}\left(E\left(G_{\left(j-1,i_{j-1}\right)}\right)\right)\right|-1\end{array}}\mathbf{P}\left[G_{\left(j-1,i_{j-1}\right)}\to G_{\left(j,i_{j}\right)}\right]\text{ ,}$ (23)
where $G_{\left(0,0\right)}\coloneqq G$ , $G_{\left(j,i_{j}\right)}\coloneqq G\setminus F_{i_{j}}$, |$\left|\left\{i_{j}\middle|1\leq j\leq h\right\}\right|=\sum\limits_{1\leq k\leq h}\left|\left\{i_{k}\middle|k=j\right\}\right|$, and $\left|E\left(G_{\left(h,i_{h}\right)}\right)\right|=\left|E\left(G\setminus F_{i_{h}}\right)\right|\leq N$. Notice that if $G_{\left(j-1,i_{j-1}\right)}\to G_{\left(j,i_{j}\right)}$ is a arrow in the tree, then $F_{i_{j}}\subseteq F_{i_{j-1}}$.
Now, one can employ the algorithmic coding theorem (as in Equation (12)) to demonstrate in Theorem B.1 that Equation (23) is an upper bound (except for an independent constant) for any other arbitrarily chosen (computable) probability (semi-)measure $\mathbf{P}^{\prime}$. In other words, for any arbitrary $\mathbf{P}^{\prime}$ capable of calculating the probability of any path in the MILS tree diagram—even if such a probability is not defined upon the assumption of independence of the subsequent perturbations like $\mathbf{P}$ is—, if a certain path maximises $\mathbf{P}^{\prime}$, then $\mathbf{P}$ in Equation (23) is also maximised. This results show that a decision tree diagram becomes subsumed into a probability tree diagram in the theoretical optimal limit no matter which probability one chooses to be maximised, which is a result that highlights the robustness of the methods employed in the present paper. In addition, Theorem B.1 also shows that the independence of events assumed in $\mathbf{P}$ does not make one loose any generality in comparison to other probability measures that one tries to maximise.
Notice that the algorithmic coding theorem promptly demonstrates a universal/maximal probability semimeasure for minimising either $C(G)$ or $C\left(G\setminus F\,\middle|\,G\right)$. Theorem B.1 then contributes by showing that this does not only holds for the unconditional and conditional classical cases directly, but also for the minimization of Equation (17) across the whole decision tree of subsequent independent events, which in turn implies the maximisation of Equation (23) across the probability tree diagram. Thus, we will prove that the algorithmic-probability-based method can be extended to Equation (17); additionally, we shall prove that this applies both globally to
$\mathbf{P}\left[G\to^{*}G_{\left(h,i_{h}\right)}\right]$
(for a path $G\to^{*}G_{\left(h,i_{h}\right)}$ between the root and a leaf) and locally to every child vertex and tree arrow along such a path; and the assumption of independent events in Equation (23) does not imply any loss of generality in the theoretical asymptotic limit. This result shows the generality and theoretical robustness of the methods based on algorithmic information theory: in particular, a decision tree based on minimisation of an index value is proven to be subsumed into a maximisation of path probabilities of (subsequent independent) destructive perturbations; in addition, this holds agnostically with respect to the choice of the computation model, programming language, or another arbitrary (computable) probability (semi-)measure that one aims at maximising (whether or not one assumes independence between parent vertices and subsequent child events in such a measure).
Theorem B.1 (Decision-Probability Tree Diagram).
[…]
$\mathbf{P}^{\prime}_{j}\left[G\to^{*}G\setminus F_{i_{j}}\right]\leq\epsilon\,\mathbf{P}\left[G\to^{*}G\setminus F_{i_{j}}\right]$ (24)
holds for every graph $G\setminus F_{i_{j}}$ in the path from $G$ to $G\setminus F_{i_{h}}$, where $j\leq h$.
Proof.
By assumption, we have that $\mathbf{P}^{\prime}_{j}\left[G\to^{*}G\setminus F_{i_{j}}\right]$ is defined for every $i_{j}\in\left\{i_{j}\middle|1\leq j\leq h\right\}$. In this way, one can employ an argument by induction on $j$. First, we have that $\mathbf{P}^{\prime}_{1}\left[G\to G\setminus F_{i_{1}}\right]$ is well defined for every $i_{1}\in\left\{i_{k}\middle|k=1\right\}$. Then, from Equation (12), we know that there is a constant $\epsilon_{1}>0$ such that
$\mathbf{P}^{\prime}_{1}\left[G\to G\setminus F_{i_{1}}\right]\leq\epsilon_{1}\mathbf{P}\left[G\to G\setminus F_{i_{1}}\right]$ (25)
holds. Suppose now that for every $m\leq n$ and every $i_{m}\in\left\{i_{k}\middle|k=m\right\}$, there is a constant $\epsilon_{m}$ such that
$\mathbf{P}^{\prime}_{m}\left[G\to^{*}G\setminus F_{i_{m}}\right]\leq\epsilon_{m}\mathbf{P}\left[G\to^{*}G\setminus F_{i_{m}}\right]$ (26)
holds. Since $\left\{\mathbf{P}^{\prime}_{j}\right\}$ is a computable list by hypothesis and $\mathbf{P}^{\prime}_{n+1}$ is well-defined, there would be another computable measure $\mathbf{P}^{\prime\prime}_{n+1}\left[G\setminus F_{i_{n}}\to G\setminus F_{i_{n+1}}\right]$ that is well-defined for every $i_{n}$ and $i_{n+1}$ such that
$\mathbf{P}^{\prime\prime}_{n+1}\left[G\setminus F_{i_{n}}\to G\setminus F_{i_{n+1}}\right]=c_{n}\,\frac{\mathbf{P}^{\prime}_{n+1}\left[G\to^{*}G\setminus F_{i_{n+1}}\right]}{\mathbf{P}^{\prime}_{n}\left[G\to^{*}G\setminus F_{i_{n}}\right]}$ (27)
holds, where $c_{n}$ is a normalising constant. Therefore, from Equation (12), we would also have that there is a constant $\epsilon_{n+1}$ such that
$\mathbf{P}^{\prime\prime}_{n+1}\left[G\setminus F_{i_{n}}\to G\setminus F_{i_{n+1}}\right]\leq\epsilon_{n+1}\mathbf{P}\left[G\setminus F_{i_{n}}\to G\setminus F_{i_{n+1}}\right]$ (28)
holds. Finally, the desired proof then follows directly by making
$\epsilon\coloneqq\max\left\{\epsilon_{k}\middle|1\leq k\leq h\right\}\text{ .}$
Thus, as proven in this Section B.8.1 (see also the forthcoming Section C.3), MILS is, by design, optimal in the computability and algorithmic-information theoretic sense, and only dependent on the method for approximating algorithmic complexity in the preservation of any possible feature of interest that contributes to the (algorithmic) information content of a network such as its degree distribution and other graph-theoretic, algebraic or topological features—even including those not necessarily captured by any graph theoretic measure or classical information approach [66, 67].
B.8.2 PyMILS and variations of MILS
Algorithm 4 Python implementation of the Minimal Information Loss Selection algorithm (PyMILS)
1:function pymils($image$)
2:     while $size(image)>size(image)_{min}$ do
3:         $bdm_{initial}=BDM(image)$
4:         Randomly select one of the following: $(rows,columns)$
5:         For some chunk size $c$, randomly select clustered groups of size $c$ of either rows or columns
6:         for each clustered group, do
7:              remove that group to get $image_{step}$
8:              measure $bdm_{step}=BDM(image_{step})$
9:              replace that group
10:         end for
11:         Remove the clustered group from $image$ that would result in a $bdm_{step}$ closest to $bdm_{initial}$.
12:     end whilereturn final image
13:end function
We also present a variation of Algorithm 2 that instead resorts in heuristically allowing a random selection of elements when they have the same information value. Notice that unlike Algorithms 3 and 2 that always tries to minimise information loss in a finer-grained fashion, the random-sampling variation in Algorithm 4 blindly chooses the elements to be removed at each iteration so as to pick the one that minimises the information loss at the end. Therefore, its time complexity overhead falls under the percentage of samplings which is a parameter specified by the user. Given a single input image, the pseudocode for pymils is described in Algorithm 4.
A user can tune pymils (Algorithm 4) with the following parameters:
  1. 1.
    $size(image)_{min}$: Float $0-1$. Minimum final image size ratio (cannot be smaller than 4 rows or columns).
  2. 2.
    Sampling: Float $0-1$. Percent of rows/columns to randomly sample at each step.
  3. 3.
    Chunk size $c$: Integer. Size of neighbouring row/column group to remove at each step.
Algorithm 4 was implemented in a Python package called pymils. This package is freely available and can be downloaded through GitHub. To test the validity of pymils, we used several different images that were converted into binary numpy arrays101010numpy is a Python package that efficiently stores and handles matrices and arrays..
To show the effectiveness of MILS as implemented through pymils, we first test its ability to compress images from the DataSaurus Dozen [48]. Figure 13 shows the 13 images as binary arrays. Each of these images have the same summary statistics, including the x/y mean, the x/y standard deviation, and linear correlation coefficient to two decimal points. This set of different images has been used as a cautionary tale against trusting summary statistics alone without looking directly at the underlying dataset.
Figure 13: The Datasaurus Dozen datasets [48] visualised as binary arrays.
Figure 14 shows the initial complexity as approximated by BDM and the subsequent change in complexity after pymils was implemented on each image. From Figure 14, it is clear that the different images have different algorithmic complexity values, in contrast with summary statistics. In addition, pymils is able to compress these images such that the resulting image is close to the same complexity as the original image.
Figure 14: Scatterplots showing the initial algorithmic complexity as approximated by BDM (top) and the change in complexity after pymils compressed each image to a 50% the original image size.
We tried various other binary array images, including shapes and 1-dimensional elementary cellular automata (ECA) patterns with random initial states. Figure 15 shows an image of a face before and after image compression via pymils. In this example, pymils removed most of the redundancy from the lower half of the original image while preserving complex features in the top half of the original image.
Figure 15: Example image of a face pasted above blank space before (top) and after (bottom) pymils, with a 50% size reduction.
To test the ability of pymils to compress an image while preserving algorithmic complexity, we compared it to two other methods: randomly removing rows/columns and PCA. We randomly removed chunks of rows or columns of the same size from images without measuring the resulting complexity values. The PCA analysis was done with two components (for binary values) on each image and then mapping the resulting values onto an image of the same resulting size of pymils (50% the original size). Results from this comparison are shown in Figure 16. As a result, pymils performs much better on average at preserving algorithmic complexity ($\Delta BDM$) during image compression. Both random deletion and PCA can introduce additional noise in the image, which results in a negative $\Delta BDM$ value.
Figure 16: Boxplots comparing pymils with random row/column deletion and PCA over various test images (top) and Elementary Cellular Automata (ECA) images (bottom). The fact that for BDM the average is closer to zero and the variance (box size) is shorter and more compact, means that the procedure is succeeding at preserving the (algorithmic) information content of the object after dimension reduction, while other methods destroy the information content and, therefore, the underlying program complexity (i.e., its causal structure).
In addition to testing whole images, the initial and resulting BDM values were tracked for the four image quadrants in each image. Results from Figure 17 suggests pymils performs relatively similar across each image quadrant.
Figure 17: Boxplots comparing pymils with random row/column deletion and PCA over various test image quadrants (top) and ECA image quadrants (bottom). In the case of the general images, quadrant averages after reduction had little effect, but for ECAs and their linear evolution, quadrants have different impact depending on the complexity of the region hence also sensitive to different features.
In future research, other variations of Algorithm 2, like Algorithm 4 may be proposed for example in order to improve the sheer cubic dependence on the network size. In this direction, such variations can stem from improvements on the embedded sorting algorithm and its usage within the loop in [Algorithm 2, Step 13]. At the expense of increasing the running time, another future variation may employ a breadth-first search or depth-first search across the edges listed in increasing order of information contribution to the parent graph. This would allow the consideration over a larger quantity of $h$-hop-length rooted paths.
B.9 MILS and other dimensionality reduction methods
Other dimension reduction methods such as UMAP, AE, VAE, t-SNE, PHATE, and others (for surveys see [21, 35, 31]) are, in a deep fundamental sense, special cases of MILS. This is because MILS implements a universal algorithm that is ultimately a generalisation of all other computable methods based on probability distributions, algebra or topology. For example, t-SNE and UMAP, are heavily based on building probability distributions thus they will suffer a similar fate to those reported in [66]. In any event, the ergodicity of the algorithmic complexity approximation implemented by the coding theorem method (CTM) over the software space implies that the BDM will eventually encompass all this space if enough computational resources are expended. In other words, even the most sophisticated algorithms implementing algebraic or topological feature selection beyond those implementing classical probability approaches will be eventually instantiated by CTM (if the algebraic or topological features are recursively enumerable) and therefore MILS can be seen as an asymptotic generalisation of all other algorithms. The question is simply which one may extract certain features of interest faster and which one will be less biased and more universal in the sense that features are not pre-selected (e.g. algebraic or topological) like when picking a method that will maximise but also bias towards some properties of the data in an embedding. MILS, in that sense, is the most universal and agnostic but will not necessarily overtake all others in all applications at the same rate or efficiency, which would only be the case with unbounded computational resources.
Appendix C Theoretical results
C.1 Algorithmic information oscillations from local perturbations
The main idea of the proof of Theorem C.1 is to construct an algorithm for naming the exact edge subset $F\subseteq\mathbb{E}(G)$ that are being deleted (or inserted). Then, we show that, once this enumeration is calculated, this general algorithm employs (or reversibly employs) the perturbation $\mathfrak{p}_{F}^{-}$ (see [13] for a formalisation of algorithmic perturbations) in order to generate $G\setminus F$ from $G$ (or a perturbation $\mathfrak{p}_{F}^{+}$ for $G$ from $G\setminus F$).
Note that the proof is made for any possible subset $F\subseteq\left|\mathbb{E}_{c}(G)\right|$ of perturbed edges, where $\left|F\right|=r\leq\left|\mathbb{E}_{c}(G)\right|$. Therefore, if one takes $\left|F\right|=1$, Theorem C.1 immediately applies to single edge deletions (or insertions) as in Definiton B.12. In addition, the reader is invited to note that the proof of Theorem C.1 holds for the classical case of graphs, but it can be extended to directed graphs.
Theorem C.1.
Let $G=(V,E)$ be an arbitrary simple graph with $E\subseteq\mathbb{E}_{c}(G)$ and $\left|V\right|=N\in\mathbb{N}$. Let $F\subseteq\mathbb{E}_{c}(G)$ be an arbitrary subset of edges, where $\left|F\right|=r\geq 1$. Then,
$\left|I(G,F)\right|\leq 2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(N)\right))\text{ ,}$
if edges in $F\neq\emptyset$, where $F\subseteq E(G)$, are being deleted, or
$\left|I^{-1}(G,F)\right|\leq 2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(N)\right))\text{ ,}$
if edges in $F\neq\emptyset$, where $F\subseteq\mathbb{E}_{c}(G)\setminus E(G)$, are being inserted.
Proof.
We will split the proof in two cases. First, we will prove
$C(G)\leq C(G\setminus F)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(N)\right))$ (29)
and then
$C(G\setminus F)\leq C(G)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(N)\right))\text{ .}$ (30)
First, let $F\subseteq\mathbb{E}_{c}(G)$ be an arbitrary subset of edges. Let $p_{1}\in\{0,1\}^{*}$ be a string that represents the algorithm running on a prefix universal Turing machine $\mathbf{U}$ such that $p_{1}$ receives $a\in\mathbf{L^{\prime}_{U}}$, $b\in\{0,1\}$, and $k,r\in\mathbb{N}$ as inputs. Then, $p_{1}$ calculates $\mathbf{U}\left(a\right)$ and check whether or not $\mathbf{U}\left(a\right)=A_{G^{\prime}}$ for some graph $G^{\prime}$, according to an arbitrarily chosen recursively bijective encoding of an arbitrary adjacency matrix into a bit string. Then, $p_{1}$ execute the following procedure:
  • If $\mathbf{U}\left(a\right)=A_{G^{\prime}}$, then:
    • if $b=0$ (i.e., edges were deleted), then:
      $p_{1}$ enumerates the subsets of size $r$ in the power set $\mathcal{P}(\overline{E}(G^{\prime}))$, which comprises all the subsets of the set
      $\overline{E}(G^{\prime})=\mathbb{E}_{c}(G^{\prime})\setminus E(G^{\prime})$
      of absent undirected edges (i.e., the symmetric zeros) in $A_{G^{\prime}}$. Then, $p_{1}$ replaces in $A_{G^{\prime}}$ every 00 with $1$ that corresponds to the elements in the $k$-th subset $X_{k}\subseteq\overline{E}(G^{\prime})$ in this enumeration of subsets, where $\left|X_{k}\right|=r$. Finally, $p_{1}$ returns this resulting adjacency matrix of graph $G^{\prime}\cup X_{k}$.
    • if $b=1$ (i.e., edges were inserted), then:
      $p_{1}$ enumerates the subsets of size $r$ in the power set $\mathcal{P}(E(G^{\prime}))$, which comprises all the subsets of the set
      $E(G^{\prime})\subseteq\mathbb{E}_{c}(G^{\prime})$
      of present undirected edges (i.e., the symmetric $1$’s) in $A_{G^{\prime}}$. Then, $p_{1}$ replaces in $A_{G^{\prime}}$ every $1$ with 00 that corresponds to the elements in the $k$-th subset $X_{k}\subseteq E(G^{\prime})$ in this enumeration of subsets. Finally, $p_{1}$ returns this resulting adjacency matrix of graph $G^{\prime}\setminus X_{k}$.
  • Otherwise, $p_{1}$ starts any arbitrarily chosen non-halting computation.
Note that, since $\left(G\setminus F\right)\cup F=G$ and $\left(G\cup F\right)\setminus F=G$ and $C(p_{1})$ only varies by $\mathbf{O}(1)$ on choosing between deletions and insertions, we can study only the edge-deletion case without loss of generality.
Also note that one can always define a recursively bijective encoding of an arbitrary adjacency matrix into a bit string. Therefore, if $a=(A_{G^{\prime}})^{*}$ is the shortest program that generates the adjacency matrix $A_{G^{\prime}}$ of an arbitrary graph $G^{\prime}$, then program $p_{1}$ always halts in the cases:
  • $b=0$, $k\leq\binom{\left|\overline{E}(G^{\prime})\right|}{r}$, and $1\leq r\leq\left|\overline{E}(G^{\prime})\right|$ or;
  • $b=1$, $k\leq\binom{\left|E(G^{\prime})\right|}{r}$, and $1\leq r\leq\left|E(G^{\prime})\right|$.
Proof of the upper bound (29):.
We have that there are only
$\left|\overline{E}(G\setminus F)\right|$ $=\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G\setminus F)\right|=$ (31)
$=\left|\mathbb{E}_{c}(G)\right|-\left|E(G\setminus F)\right|=$
$=\frac{N^{2}-N}{2}-\left|E(G)\right|+\left|F\right|=$
$\leq 2\left(\frac{N^{2}-N}{2}\right)=\mathbf{O}\left(N^{2}\right)$
possible absent edges in $G\setminus F$. We know that the exact collection $F$ of edges that were deleted in $G$ can be univocally encoded in less than or equal to
[…]
bits. Note that
$\binom{\left|\overline{E}(G\setminus F)\right|}{r}=\frac{\left|\overline{E}(G\setminus F)\right|!}{r!(\left|\overline{E}(G\setminus F)\right|-r)!}<2^{\left|\overline{E}(G\setminus F)\right|}\text{ .}$ (33)
In addition, we have from Equation (31) that, if $r=\left|F\right|$, then
$\frac{\left|\overline{E}(G\setminus F)\right|!}{r!(\left|\overline{E}(G\setminus F)\right|-r)!}=\\ \frac{\left(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|+\left|F\right|\right)!}{\left|F\right|!(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|+\left|F\right|-\left|F\right|)!}=\\ =\frac{\left(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|+\left|F\right|\right)!}{\left|F\right|!(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|)!}=\\ =\frac{\left(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|+1\right)\cdots\left(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|+\left|F\right|\right)}{\left|F\right|\cdots 1}\leq\\ \leq\left(\left|\mathbb{E}_{c}(G)\right|-\left|E(G)\right|+\left|F\right|\right)^{\left|F\right|}\leq\\ \leq\left(2\left(\frac{N^{2}-N}{2}\right)\right)^{\left|F\right|}\leq\left(N^{2}\right)^{\left|F\right|}\text{ .}$ (34)
Now, let $X_{k}$ corresponds to the exact subset $F$ of edges, $a=(A_{G\setminus F})^{*}$, $b=0$, and $r=\left|F\right|$. Note that, by supposition and construction, $\left|F\right|\leq\left|\overline{E}(G\setminus F)\right|$ and $\left|F\right|\geq 1$. Therefore, by the minimality of the prefix algorithmic complexity [43, 34, 26], the construction of program $p_{1}$, and Equations (32), (33), and (34) we will have that
$C(G)\leq l\left(\left<a,b,k,r,p_{1}\right>\right)\leq$ (35)
$\leq C(G\setminus F)+\log_{2}\left(\binom{\left|\overline{E}(G\setminus F)\right|}{r}\right)+\\ \mathbf{O}\left(\log_{2}\left(\log_{2}\left(\binom{\left|\overline{E}(G\setminus F)\right|}{r}\right)\right)\right)+\\ \log_{2}\left(\left|F\right|\right)+\mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))+\mathbf{O}(1)\leq$ (36)
$\leq C(G\setminus F)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(2\left|F\right|\log_{2}\left(N\right)\right)\right)+\log_{2}\left(\left|F\right|\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))+\mathbf{O}(1)\leq$ (37)
$\leq C(G\setminus F)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\mathbf{O}(\log_{2}\left(\log_{2}(N)\right))$ (38)
Proof of the lower bound (30):.
We have that there are only
$\left|E(G)\right|\leq\frac{N^{2}-N}{2}=\mathbf{O}\left(N^{2}\right)$ (39)
possible present edges in $G$. We know that the exact collection $F$ of edges that were deleted in $G$ (or inserted in $G\setminus F$) can be univocally encoded in
$\log_{2}\left(\binom{\left|E(G)\right|}{r}\right)+\mathbf{O}\left(\log_{2}\left(\log_{2}\left(\binom{\left|E(G)\right|}{r}\right)\right)\right)$ (40)
bits. From Equation (39), we also analogously have as in Equation (31) that, if $r=\left|F\right|$, then
$\binom{\left|E(G)\right|}{r}$ $=\frac{\left|E(G)\right|!}{\left|F\right|!(\left|E(G)\right|-\left|F\right|)!}$ $=$ (41)
$=\frac{\left(\left|E(G)\right|-\left|F\right|+1\right)\cdots\left(\left|E(G)\right|\right)}{\left|F\right|\cdots 1}\leq$
$\leq\left(\left|E(G)\right|\right)^{\left|F\right|}\leq$
$\leq\left(\frac{N^{2}-N}{2}\right)^{\left|F\right|}\leq\left(N^{2}\right)^{\left|F\right|}\text{ .}$
Now, let $X_{k}$ corresponds to the exact subset $F$ of edges, $a=(A_{G})^{*}$, $b=1$, and $r=\left|F\right|$. Note that, by supposition, $\left|F\right|\leq\left|E(G)\right|$ and $\left|F\right|\geq 1$. Therefore, by the minimality of the prefix algorithmic complexity [43, 34, 26], the construction of program $p_{1}$, and Equations (40) and (41), we will have that
$C(G\setminus F)\leq l\left(\left<a,b,k,r,p_{1}\right>\right)\leq$ (42)
$\leq C(G)+\log_{2}\left(\binom{\left|E(G)\right|}{r}\right)+\mathbf{O}\left(\log_{2}\left(\log_{2}\left(\binom{\left|E(G)\right|}{r}\right)\right)\right)+\\ \log_{2}\left(\left|F\right|\right)+\mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))+\mathbf{O}(1)\leq$ (43)
$\leq C(G)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(2\left|F\right|\log_{2}\left(N\right)\right)\right)+\log_{2}\left(\left|F\right|\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))+\mathbf{O}(1)\leq$
$\leq C(G)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\mathbf{O}(\log_{2}\left(\log_{2}(N)\right))$
If one wants tighter lower and upper bounds for multiple edge deletions as a function of the edge density (i.e., the size of $\left|E(G)\right|$ in comparison to $\frac{N^{2}-N}{2}$), then one can re-write the statement of Theorem C.1 and achieve:
$-\left|F\right|\log_{2}\left(\left|E(G)\right|\right)-\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)-\\ \mathbf{O}\left(\log_{2}\left(\log_{2}\left(\left|E(G)\right|\right)\right)\right)-\mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))-\mathbf{O}(1)\\ \leq C(G)-C(G\setminus F)\leq\\ \left|F\right|\log_{2}\left(\frac{N^{2}-N}{2}-\left|E(G)\right|+\left|F\right|\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\\ +\mathbf{O}\left(\log_{2}\left(\log_{2}\left(\frac{N^{2}-N}{2}-\left|E(G)\right|+\left|F\right|\right)\right)\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))+\mathbf{O}(1)$ (44)
Note that, if the number of perturbations are about the same order of the present edges in $G$, then Equation (44) would tell us that the algorithmic information loss is allowed to be logarithmically stronger than the algorithmic information gain, should $\left|E(G)\right|\ll\frac{N^{2}-N}{2}$ holds.
As one might intuitively expect, when the frequency of edge occurrences is nearly equidistributed (in particular, when the graph is incompressible), we can also retrieve from algorithmic information theory [26, 43, 34, 24] some of the particular cases in which the loss and gain of algorithmic information are both dominated by a slightly more-than-linear factor of the number of destructive perturbations:
Corollary C.2.
Let $G=(V,E)$ be an arbitrary $\mathbf{O}(1)$-K-random simple graph (as in Definition B.11) with $E\subseteq\mathbb{E}_{c}(G)$ and $\left|V\right|=N$, where $N\in\mathbb{N}$ is large enough such that
$\left|\left|E(G)\right|-\frac{N^{2}-N}{4}\right|\leq\mathbf{O}\left(\sqrt{\frac{N^{2}-N}{2}\log_{2}\left(\frac{N^{2}-N}{2}\right)}\right)$ (45)
holds. Let $F\subseteq\mathbb{E}_{c}(G)$ be an arbitrary subset of edges, where $\left|F\right|=r\geq 1$. Then,
[…]
$C(G)-C(G\setminus F)\leq$
[…]
if edges in $F\neq\emptyset$, where $F\subseteq E(G)$, are being deleted.
Proof.
From [9], as stated in Section B.6, we have that
$C(E(G))=C(G)\pm\mathbf{O}(1)=C(x)\pm\mathbf{O}(1)\text{ ,}$ (47)
where $x$ is the characteristic string of $G$. Therefore, in accordance with Definitions B.8 and B.11, we will have that
$C(x)\geq l(x)-\mathbf{O}(1)$
In this case, from the Borel normality of the string $x$ [43, 24], we know there is a large enough characteristic string of length $\frac{{N_{0}}^{2}-N_{0}}{2}$ such that the condition in Equation (45) is always satisfied for every $N\geq N_{0}\in\mathbb{N}$. Now, remember Equation (31). Therefore, to achieve the desired proof, just replace $\left|E(G)\right|$ with
$\frac{N^{2}-N}{4}\pm\mathbf{O}\left(\sqrt{\frac{N^{2}-N}{2}\log_{2}\left(\frac{N^{2}-N}{2}\right)}\right)$
into Equation (44). ∎
One can modify Theorem C.1 (or Equation 44) and find an alternative form that does not depend on $|F|$ and edge density (which in fact corresponds to the worst-case information difference in a straightforward encoding of the set $F$ in $\sim\log_{2}\left(2^{\frac{N^{2}-N}{2}}\right)$ bits, thereby simpler than the encoding applied in Theorem C.1):
Corollary C.3.
Let $G=(V,E)$ be an arbitrary simple graph (as in Definition B.3) with $E\subseteq\mathbb{E}_{c}(G)$ and $\left|V\right|=N\in\mathbb{N}$. Let $F\subseteq\mathbb{E}_{c}(G)$ be an arbitrary subset of edges, where $\left|F\right|=r\geq 1$. Then,
$\left|I(G,F)\right|=\\ \mathbf{O}(N^{2})\text{ ,}$
if edges in $F\neq\emptyset$, where $F\subseteq E(G)$, are being deleted, or
$\left|I^{-1}(G,F)\right|=\mathbf{O}(N^{2})\text{ ,}$
if edges in $F\neq\emptyset$, where $F\subseteq\mathbb{E}_{c}(G)\setminus E(G)$, are being inserted.
Proof.
First, remember that
$\binom{n}{k}\leq 2^{n}$ (48)
for very $n\in\mathbb{N}$ and $0\leq k\leq n$. As described in the beginnig of the proof of Theorem C.1, we can focus on the edge deletion case without loss of generality, so that the edge insertion case derives in an analogous inverse manner. This way, we have by definition that
$\left|F\right|\leq\left|E(G)\right|\leq\frac{N^{2}-N}{2}$ (49)
and
$\left|\overline{E}(G\setminus F)\right|\leq\frac{N^{2}-N}{2}\text{ .}$ (50)
Therefore, the proof follows analogously from Theorem C.1, except for subsequentially replacing Equations (32) and (40) with
$\left|\overline{E}(G\setminus F)\right|+\mathbf{O}\left(\log_{2}\left(\left|\overline{E}(G\setminus F)\right|\right)\right)$ (51)
and
$\left|E(G)\right|+\mathbf{O}\left(\log_{2}\left(\left|E(G)\right|\right)\right)\text{ ,}$ (52)
respectively in the proof of Theorem C.1. ∎
Moreover, for single edge perturbations (i.e., $\left|F\right|=1$), these upper and lower bounds become really tight in general, as expected:
Corollary C.4.
Let $G=(V,E)$ be an arbitrary simple graph with $E\subseteq\mathbb{E}_{c}(G)$ and $\left|V\right|=N\in\mathbb{N}$. Then,
$\left|I(G,e_{i})\right|=\left|C(G)-C(G\setminus e_{i})\right|=\mathbf{O}(\log_{2}(N))\text{ ,}$
if an edge is being deleted, or
$\left|I^{-1}(G,e_{i})\right|=\left|C(G)-C(G\cup e_{i})\right|=\mathbf{O}(\log_{2}(N))\text{ ,}$
if an edge is being inserted.
Proof.
First, note that in this particular case we have $r=|F|=1$. Therefore, the proof follows directly from Theorem C.1, where
$2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\mathbf{O}(\log_{2}\left(\log_{2}(N)\right))\leq\\ 2\log_{2}\left(N\right)+\mathbf{O}\left(1\right)+\mathbf{O}(\log_{2}\left(\log_{2}(N)\right))\text{ .}$
C.2 Eventual optimal values for graph complexity from edge deletions
The semi-computability of the algorithmic complexity is a well-known phenomenon in algorithmic information theory and theoretical computer science [24, 26, 30, 43]: one can approximate the value of $C(x)$ from above with computable enumerations, but there is no computable function that return the value of $C(x)$ in general for every $x$. In other words, one equivalently cannot have an algorithm that in general can calculate how much the error between the approximate value to algorithmic complexity and its actual value is decreasing over time (as the enumerations of halting programs unfolds in the approximation from above). This contrasts with the result we will present in Theorem C.8, where we prove the existence of a single edge that carries maximum information about the whole network (i.e., it is maximally redundant with respect to the information content of the network) whose actual size of algorithmic information content (i.e., the exact algorithmic complexity value) can be realistically approximated.
Furthermore, in this section we will explore the possibility of taking advantage of a single-edge perturbation on $G$ with purpose of calculating the value of $C(G)$. In other words, although no algorithm can determine in general the exact edge deletions that produce the desired phenomenon, we will present some particular cases in which an edge deletion works like an oracle that carries the sufficient amount of information for calculating the value of $C(G)$. Thus, algorithms for approximations of some arbitrary measures defined upon algorithmic information content, such as MILS algorithms, can eventually harvest this information in order to achieve these optimal values, should enough computational resources be expended. See Theorem C.9. Contributing to theoretical foundations of the methods employed in the present article, these eventually optimal calculations, although unpredictable in general, highlight the empirical significance of general methods based on algorithmic complexity, such as BDM, when compared with other lossless compression methods that do not (at least in a computable enumeration manner) asymptotically cover the space of partial recursive functions for compressing data.
We will study a case in which the characteristic string (i.e., a bit string that univocally determines the presence or absence of the edges [11, 12]) of a compressible and dense enough graph already contains useful information for identifying where the edge deletion occurred, so that the deleted edge’s positioning becomes information itself.
So, the aim is to prove a particular case and, thus, show that the phenomenon exists and may occur more often. To achieve Theorem C.9 with this purpose, the main idea of its proof is to construct some particular cases of simple graphs $G$ whose each characteristic string’s prefix—in the form $x\upharpoonright_{k}$ (i.e., a $k$-bit-length initial segment of $x$)—, respective to each $G$, carries information sufficient for computing the entire characteristic string $x$, so that this prefix is logarithmically much smaller than the characteristic string (i.e., $k=\mathbf{o}\left(\log_{2}\left(l(x)\right)\right)$), where $l\left(x\right)$ denotes the length of the string $x$. Moreover, these graphs should be sufficiently dense. Thus, there will be a program that, given the perturbed graph as input, can always decide where a perturbation landed onto the characteristic string, should the edge deletion occur in the characteristic string’s suffix. Hence, one just need to choose the actual index $i$ of the edge $e_{i}$ destroyed (i.e., the bit in the characteristic string in which $1$ was replaced with 00) that has
$C\left(i\right)\leq\mathbf{o}\left(\log_{2}\left(\left|V(G)\right|\right)\right)$
bits of algorithmic complexity. In addition, the information in the binary encoding of $i$ carries all the necessary information of the minimum generating program of $G$, so that
$C\left(G^{*}|i\right)=\mathbf{O}\left(1\right)$
and
$C\left(i|\left(G\setminus e_{i}\right)\right)=\mathbf{O}\left(1\right)\text{ .}$
In this manner, we will get that one becomes able to compute $G^{*}$ from $G\setminus e_{i}$, and vice versa.
Before reaching the desired theorem, we introduce the concept of prefix self-determined graphs. Let $F_{G}$ be an infinite family of simple graphs. Let $p\in\{0,1\}^{*}$ only depend on the choice of $F_{G}$. By prefix self-determined, we mean that there is a program $p$ that can return the string $x$ given only a proper prefix of the characteristic string $x$ of $G$ as input. Formally, as in Definition C.1, we say a graph $G\in F_{G}$ is prefix self-determined iff there is $y=x\upharpoonright_{n}$ with $n<l(x)$ such that $\mathbf{U}\left(\left<y,p\right>\right)=x$, where $x$ is the characteristic string of $G$. In this article, we are applying the notion of prefix self-determination to graphs through its characteristic strings, but prefix self-determination is a property that can be clearly assigned to strings (or any other type of encodable object) in general.
Definition C.1.
Let $F_{G}$ be an infinite family of simple graphs. Let $p\in\{0,1\}^{*}$ only depend on the choice of $F_{G}$. We say a graph $G\in F_{G}$ is prefix self-determined iff there is $y=x\upharpoonright_{n}$ with $n<l(x)$ such that $\mathbf{U}\left(\left<y,p\right>\right)=x$, where $x$ is the characteristic string of $G$.
Thus, prefix self-determination is a stronger notion in comparison to the widely employed self-delimitation [24, 26] from proper prefixes in algorithmic information theory. From the prefix of a prefix self-determined string, one might not only get the length of the string, but also promptly compute the very string. In other words, a prefix self-determined string contains a compression of itself. It is immediate to see that, for an infinite family of prefix self-determined strings, if $y$ is the respective proper prefix of string $s$, then
$C(s)\leq C(y)+\mathbf{O}(1)\leq l(y)+\mathbf{O}(1)\leq l(s)+\mathbf{O}(1)\text{ .}$ (53)
As a natural extension, one may want to limit a maximum length for the prefix with respect to the string. This is the case of prefix self-determined graphs whose initial segment $y=x\upharpoonright_{n}$ of its characteristic string has length $n=\mathbf{O}\left(\log_{2}\left(l(x)\right)\right)$. Let $F_{G}$ be an infinite family of simple graphs. As in Definition C.2, we say a graph $G\in F_{G}$ is logarithmically prefix self-determined iff $G$ is prefix self-determined and $n=\mathbf{O}\left(\log_{2}\left(l(x)\right)\right)$. Thus, a logarithmically prefix self-determined graph has a very small subset of edges (which are encoded into the initial segment of the characteristic string) that algorithmically determine the whole graph. This property grasps the idea of a small subgraph carrying all the information sufficient for computably determining the graph, should the initial segment of the characteristic string be in turn a characteristic string of a vertex-induced subgraph.
Definition C.2.
Let $F_{G}$ be an infinite family of simple graphs. We say a graph $G\in F_{G}$ is logarithmically prefix self-determined iff $G$ is prefix self-determined and $n=\mathbf{O}\left(\log_{2}\left(l(x)\right)\right)$.
Remember the usual notion of a family of dense graphs being defined as an infinite family $F_{G}^{{}^{\prime}}$ of graphs in which every member $G\in F_{G}^{{}^{\prime}}$ has $\left|E(G)\right|=\mathbf{\Theta}\left(\left|V(G)\right|^{2}\right)$.
In order to prove that Definition C.2 and, as a direct consequence, Definition C.1 are not vacuous (i.e, that they are satisfiable), we can choose a particular type of self-delimiting program that returns itself concatenated with repetitions of itself (in addition, we choose a prefix that is also nearly incompressible):
Lemma C.5.
There is an infinite family $F_{G}$ of logarithmically prefix self-determined dense graphs whose respective proper prefixes $v\in\{0,1\}^{*}$ follows
$\left|C(v)-l(v)\right|=\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)$
and
$l(v)=\mathbf{o}\left(\log_{2}\left(l(x)\right)\right)\text{ ,}$
where $x$ is the characteristic string of $G$.
Proof.
This proof is based on constructing characteristic strings of the form $w\;r\upharpoonright_{k}w^{\prime}$, where $r\in\left[0,1\right]\subseteq\mathbb{R}$ is an arbitrary $\mathbf{O}(1)$-K-random (i.e., prefix algorithmically random as in Definition B.9) real number111111 For example, one may choose $r$ to be the halting probability $\Omega$ without loss of generality in this proof., and $w,r\upharpoonright_{k},w^{\prime}\in\{0,1\}^{*}$. Moreover, we want that
$l(w)+k=\mathbf{o}\left(\log_{2}\left(l(w\;r\upharpoonright_{k}w^{\prime})\right)\right)=\mathbf{O}\left(\log_{2}\left(l(w\;r\upharpoonright_{k}w^{\prime})\right)\right)\text{ ,}$
$\left|E(G)\right|=\mathbf{\Theta}\left(\left|V(G)\right|^{2}\right)\text{ ,}$
and that there is a program $p_{2}$ such that
$\mathbf{U}\left(\left<w\,r\upharpoonright_{k},p_{2}\right>\right)=w\;r\upharpoonright_{k}w^{\prime}\text{ .}$
To this end, first let $p_{2}\in\{0,1\}^{*}$ be a string that represents the algorithm running on a prefix universal Turing machine $\mathbf{U}$ that receives $x\in\{0,1\}^{*}$ as input and:
  1. (I)
    first, decides whether or not there are $w\in\mathbf{L^{\prime}_{U}}$ and $y\in\{0,1\}^{*}$ such that $x=wy$, where $l(w)<l(x)$ ;
  2. (II)
    then,
    1. (a)
      if $x=wy$, and this holds where $l(w)$ is the minimum length of a string $w$ for which the above Step (I) holds:
      1. i.
        build the string
        $z=y\,y\,y\cdots\,\left(y\upharpoonright_{h}\right)$
        by concatenating the string $y$ until finding $h\in\mathbb{N}$, where $h\leq l(y)$, for which
        $l(wz)=\frac{n^{2}-n}{2}\text{ ,}$
        where
        $n=\min\left\{t\middle|\begin{array}[]{l}t\in\mathbb{N}\\ \frac{t^{2}-t}{2}\geq 2^{{l(y)}^{2}}+l(w)\end{array}\right\}\text{ ;}$
      2. ii.
        return $wz$.
    2. (b)
      Otherwise, return “$error$”.
Note that $p_{2}$ always halts for every input and we immediately have that either $\mathbf{U}\left(\left<x,p_{2}\right>\right)=wz$ or $\mathbf{U}\left(\left<x,p_{2}\right>\right)=\text{ ``$error$'' }$ ”. Since $w$ and $y$ were arbitrary in the construction of $p_{2}$, let $w$ encode the length of $y$ in language $\mathbf{L^{\prime}_{U}}$. Hence, we will have that
$l(wy)$ $=l(w)+l(y)\leq$ (54)
$\leq\mathbf{O}\left(\log_{2}\left(l(y)\right)\right)+\,l(y)\leq$
$\leq\mathbf{o}\left(l(y)\right)+\,l(y)=$
$=\mathbf{o}\left({l(y)}^{2}\right)$
$=\mathbf{o}\left(\log_{2}\left(l(w)+2^{{l(y)}^{2}}\right)\right)=$
$=\mathbf{o}\left(\log_{2}\left(l(wz)\right)\right)\text{ .}$
From [9, 11, 12], we know the family of all possible simple graphs can be recursively labeled with a unique way of ordering the sequence of all possible edges. Thus, since by construction there is $n\in\mathbb{N}$ such that $l(wz)=\frac{n^{2}-n}{2}$, let $G$ be a simple graph whose characteristic string is $wz=\mathbf{U}\left(\left<wy,p_{2}\right>\right)$. This way, we have that
$\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}=l(wz)\geq 2^{{l(y)}^{2}}+l(w)$ (55)
and, by the construction of $p_{2}$ and the minimality of $|V(G)|$,
$\mathbf{O}\left(2^{{l(y)}^{2}}+l(w)\right)+2^{{l(y)}^{2}}+l(w)\geq\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}=l(wz)$ (56)
Until here, we proved the logarithmically prefix self-determination part and proved that
$l(wy)=\mathbf{o}\left(\log_{2}\left(l(wz)\right)\right)\text{ .}$
Now, let $r\in\left[0,1\right]\subseteq\mathbb{R}$ be an arbitrary $\mathbf{O}(1)$-K-random real number. From the Borel normality of $r$ (as in the proof of Corollary C.2), we know there is a large enough $k\in\mathbb{N}$ such that the number #$\#_{1}(r\upharpoonright_{k})$ of occurrences of $1$’s in $r\upharpoonright_{k}$ follows
|#$\left|\#_{1}(r\upharpoonright_{k})-\frac{k}{2}\right|=\mathbf{O}\left(\sqrt{k\log_{2}\left(k\right)}\right)\text{ .}$ (57)
Since $w$ and $y$ were arbitrary in the construction of $p_{2}$, let $y\coloneqq r\upharpoonright_{k}$.121212 Note that, for any large enough length $l(wz)$, the string $wz$ cannot ever be $\mathbf{O}(1)$-K-random, since it is based on multiple repetitions of the same string. Therefore, since $k>k_{0}$ can be arbitrarily large given a large enough $k_{0}$, we will obtain an infinite family $F_{G}$ of distinct graphs with characteristic strings of the form $wz$. Remember that, by construction, $w$ encodes the integer value $k$ and that $k>k_{0}$ can be arbitrarily large given that large enough $k_{0}$. Therefore, from Equations (55), (56), and (57), by construction of $p_{2}$ and since $l(w)=\mathbf{O}\left(\log_{2}\left(k\right)\right)$, we will have that
$\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}&\geq\left|E(G)\right|\\ &\geq\frac{2^{k^{2}}-k}{2}-\left(\frac{2^{k^{2}}-k}{k}\right)\,\mathbf{O}\left(\sqrt{k\log_{2}\left(k\right)}\right)\\ &=\mathbf{\Omega}\left(2^{k^{2}}+l(w)\right)\\ &=\mathbf{\Omega}\left(\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}\right)\text{ .}\end{aligned}$ (58)
Notice that Equation (58) holds because
$\frac{2^{k^{2}}-k}{2}-\left(\frac{2^{k^{2}}-k}{k}\right)\,\mathbf{O}\left(\sqrt{k\log_{2}\left(k\right)}\right)\gg\frac{2^{k^{2}}-k}{2}-\frac{2^{k^{2}}-k}{4}$ (59)
and
$\frac{2^{k^{2}}-k}{2}>\frac{2^{k^{2}}+l(w)}{4}$ (60)
hold for sufficiently large $k$. Therefore, Equation (58) proves $\left|E(G)\right|=\mathbf{\Theta}\left(\left|V(G)\right|^{2}\right)$. Now, the fact that $C(r\upharpoonright_{k})=C(y)\geq k-\mathbf{O}(1)$ holds follows directly from Definition B.9. In addition, from Lemma B.1, we have that
$C\left(wy\right)\leq l\left(r\upharpoonright_{k}\right)+\mathbf{O}\left(\log_{2}\left(k\right)\right)+\mathbf{O}\left(\log_{2}\left(\log_{2}\left(k\right)\right)\right)\\ \leq k+\mathbf{O}\left(\log_{2}\left(k\right)\right)$ (61)
and
$C(y)\leq C(wy)+\mathbf{O}\left(\log_{2}\left(\log_{2}\left(k\right)\right)\right)$ (62)
Therefore, by making $v\coloneqq wy$, where $l(y)=k$ and $l(w)=\mathbf{O}\left(\log_{2}\left(k\right)\right)$, and achieving
$k-\mathbf{O}\left(\log_{2}\left(\log_{2}\left(k\right)\right)\right)\leq C(v)\leq k+\mathbf{O}\left(\log_{2}\left(k\right)\right)$ (63)
from the incompressibility of $r$ and Equations (61) and (62), we will have it that
$l(v)-\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)-\mathbf{O}\left(\log_{2}\left(\log_{2}\left(l(v)\right)\right)\right)\leq C(v)\leq\\ l(v)+\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)$ (64)
Now we show in Lemma C.6 that the compression of the characteristic string and its density are so strong that enables one to pick and delete the right edge in a way such that this very deletion encodes the optimal compression of $G$:
Lemma C.6.
There is an infinite family $F_{G}$ of logarithmically prefix self-determined dense simple graphs (in particular, any family satisfying Lemma C.5) such that, for every large enough member $G\in F_{G}$, there is at least one edge deletion so that
$C\left(G^{*}|\left(G\setminus e_{i}\right)\right)=\mathbf{O}(1)\text{ .}$
Proof.
Let $F_{G}$ be an arbitrary infinite family of graphs satisfying Lemma C.5. Thus, there is $v=x\upharpoonright_{n}$ with $n<l(x)$ such that $\mathbf{U}\left(\left<v,p_{2}\right>\right)=x$, $n=\mathbf{o}\left(\log_{2}\left(l(x)\right)\right)$, and $\left|C(v)-l(v)\right|=\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)$, where $x$ is the characteristic string of $G$. Let $p_{3}\in\{0,1\}^{*}$ be a string that represents the algorithm running on a prefix universal Turing machine $\mathbf{U}$ that receives the adjacency matrix $A_{G^{\prime}}$ and $p_{2}$ as inputs and:
  1. 1.
    applies any previous arbitrarily chosen method of recursively encoding a matrix into a string and check whether or not $A_{G^{\prime}}$ is a matrix;
  2. 2.
    if $A_{G^{\prime}}$ is a matrix, then:
    1. (a)
      reversively applies the previous arbitrarily chosen method of recursively encoding a matrix into a string to produce the correspondent characteristic string $x^{\prime}$;
    2. (b)
      decides whether or not there are $w^{\prime}\in\mathbf{L^{\prime}_{U}}$ and $z^{\prime}\in\{0,1\}^{*}$ such that $x^{\prime}=w^{\prime}\,z^{\prime}$, where $l(w^{\prime})<l(x^{\prime})$ and $w^{\prime}$ encodes the binary representation of a natural number $h\in\mathbb{N}$ in language $\mathbf{L^{\prime}_{U}}$;
    3. (c)
      if $x^{\prime}=w^{\prime}\,z^{\prime}$ holds such that $l(w^{\prime})$ is the minimum prefix for which this holds, then:
      1. i.
        makes $v\coloneqq w^{\prime}\,y^{\prime}$, where $y^{\prime}$ is the string of the next $\mathbf{U}\left(w^{\prime}\right)$ bits in $x^{\prime}$ right after the end of $w^{\prime}$;
      2. ii.
        makes $x\coloneqq\mathbf{U}\left(\left<v,p_{2}\right>\right)$  ;
      3. iii.
        decides whether or not $x^{\prime}=x$ ;
      4. iv.
        if $x^{\prime}<x$ and $l(x)=l(x^{\prime})$, then:
        1. A.
          enumerates all the $1$’s in $x$;
        2. B.
          search for the first bit of $x$ that displays value $1$ while the same bit of $x^{\prime}$ displays value 00;131313 Notice that this step evinces one of the advantages of using the characteristic string formalism from [9, 11, 12].
        3. C.
          calculates de index $i\in\mathbb{N}$, where the $i$-th $1$ in $x$ corresponds to this first bit in which the values differ from each other;
        4. D.
          returns the binary representation $\left(i\right)_{2}\in\{0,1\}^{*}$ of the integer $i\in\mathbb{N}$.
      5. v.
        otherwise, returns “$error$”.
    4. (d)
      otherwise, returns “$error$”.
  3. 3.
    otherwise, returns “$error$”.
Note that, if $\mathbf{U}$ halts on $v,p_{2}\right>$, then $p_{3}$ always halts. We will have it that by hypothesis, the minimality of algorithmic complexity, and Lemmas B.1 and B.2,
$C(x)$ $\leq C(v)+\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)+\mathbf{O}(1)\leq$ (65)
$\leq l(v)+\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)+\mathbf{O}(1)=$
$=\mathbf{o}\left(\log_{2}\left(l(x)\right)\right)$
holds. This occurs because $\mathbf{U}\left(\left<v,p_{2}\right>\right)=x$ and $l(v)=\mathbf{o}\left(\log_{2}\left(l(x)\right)\right)$ hold. Therefore, once a method of recursively encoding a matrix into a string to produce the correspondent string $x^{\prime}$ is fixed, we will have that, from Lemmas B.1 and B.2 and the fact that $x$ is the characteristic string of $G$,
$l(G^{*})=C(G)\leq\mathbf{o}\left(\log_{2}\left(\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}\right)\right)+\mathbf{O}(1)$ (66)
holds. Therefore, since there are only up to
$2^{\mathbf{o}\left(\log_{2}\left(\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}\right)\right)+\mathbf{O}(1)}$ (67)
possible bit strings of length $\leq l(G^{*})$ and $G$ is dense, then there is at least one present edge that is the $i$-th bit of $x$, where $i>l(v)$, and $i\in\mathbb{N}$ corresponds to the program $G^{*}\in\mathbf{L^{\prime}_{U}}$. Then, $\mathbf{U}\left(\left<A_{(G\setminus e_{i})},p_{2},p_{3}\right>\right)=G^{*}$ and, since $p_{2}$ and $p_{3}$ are fixed, $C\left(G^{*}|(G\setminus e_{i})\right)=\mathbf{O}(1)$. ∎
Notice that the procedure in the proof of Lemma C.6 was designed in such a way that it can be recursively inverted so as to make the perturbed graph be computably retrievable from $G^{*}$ too:
Lemma C.7.
There is an infinite family $F_{G}$ of logarithmically prefix self-determined dense simple graphs (in particular, any family satisfying Lemma C.5) such that, for every large enough member $G\in F_{G}$, there is at least one edge $e_{i}$ that satisfies Lemma C.6 such that
$C\left(\left(G\setminus e_{i}\right)|G^{*}\right)=\mathbf{O}(1)$
also holds.
Proof.
The idea of the proof is to construct a program that, with entry $G^{*}$ as input, performs the inverted operation that program $p_{3}$ from the proof of Lemma C.6 does. Then, it builds the correspondent adjacency matrix of the perturbed graph, that is, the adjacency matrix of $A_{(G\setminus e_{i})}$ that satisfies Lemma C.6. To this end, let $p^{\prime}_{3}\in\{0,1\}^{*}$ be a string that represents the algorithm running on a prefix universal Turing machine $\mathbf{U}$ that receives the $\left(i\right)_{2}$ as input and:
  1. 1.
    makes $A_{G}\coloneqq\mathbf{U}\left(\left(i\right)_{2}\right)$  ;
  2. 2.
    builds the characteristic string $x$ of $A_{G}$;
  3. 3.
    replace only the $i$-th $1$ in the characteristic string $x$ with 00 and builds the resulting string $x^{\prime}$, where $x^{\prime}<x$;
  4. 4.
    returns $A_{G^{\prime}}$ for which $x^{\prime}$ is its characteristic string.
If $i$ corresponds to the program $G^{*}$ in the proof of Lemma C.6, then we will have that $p^{\prime}_{3}$ halts and $A_{G^{\prime}}=A_{(G\setminus e_{i})}$. Therefore, we will have that, from the minimality of algorithmic complexity,
$C\left(\left(G\setminus e_{i}\right)|G^{*}\right)=\mathbf{O}(1)$ (68)
holds. ∎
Now remember that a neutral information edge $e$ is one such that the information difference is minimal so that it contributes less to the whole information content of $G$. By combining the previous two Lemmas C.6 and C.7 one can easily obtain the following result:
Theorem C.8 (Maximally neutral information edge).
There is an infinite family $F_{G}$ of logarithmically prefix self-determined dense simple graphs (in particular, any family satisfying Lemma C.5) such that, for every large enough member $G\in F_{G}$, there is at least one (maximally) neutral information edge $e_{i}\in E(G)$ such that
$\left|I(G,e_{i})\right|=\left|C(G)-C(G\backslash e_{i})\right|=\mathbf{O}(1)$ (69)
holds.
Proof.
The proof follows directly from Lemmas C.6 and C.7 and Definition B.12 by employing the basic properties of AIT in Lemmas B.1 and B.2. ∎
Identifying complex systems that displays such a phenomenon, as the graphs in Theorem C.8, may reveal those systems that are capable of carrying the sufficient information about where the perturbation occur and, hence, it enables future investigation on how this correlates with robustness or self-repair, two concepts which may be also related to the preservation of network topological properties studied in the present article.
Moreover, one can take advantage of Lemma C.5 in order to obtain one of the particular cases in which a computer can already find nearly optimal values of algorithmic complexity in polynomial computation time (i.e., running time) even with brute force, as shown in Theorem C.9.
The proof of Theorem C.9 is based on the fact that, although program $p_{2}$ demands an exponential computation time with input $wy$, the computation time of emulating program $wy,p_{2}\right>$ is still polynomial with respect to $\left|E(G)\right|$, since $G$ is dense from Lemma C.5. This way, from the “almost” incompressibility of the prefix $wy$ also demonstrated in Lemma C.5, we will achieve the desired upper bound for the error with respect to the optimal value $C(G)$.
Theorem C.9.
Let $G$ be a large enough simple graph that belongs to an infinite family $F_{G}$ of logarithmically prefix self-determined dense simple graphs. Then, there is a polynomially time-bounded exhaustive search algorithm that, with input $G\in F_{G}$, can approximate the integer value $C(G)$ with an error bounded by $\mathbf{o}\left(\log_{2}\left(\log_{2}\left(\left|V(G)\right|\right)\right)\right)$ bits.
Proof.
First, we choose without loss of generality a self-delimiting universal programming language $\mathbf{L^{\prime}_{U}}$ that is recognizable in polynomial time as in Section B.6. Let $\mathbf{P}$ be the usual time complexity class of polynomially time-bounded deterministic Turing machines [18, 52]. As a direct consequence, then we will have that Steps (I) or (II)b in the proof of Lemma C.5 can be solved in $\mathbf{P}$. Note that Step (II)(a)ii is basically the concatenation of the string $w$ obtained from Step (I) with the string resulting from Step (II)(a)i. Thus, since we know the concatenation operation [50] and Step (I) are in $\mathbf{P}$, we can just focus on analyzing the time complexity of Step (II)(a)i, which is indeed where the computation is most demanding. In addition, from the minimality of $n$, construction of $p_{2}$, and Equation (56), we have that at most
$\frac{\mathbf{O}\left(2^{{l(y)}^{2}}+l(w)\right)}{l(y)}$
concatenations will be needed to find the minimum $n$ with
$l(wz)=\frac{n^{2}-n}{2}\geq 2^{{l(y)}^{2}}+l(w)\text{ .}$
Now, given $k\geq 4$ tapes with one of the tapes filled with the binary representation of this maximum number of concatenations, one of the tapes filled with $w$, and another tape filled with $y$ (which leaves at least one of the remaining tapes empty or blank), we will have that there is a $k$-tape Turing machine that returns $wz$ in at most
$\mathbf{O}\left(\left(\frac{2^{{l(y)}^{2}}+l(w)}{l(y)}\right)\,\left(\max\left\{l(y),l(w)\right\}\right)\right)$ (70)
computation steps. This holds because the tape in which one informs the number of concatenations (needed to calculate $n$) can be filled exactly with $\frac{\mathbf{O}\left(2^{{l(y)}^{2}}+l(w)\right)}{l(y)}$ as many contiguously repeated $1$’s; and this tape works as a counter to the number of concatenations the $k$-tape Turing machine must perform in the remaining empty tape in order to halt when this tape is filled with $wz$. Now, note that the worst-case time complexity for another machine to construct the counter string (i.e., the one constituted by $1$’s with length $\frac{\mathbf{O}\left(2^{{l(y)}^{2}}+l(w)\right)}{l(y)}$) from input $wy$ is upper bounded by an exponential term $2^{\mathbf{O}\left({l(y)}^{2}\right)}$. Therefore, since the worst-case time complexity of simulating any $k$-tape Turing machine is also in $\mathbf{P}$ [18, 52], then the time complexity of $p_{2}$ is upper bounded by an exponential term as a function of $l(wy)$ (i.e., $p_{2}$ represents an algorithm in $\mathbf{EXPTIME}$). In fact, since $wz$ is the output of $p_{2}$, we will necessarily also have a best-case time complexity (for one-way single-tape Turing machines) of
$2^{{l(y)}^{2}}+l(w)-l(wy)$
by just comparing the lengths of the output and input. Thus, since $l(wy)=\mathbf{o}\left(\log_{2}\left(l(wz)\right)\right)$, $p_{2}$ is also exponential in the best case scenario, which thereby implies that the worst-case time complexity of $p_{2}$ is indeed exponential. Remember by hypothesis that $G$ is dense, i.e., $\left|E(G)\right|=\mathbf{\Theta}\left(\left|V(G)\right|^{2}\right)$. Then, we will have that $\left|E(G)\right|=\mathbf{\Theta}\left(l(wz)\right)$. Now, note that there are only up to $2^{l(\left<wy,p_{2}\right>)}$ possible bit strings with length at most $l(\left<wy,p_{2}\right>)$. Since the worst-case time complexity of simulating any Turing machine in a universal Turing machine is polynomially bounded with respect to the original computation time of the machine being simulated [18, 52], then we have that exhaustively searching all possible outputs of exponentially time-bounded programs with length $\leq l(\left<wy,p_{2}\right>)$ has a worst-case computation time bounded by
[…]
for some large enough constant $C\in\mathbb{N}$. Remember our choice of the pairing function in Notation B.5, so that, since $p_{2}$ is fixed, we will have that
$l\left(\left<wy,p_{2}\right>\right)\leq l(wy)+l(p_{2})+\mathbf{O}\left(\log_{2}\left(l(wy)+l(p_{2})\right)\right)=\\ \mathbf{O}\left(l(wy)\right)\text{ .}$
Therefore, since $l(wy)=\mathbf{o}\left(\log_{2}\left(wz\right)\right)$ by hypothesis, Equation (71) holds, and as we have shown, $p_{2}$ represents an algorithm in $\mathbf{EXPTIME}$ given $wy$ as input, we will have that there is an exhaustive search algorithm with $G$ as input that will eventually test $wy,p_{2}\right>$ as the generating program of $G$ in worst-case computation time upper bounded by
$\mathbf{o}\left(\left(l\left(wz\right)\right)^{c^{\prime}}\right)\leq\mathbf{O}\left(\left|E(G)\right|^{c^{\prime\prime}}\right)\text{ ,}$ (72)
for some large enough constants $c^{\prime},c^{\prime\prime}\in\mathbb{N}$. Therefore, such an algorithm returns a value $C^{\prime}(G)\leq l\left(\left<wy,p_{2}\right>\right)$ in polynomial computation time with input $G$. Notice that, from Lemma C.5, we have that $\left|C(wy)-l(wy)\right|=\mathbf{O}\left(\log_{2}\left(l(wy)\right)\right)$, $G$ is dense, and $l(wy)=\mathbf{o}\left(\log_{2}\left(l(wz)\right)\right)$ hold, and $w$ encodes the length of $y$. Then, from Lemma C.5 and the basic properties of algorithmic information theory in Section  B.6, we will have it that
$C(G)$ $\leq C^{\prime}(G)+\mathbf{O}(1)$ (73)
$\leq l(wy)+\mathbf{O}(1)$
$\leq l(y)+\mathbf{O}\left(\log_{2}\left(l(y)\right)\right)$
$\leq l(y)+\mathbf{o}\left(\log_{2}\left(\log_{2}\left(l(wz)\right)\right)\right)\text{ .}$
We also obtain that
$C^{\prime}(G)+\mathbf{O}(1)$ $\geq C(y)$ (74)
$\geq l(y)-\mathbf{O}(1)$
$\geq C(y)-\mathbf{O}\left(\log_{2}\left(l(y)\right)\right)-\mathbf{O}(1)$
$\geq C(y)-\mathbf{o}\left(\log_{2}\left(\log_{2}\left(l(wz)\right)\right)\right)-\mathbf{O}(1)\text{ .}$
Finally, from Equations (73) and (74) and the fact that $l(wz)=\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}$, we will have it that
$\left|C^{\prime}(G)-C(G)\right|=\mathbf{o}\left(\log_{2}\left(\log_{2}\left(\left|V(G)\right|\right)\right)\right)$
holds, which finishes the proof of Theorem C.9. ∎
Thus, Theorem C.9 shows that there is an algorithm that, with $G$ as input, calculates a value $C^{\prime}(G)$ that could be considered as a good approximation of the algorithmic complexity of $G$ in terms of compression rate. Although the actual $C(G)$ and the approximation $C^{\prime}(G)$ may differ by a number of bits of about a logarithmic term of its own value $C(G)$ (or of the order of a double logarithmic term of the input size), they are both asymptotically comparable with respect to $\left|V(G)\right|$. This phenomenon is expected to have an even faster speedup for algorithms that approximate algorithmic complexity values, such as BDM that is based on the decomposition into blocks of objects whose complexity is approximated using the Coding Theorem Method (CTM) [32, 56, 55, 62], which in turn precomputes lookup tables so as to exchange computation time for space/memory [68]. In this sense, BDM can reach optimality for even larger families of graphs beyond that of Theorem C.9. This is because one can easily extend the latter in order to demonstrate that BDM not only can find the nearly optimal approximations for logarithmically prefix self-determined graphs, but also can find those for graphs that are self-determined by logarithmically size-bounded $2$D- or $3$D-partitions of most (if not all) of its constituent subgraphs. Additionally, the precomputation of the lookup table enables BDM to take more advantage of larger quantities of constituent blocks in comparison to other brute-force methods of approximating algorithmic complexity from scratch every time. We suggest that this theoretical finding might be one of the reasons MILS is performing better than other size reduction methods. Future research should be pursued with the purpose of revealing other intrinsic structural characteristics underlying MILS and other data size reduction techniques, particularly with respect to modularity [72, 38, 67].
The following corollary follows directly from the combination of Theorems C.8 and C.9.
Corollary C.10.
Let $G$ be a large enough simple graph that belongs to an infinite family $F_{G}$ of logarithmically prefix self-determined dense simple graphs. Then, there is a polynomially time-bounded exhaustive search algorithm such that for every large enough member $G\in F_{G}$, there is at least one edge $e_{i}\in E(G)$ such that this algorithm can approximate both the integer values $C(G)$ and $C(G\setminus e_{i})$ with an error bounded by $\mathbf{o}\left(\log_{2}\left(\log_{2}\left(\left|V(G)\right|\right)\right)\right)$ bits.
Notice that Theorem C.9 and Corollary C.10 contribute to the theoretical foundations of the methods employed in the main article by showing that these eventually optimal calculations, although unpredictable in general, highlight the empirical significance of general methods based on algorithmic complexity, such as BDM, when compared with other lossless compression methods that do not (at least in a computable enumeration manner) asymptotically cover the space of partial recursive functions for compressing data.
C.3 Time Complexity of Minimal Information Loss Selection (MILS)
For MILS to be well-defined, we need to guarantee that the deletion of an element (e.g., a vertex or edge in a graph) produces a unique object in a deterministic fashion. In the case of edges in networks, the problem is when there are elements $e_{i},e_{j}\in E$ such that $I(G,e_{i})=I(G,e_{j})$ and the algorithm cannot uniquely decide whether to remove $e_{i}$ or $e_{j}$ first, which potentially (and likely) leads to the production of different objects. Algorithm 2 avoids this problem and shows that the procedure is robust. Note that this algorithm and Theorem C.11 are stated for graphs, but they can be generalized for multidimensional objects $\mathbb{U}$ in general.
Theorem C.11.
Let $T_{ID}(G)$ be the worst-case computation time of calculating the value $I(G,e_{i})$ for any $e_{i}\in E\left(G^{\prime}\right)$ by the chosen algorithmic-complexity approximating algorithm, where $G^{\prime}$ is any edge-induced subgraph of $G$141414 For example, by choosing the BDM, $T_{ID}(G)$ is the worst-case computation time of calculating the value of $K_{log}(G^{\prime})-K_{log}(G^{\prime}\setminus e_{i})$ for every $e_{i}\in E\left(G^{\prime}\right)$, where $G^{\prime}$ is any edge-induced subgraph of $G$. See Definition B.12 and Equation (16). Then, MILS (Algorithm 2) is a deterministic algorithm that has a cubic worst-case time complexity with input $E(G)$ and linear worst-case time complexity in $T_{ID}(G)$.
Proof.
Let $e_{i}$ and $e_{j}$ be any two edges of a graph $G$ (whose size is to be reduced by the application of MILS) with $I(G,e_{i})=I(G,e_{j})$, then MILS will remove $e_{i}$ and $e_{j}$ from $G$ simultaneously. This element deletion condition helps MILS to always produce the same output for a given input, which makes it a deterministic algorithm. To prove this claim, note that $I(G,e_{i})$ calculated by any algorithmic-complexity approximating algorithm (e.g., by the BDM) always produce an integer, ordering any set of integers is decidable and always produce unique outputs for the same set of integers. Therefore, for every iteration of the loop in [Algorithm 2, Step 13], there will be a unique size-reduced graph $G_{n-i-1}$ from a previous graph $G_{n-i}$ by deleting the edges in $\{e_{i}\in E(G_{n-i}):I(G_{n-i},e_{i})=\textit{minLoss}\}$. In fact, MILS can only vary the output for the same input if more computation time is expended on the CTM’s precomputation before the BDM. In order to study the time complexity of MILS, note that the worst-case of function $\text{MILS}(G,N)$ is given when $N=1$ and InfoRank(G) outputs a list of $\left|E(G)\right|$ distinct entries. This way, since one has $\left|E(G)\right|$ iterations in the loop in [Algorithm 2, Step 4] by insertion (thus, an operation in $\mathbf{O}(1)$) of the result of the approximation of $I(G,e_{i})$ in the array informationLoss, which has size $\left|E(G)\right|$, and sorting algorithms have quadratic worst-case time complexity, we will have a worst-case time complexity of function InfoRank(G) in
$\mathbf{O}\left(\left|E(G)\right|\left(T_{ID}(G)\right)\right)+\mathbf{O}\left(\left|E(G)\right|^{2}\right)\text{ .}$ (75)
In addition, since one has up to $\left|E(G)\right|$ iterations in the loop, finding minimums in the first dimension of arrays takes a linear worst-case computation time, and deleting elements in an array also takes a linear worst-case computation time, we will have a worst-case time complexity of function $\text{MILS}(G,N)$ in
$\mathbf{O}\left(\left|E(G)\right|\,\left(T_{IR}(G)+\left|E(G)\right|\right)\right)\text{ ,}$ (76)
where $T_{IR}(G)$ is the worst-case computation time of function InfoRank(G) for any $G$ (which was already studied in Equation (75) so that
$T_{IR}(G)=\mathbf{O}\left(\left|E(G)\right|\left(T_{ID}(G)\right)+\left|E(G)\right|^{2}\right)$ (77)
holds). Now, suppose without loss of generality that the inputs of Algorithm 2 are $E(G)$ and $N$. Therefore, since composing functions is closed under polynomial computation time as in [Algorithm 2, Step 15] and function InfoRank(G) is called recursively, we will have that the worst-case time complexity of Algorithm 2 is
$\mathbf{O}\left(\left|E(G)\right|\,\left(\left|E(G)\right|\left(T_{ID}(G)\right)+\left|E(G)\right|^{2}+\left|E(G)\right|\right)\right)=$ (78)
$\mathbf{O}\left(\left|E(G)\right|^{3}+\left|E(G)\right|^{2}\left(T_{ID}(G)\right)\right)$
Algorithm 2 also produces a speedup with respect to Algorithm 3 when all elements have the same information value, and thus are to be deleted at the same time: while in Algorithm 3 there would be a time complexity overhead to perform this task, in Algorithm 2, given a previously calculated InfoRank(G), it will be required just one step in the loop and, therefore, just a linear computation time in $\mathbf{O}\left(\left|\textit{informationLoss}\right|\right)$, where informationLoss is the output of InfoRank(G). For example, any attempt to reduce the dimension of the complete graph (either by e.g. single-vertex or single-edge deletion) will produce an empty graph.
Thus, MILS’ Algorithm 2 scales in a tractable manner with respect to the complexity approximating method for very large networks. This is because the worst-case running time is basically dominated by the amount of computational resources expended in the sole approximation to the values of algorithmic complexity as this amount only contributes linearly to the overall time complexity when the network size increases.

### v12

@@
 Abstract
-We present a novel, domain-agnostic, model-independent, unsupervised, and universally applicable approach for data summarization.
-Specifically, we focus on addressing the challenge of reducing certain dimensionality aspects, such as the number of edges in a network, while retaining essential features of interest.
-These features include preserving crucial network properties like degree distribution, clustering coefficient, edge betweenness, and degree and eigenvector centralities.
-Our approach outperforms state-of-the-art network reduction techniques by achieving an average improvement in feature preservation.
-Previous methods grounded in statistics or classical information theory have been limited in their ability to capture more intricate patterns and features, particularly nonlinear patterns stemming from deterministic computable processes.
-Moreover, these approaches heavily rely on a priori feature selection, demanding constant supervision.
-Our findings demonstrate the effectiveness of the algorithms proposed in this study in overcoming these limitations, all while maintaining a time-efficient computational profile.
-In many instances, our approach not only matches but also surpasses the performance of established network reduction algorithms.
-Furthermore, we extend the applicability of our method to lossy compression tasks involving images or any bi-dimensional data.
-This highlights the versatility and broad utility of our approach in various domains.
-Keywords: data summarisation; lossy compression; data dimensionality reduction; network complexity; image segmentation; algorithmic machine learning
+We present a novel, domain-agnostic, model-independent, unsupervised, and universally applicable Machine Learning approach for dimensionality reduction based on the principles of algorithmic complexity.
+Specifically, but without loss of generality, we focus on addressing the challenge of reducing certain dimensionality aspects, such as the number of edges in a network, while retaining essential features of interest.
+These features include preserving crucial network properties like degree distribution, clustering coefficient, edge betweenness, and degree and eigenvector centralities but can also go beyond edges to nodes and weights for network pruning and trimming.
+Our approach outperforms classical statistical Machine Learning techniques and state-of-the-art dimensionality reduction algorithms by preserving a greater number of data features that statistical algorithms would miss, particularly nonlinear patterns stemming from deterministic recursive processes that may look statistically random but are not.
+Moreover, previous approaches heavily rely on a priori feature selection, which requires constant supervision.
+Our findings demonstrate the effectiveness of the algorithms in overcoming some of these limitations while maintaining a time-efficient computational profile.
+Our approach not only matches, but also exceeds, the performance of established and state-of-the-art dimensionality reduction algorithms.
+We extend the applicability of our method to lossy compression tasks involving images and any multi-dimensional data.
+This highlights the versatility and broad utility of the approach in multiple domains.
+keywords:
+Machine Learning , lossy algorithmic complexity, recursive compression, data dimensionality reduction, network complexity, algorithmic image segmentation
+journal: Information Sciences
+[…]
+[KCL]organization=Department of Biomedical Computing, Deparment of Digital Twins, School of Biomedical Engineering and Imaging Sciences, addressline=King’s Institute for AI, King’s College London, country=U.K.
+[…]
+[Turing]organization=The Alan Turing Institute, addressline=British Library, country=U.K.
+[…]
+[OxfImAl]organization=Oxford Immune Algorithmics, Oxford University Innovation and London Institute for Healthcare Engineering, country=U.K.
+[…]
+[ADL]organization=Algorithmic Dynamics Lab and Department of Oncology-Pathology, Center of Molecular Medicine, addressline=Karolinska Institutet, country=Sweden
+[…]
+[Cross]organization=Cross Labs, addressline=Cross Compass, country=Japan
+[…]
+[OncPat]organization=Karolinska Institutet, addressline=Department of Oncology-Pathology, Center of Molecular Medicine, city=Stockholm, country=Sweden
+[…]
+[WisMad]organization=University of Wisconsin-Madison, addressline=John W. and Jeanne M. Rowe Center for Research in Virology, Morgridge Institute for Research, country=USA
+[…]
+[Unicamp]organization=University of Campinas (UNICAMP), addressline=Centre for Logic, Epistemology and the History of Science (CLE), country=Brazil
+[…]
+[LNCC]organization=National Laboratory for Scientific Computing (LNCC), addressline=Data Extreme Lab, country=Brazil
+[…]
+[Vene1]organization=Universidad Central de Venezuela, addressline=Instituto Nacional de Bioingeniería, country=Venezuela
+[…]
+[Vene2]organization=Universidad Central de Venezuela, addressline=Department of Mathematics, country=Venezuela
+[…]
+[KAUST]organization=King Abdullah University of Science and Technology (KAUST), addressline=Biological and Environmental, Science and Engineering Division, country=Kingdom of Saudi Arabia
 1 Introduction
-The study of large and complex datasets, or Big Data, organised as networks, has emerged as one of the central challenges in most areas of science and technology, cellular and molecular networks in biology being one of the prime examples.
+The study of large and complex datasets such as networks, has emerged as one of the central challenges in most areas of complex systems and Machine Learning, cellular and molecular networks in biology being one of the prime examples.
 The analysis of large-scale networks poses significant computational and analytical challenges.
-Hence, network sparsification methods have emerged as indispensable tools to address these challenges by reducing the complexity of networks while preserving essential structural or functional properties [14, 19, 21, 27, 31, 35, 46].
-Moreover, it’s worth noting that data reduction techniques extend beyond network data.
-Methods for image sparsification also play pivotal roles in managing and analysing large datasets.
+Network sparsification methods have become indispensable tools[1, 2] for addressing these challenges by reducing the complexity of networks[3, 4] while preserving essential structural [5] or functional properties [6, 7].
+Data reduction techniques extend beyond network data.
+Methods for image sparsification also play an essential role in managing and analysing large datasets.
 In essence, data reduction, in a broader sense, involves transforming various types of numerical, textual, or visual digital information into more compact representations while retaining specific properties of significance.
 This overarching concept of data reduction encompasses a wide range of methods, including network sparsification, image compression, and simplifying cellular automata rules.
-The common thread among these methods is the pursuit of more efficient data representation, enabling streamlined analysis and storage across diverse fields.
-The primary objective of data dimensionality reduction is twofold, much like two sides of the same coin.
-One facet emphasises minimising information loss, ensuring that valuable insights and nuances are preserved as data is transformed into a more manageable form.
-The other aspect underscores the importance of preserving the most ’meaningful’ features that uniquely characterise an object, a concept commonly referred to as feature selection.
+The common thread among these methods is the pursuit of more efficient data representation, enabling streamlined analysis and storage across various fields.
+The primary objective of data dimensionality reduction in Machine Learning is twofold, much like two sides of the same coin.
+One facet emphasizes minimizing information loss, ensuring that valuable insights and nuances are preserved as data is transformed into a more manageable form.
+The other aspect underscores the importance of preserving the most ‘meaningful’ features that uniquely characterise an object, a concept commonly referred to as feature selection.
 This dual objective serves as a fundamental principle in data reduction, often constituting a trade-off to be tackled in the data analysis phase.
-Traditionally, identifying these meaningful features relies on an user-centric, subjective criterion.
-Specific analytical goals and human expertise guide selection processes.
+Traditionally, identifying these meaningful features is based on a subjective criterion that is user-centric.
+Specific analytical goals and human experience guide the selection processes.
 For example, linear algebraic (e.g., matrix analysis) and statistically-based dimensionality reduction techniques attempt to minimise statistical information loss under certain algebraic (interpreted as signal and noise) conditions, consequently maximizing the statistical mutual information between the desired information and the dimensionally-reduced output.
 The primary objective of dimensionality reduction in network analysis is to approximate a complex network with a sparser representation.
-Numerous methods for graph sparsification, a form of graph summarization, have been documented in the literature [46].
+Numerous methods for graph sparsification, a form of graph summarisation, have been documented in the literature [7].
 Some earlier approaches, such as Chew’s [4], relied on criteria like the shortest path distance between pairs of vertices for network sparsification.
 Others, like Benczur and Karger [8], utilized cut problems as a basis for sparsification.
-More recently, techniques such as spectral similarity of graph Laplacians have been employed by Spielman et al.
+Techniques such as spectral similarity of graph Laplacians have been employed by Spielman et al.
 [9].
@@
 Consequently, data analysts often face the daunting task of making arbitrary choices regarding which features to prioritize, as exemplified in previous research [11, 12].
-This paper’s main goal is to develop an unsupervised method that remains theoretically optimal and is independent of computational models, programming languages, physical implementations, and available resources.
-Furthermore, this method should be agnostic to selecting features of interest, encompassing all computably enumerable parts within a given dataset earmarked for reduction.
+This paper’s main goal is to develop an unsupervised method that is theoretically sound and optimal in distinguishing noise from data, efficient, and independent from traditional statistical Machine Learning approaches based on classical information theory.
+Furthermore, the methods introduced are agnostic in selecting features of interest, encompassing all estimated computably enumerable parts within a given dataset earmarked for reduction.
 We introduce a family of semi-computable algorithms designed to preserve computable properties, encompassing both statistical and algorithmic aspects.
-This approach is a generalization of various dimension reduction techniques since statistically definable, generable, or recognizable patterns constitute a subset of those that algorithmic approaches can theoretically handle.
+This approach is a generalisation of various dimension reduction techniques since statistically definable, generable, or recognizable patterns constitute a subset of those that algorithmic approaches can theoretically handle.
 Our method represents a novel approach to designing theoretically optimal lossy compression techniques, drawing inspiration from principles and estimations related to theoretical optimal lossless compression [14, 13].
-By employing efficient (polynomial) approximations to algorithmic complexity values from recent numerical methods of algorithmic probability [32, 56, 68, 55, 62, 69], we demonstrate how these algorithms can preserve structure properties, performing similarly to, if not consistently outperforming, state-of-the-art algorithms in e.g.
-the area of network dimension reduction.
-We test our algorithms on non-trivial cases against transitive and non-linear (spectral) methods involving simple graphs where statistical regularities are even easier to conceal more intricate computable processes or embedded encodings, and thus may easily fool weaker, linear and computable measures [66, 65, 11, 12].
+By employing efficient (polynomial) approximations [16, 15] to algorithmic complexity [17] values from recent numerical methods of algorithmic probability, like in the works of  [18] [19] and [20], we demonstrate how these algorithms can preserve structural properties, performing similarly to, if not consistently outperforming, state-of-the-art algorithms in, e.g., the area of network dimension reduction.
+We test our algorithms on non-trivial cases against transitive and non-linear (spectral) methods involving simple graphs where statistical regularities are even easier to conceal more intricate computable processes or embedded encodings, and thus may easily fool weaker [11], linear and computable measures [10, 21].
 We will demonstrate how our method preserves critical network topological properties, including degree distribution, clustering coefficient, edge betweenness, degree, and eigenvector centralities compared to other methods.
-These fundamental properties provide essential insights into network structure and functionality.
+These fundamental properties provide essential information about the structure and functionality of the network.
 Degree distribution, for instance, reveals how nodes are interconnected, helping us understand the overall network architecture and the distribution of connections.
 It is indispensable for identifying hubs and assessing network robustness.
-The clustering coefficient quantifies the tendency of nodes to form tightly connected clusters.
+The clustering coefficient quantifies the tendency of the nodes to form tightly connected clusters.
 This metric is invaluable for detecting network communities and assessing resilience against failures or attacks.
 Edge betweenness measures the significance of edges in facilitating efficient communication between nodes.
-Identifying bottleneck edges is vital for understanding network flow and vulnerability.
-Degree and eigenvector centralities help pinpoint influential nodes within a network.
+Identifying bottleneck edges is vital to understanding network flow and vulnerability.
+The degree and eigenvector centralities help identify influential nodes within a network.
 High-degree nodes are hubs, while high-eigenvector centrality nodes are well-connected to other influential nodes.
-These centralities are vital for identifying key players in various network types.
-By preserving these critical properties, our approach strives to provide more meaningful and accurate representations of complex networks [67, 11, 12] while addressing the challenges inherent in data reduction techniques.
+These centralities are vital for identifying key players in various types of networks.
+By preserving these critical properties, our approach strives to provide more meaningful and accurate representations of complex networks [12, 22] while addressing the challenges inherent in data reduction techniques.
 Central to our approach are perturbations in the form of edge deletions aimed at minimizing the loss of algorithmic information content [23].
@@
 They can be implemented using various approaches, including entropy or lossless compression.
-However, our description employs a technique based on [68, 72, 63], which not only encompasses but also enhances other statistics- and entropic-based methods due to its intrinsic structural characteristics.
-In Section 2.2, where we detail the algorithm, it’s essential to recognise that substituting the underlying methods for approximating algorithmic information content with alternatives like Shannon entropy or lossless compression algorithms represents special cases within the broader algorithm based on algorithmic complexity.
-This comprehensive approach effectively covers all these less powerful cases because the block decomposition method (BDM) calculation we employ subsumes entropy-based approximations [68].
+However, our description employs a technique based on [24, 14], which not only encompasses but also enhances other statistics- and entropic-based methods due to its intrinsic structural characteristics.
+It is essential to notice that substituting the underlying methods to approximate algorithmic information content with alternatives such as Shannon entropy or lossless compression algorithms represents special cases within the broader algorithm based on algorithmic complexity.
+This comprehensive approach effectively covers all these less powerful cases, because the block decomposition method (BDM) calculation we employ subsumes entropy-based approximations [16].
 BDM presents a novel approach to understanding and quantifying the complexity of biological networks and systems.
@@
 This results in efficient estimations, albeit with some loss of accuracy.
-BDM has been applied in various fields and applications outperforming other complexity measures [28, 67], for example to evolutionary systems [38], artificial intelligence [41, 17], and the detection of bio- and technosignature [26, 29].
-Significantly, BDM’s application to graphs and causal calculus has enabled the deconvolution of complex network behaviors, identifying distinct topological generating mechanisms [33, 63, 18, 61].
-This approach uses a perturbation-based causal calculus to infer model representations [13, 60, 40], effectively tackling the challenges of causation in complex systems.
-It has been demonstrated by us and others how by analyzing data from discrete dynamical systems like cellular automata and complex networks, BDM-based algorithms provides a deeper understanding of the intricate interplay of elements within these systems, revealing hidden patterns and relationships, and contributing significantly to the field of algorithmic complexity in biological networks[33, 18, 64, 70, 72, 63].
-Moreover, algorithmic complexity, being an asymptotically optimal measure of irreducible information content [26, 34, 43, 24], holds the unique advantage of capturing algorithmically determined patterns that defy reduction to statistical patterns [28, 66].
+BDM has been applied in various fields and applications that outperform other complexity measures [25, 12], for example, to evolutionary systems [26], artificial intelligence [28, 27], and the detection of bio- and technosignature [29].
+Significantly, the application of BDM to graphs and causal calculus has enabled the deconvolution of complex network behaviors [30], identifying distinct topological generating mechanisms [13].
+This approach uses a perturbation-based causal calculus to infer model representations [31, 32], effectively tackling the challenges of causation in complex systems.
+We and others have demonstrated how by analysing data from discrete dynamical systems such as cellular automata and complex networks, BDM-based algorithms provide a deeper understanding of the intricate interplay of elements within these systems, revealing hidden patterns and relationships, and contributing significantly to the field of algorithmic complexity [33] in biological networks [34, 24].
+Moreover, algorithmic complexity, being an asymptotically optimal measure of irreducible information content [35, 36], has the unique advantage of capturing algorithmically determined patterns that defy reduction to statistical patterns [25].
 Our implementation of the MILS algorithm, rooted in algorithmic complexity, offers the added benefit of avoiding certain low-complexity deceptive phenomena.
 Traditional methods based solely on entropy and statistics are prone to such issues [25, 11].
-Significantly, this avoidance does not compromise running time efficiency or data size reduction performance, as evidenced by our results in Section 4.
+Significantly, this avoidance does not compromise running time efficiency or data size reduction performance for graphs, as evidenced by our results in Section 5.
+On the other hand, when implemented in ML pipelines, computational time still presents challenges.
 These results demonstrate that we either match or, more often than not, surpass the performance of the best current algorithms, both for local and global graph properties.
-The approach presented in this article lies in its potential to assess the effectiveness of all reduction techniques while striving for optimal reduction by minimising algorithmic information loss.
+In image-compression tasks, we show that MILS presents the highest mean classification accuracy per bit of compressed images when such images are used in a particular classification task.
+The approach presented in this article lies in its potential to assess the effectiveness of all reduction techniques while striving for optimal reduction by minimizing algorithmic information loss.
 This represents a nonlinear generalisation encompassing a broader range of techniques beyond merely preserving statistical or domain-specific algebraic properties.
-While our focus in this article centers on edge deletions in graphs to describe the MILS algorithm, it’s essential to emphasise that MILS can be applied to a wide array of objects and various types of constructive or destructive perturbations [13, 60, 72, 70] on their constituent elements.
-For instance, MILS can seamlessly extend its applicability to node deletion, offering versatility in addressing diverse algorithmic perturbations [13] aimed at minimising information differences between objects.
+While our focus in this article centers on edge deletions in graphs to describe the MILS algorithm, it’s essential to emphasise that MILS can be applied to a wide array of objects and various types of constructive or destructive perturbations [24, 34] on their constituent elements.
+For instance, MILS can seamlessly extend its applicability to node deletion, offering versatility in addressing diverse algorithmic perturbations [23] aimed at minimizing information differences between objects.
+2 Preliminaries on complexity measures and algorithmic probability
+A comprehensive discussion of complexity measures and algorithmic probability is provided in the Supplementary Material.
+Meanwhile, this section introduces key preliminaries to establish the foundational concepts necessary for the main discussion.
+Using the principles of classical information theory, Block Decomposition Method (BDM) or BDM combines the calculation of the global Shannon Entropy rate of the object with local estimations to algorithmic complexity of smaller blocks in which the object is decomposed for which values are found in a pre-computed database of direct approximations of algorithmic probability [16].
+The BDM relies on the following assumptions:
+  1. 1.
+    In the case of small enough objects (e.g., binary strings), their algorithmic complexity can be approximated using exhaustive search.
+  2. 2.
+    For larger objects, breaking them into smaller parts allows for the approximation of the overall complexity by summing the complexity of individual blocks, with a correction factor to account for interactions between the blocks.
+  3. 3.
+    For every other length, values of Shannon Entropy rates are calculated and combined with the previous values by using the same principles of information theory.
+Formally, let $x$ be a string divided into blocks $x_{i}$, with $x=x_{1}\oplus x_{2}\oplus\dots\oplus x_{n}$, where $\oplus$ denotes a concatenation operator.
+The BDM complexity of $x$, denoted $\text{BDM}(x)$, is:
+$\text{BDM}(x)=\sum_{i=1}^{n}\text{CTM}(x_{i})+\log n$
+Where:
+  • $\text{CTM}(x_{i})$ is the algorithmic complexity approximation for block $x_{i}$, derived from the Coding Theorem Method (CTM).
+  • $\log n$ is a correction factor accounting for the interactions between the blocks.
+The Coding Theorem Method (CTM) is a method based on the Coding Theorem and Algorithmic Probability [36], which connects classical probability to Kolmogorov complexity.
+The CTM maps sets of micro programs (e.g., small Turing machines) to small assembly objects for which it can empirically estimate the algorithmic probability of an object based on the following relationship [15]:
+$K(s)\approx-\log P(s)$
+Where:
+  • $K(s)$ is the Kolmogorov complexity of string $s$.
+  • $P(s)$ is the algorithmic probability of string $s$, as defined by Solomonoff’s distribution.
+The CTM leverages pre-computed distributions from small Turing machines to approximate the algorithmic complexity of small strings by estimating their algorithmic probability [19, 20].
+BDM and CTM can be seen as practical approximations of algorithmic probability and Kolmogorov complexity, especially in scenarios where exact computation is impossible.
+The relationship can be summarised as follows:
+On the one hand, CTM provides an approximation to algorithmic probability $P(s)$ by connecting the empirical frequency of a string’s generation in a small Turing machine model to its Kolmogorov complexity $K(s)$, using the relation $K(s)\approx-\log P(s)$.
+On the other hand, BDM offers a method to map micro programs to their respective small pieces of an object requiring an algorithmic explanation.
+2.1 Algorithmic information theory, data compression and feature selection
+BDM, CTM and Algorithmic Probability offer a principled approach to data compression and feature selection by quantifying algorithmic complexity rather than relying on traditional statistical methods.
+In data compression, these methods allow for data decomposition into patches and estimating their Kolmogorov complexity, ultimately selectively removing redundant regions while preserving structurally meaningful information.
+This empowers compression routines which prioritize algorithmically significant structures rather than merely frequent ones, resulting in compression that minimises storage requirements while retaining essential data properties.
+In feature selection, these methods provide a more fundamental measure of information content than conventional statistical metrics like correlation or variance.
+These methods help identify which features contain essential computational information, enabling the removal of redundant or low-complexity features without significant loss of structure.
+Thus, their ability to optimize data representation while retaining essential information makes them valuable tools for scientific computing, machine learning, and complex data-driven applications.
 3 Method
 3.1 Calculating information difference values
-All methods employed in this article are founded on the information difference among an object’s elements.
+All methods employed in this article are based on information difference among an object’s elements.
 Specifically, we focus on the contribution of each element to the system’s overall information content, such as the impact of nodes or links in a network context.
 Formally, this involves assessing the impact of local perturbations (changes made to individual elements) on the algorithmic information content of the entire object, compared to its prior state.
-These methods are grounded in algorithmic/causal perturbation analysis, a key component of algorithmic information dynamics (AID) [13, 60, 40].
-In the specific context of this article, our approach entails the perturbation of system elements through the removal of components, followed by measuring and ranking the effects on the system’s algorithmic information content.
-This dynamic approach to studying and calculating the potential changes in an object for evaluating the contribution of its components was initially introduced in [61, 64, 72, 71].
+These methods are grounded in algorithmic/causal perturbation analysis, a key component of algorithmic information dynamics (AID) [31, 32].
+In this contribution, our approach involves the perturbation of the elements of the system through the removal of components, followed by measuring and ranking the effects on the algorithmic information content of the system.
+This dynamic approach to studying and calculating the potential changes in an object to assess the contribution of its components was initially introduced in [38, 37].
 We extend these concepts to data/network dimension reduction and lossy compression of bi-dimensional data, such as images.
@@
 Notice that this difference could be defined for perturbations on the graph’s vertices, but here, we will restrict our attention to perturbations on edges.
-In the present article, as introduced in the next Section 2.2, we focus on the empirical investigation of destructive perturbations (i.e., the edge deletion problem) to perform network reduction while minimising information loss.
+In the present article, as introduced in the next Section 3.2, we focus on the empirical investigation of destructive perturbations (i.e., the edge deletion problem) to perform network reduction while minimizing information loss.
 However, in the Sup.
@@
 Let $G$ be a graph, and $\left|E\left(G\right)\right|$ denote its number of edges.
-MILS seeks to reduce $G$ to a graph on $N<\left|E\left(G\right)\right|$ edges so that the loss of information is minimized.
+MILS seeks to reduce $G$ to a graph on $N<\left|E\left(G\right)\right|$ edges so that the loss of information is minimised.
 Ideally, one would do so by calculating the information differences between $G$ and $G\backslash S$, where $\left|S\right|\geq 1$, and then finding the subset $F\subset E(G)$ such that $I(G,F)\leq I(G,S)$ for all non-empty proper subsets of edges $S\subset E(G)$, repeating this task with $G\coloneqq G\backslash F$ until the target size is reached (i.e., when $|E(G)|=N$).
@@
 performs such a sequential deletions of edges, and then removes the edge that contributes less to the information content of the graph before moving to the next step.
-When $e$ is such that $\left|I(G,e)\right|$ is minimal, where $I(G,e)$ denotes $I(G,\{e\})$ (i.e., the information difference from deleting only a single edge), we call it a neutral information edge [72] because it is an edge that contributes the least to the information content of $G$ (in particular, it minimizes information loss or the introduction of spurious information into the network according to the information difference when removed from the original network).
+When $e$ is such that $\left|I(G,e)\right|$ is minimal, where $I(G,e)$ denotes $I(G,\{e\})$ (i.e., the information difference from deleting only a single edge), we call it a neutral information edge [24] because it is an edge that contributes the least to the information content of $G$ (in particular, it minimises information loss or the introduction of spurious information into the network according to the information difference when removed from the original network).
 The existence of such edges whose information is maximally neutral (or redundant) is demonstrated in Theorem $3.8$ in the Sup.
@@
 2 shows that a decision process of minimisation of information differences implies a maximization of (algorithmic) probabilities both locally for each arrow and globally for each path from the root to a leaf.
-This holds theoretically independent of the choice of the computation model, programming language, or any arbitrary (computable) probability (semi-)measure that one aims at maximising (whether or not one assumes independence between parent vertices and subsequent child events in such a measure).
+This holds theoretically independent of the choice of the computation model, programming language, or any arbitrary (computable) probability (semi-)measure that one aims at maximizing (whether or not one assumes independence between parent vertices and subsequent child events in such a measure).
 See Section $2.8.1$ in the Sup.
@@
 That is, the algorithms return smaller networks whose respective edge count reductions are done so that the information difference between the original network and the smaller one is minimised according to those optimal upper and lower bounds described in the above paragraphs.
-This holds not only for the ideal EXPTIME version of MILS mentioned in Section 2.2, whose iterations cover the power set of the set $E(G)$ of present edges, but our results show that switching to a summarization procedure based on successive single-edge deletions (like in Algorithm 1) is sound.
+This holds not only for the ideal EXPTIME version of MILS mentioned in Section 3.2, whose iterations cover the power set of the set $E(G)$ of present edges, but our results show that switching to a summarisation procedure based on successive single-edge deletions (like in Algorithm 1) is sound.
 In particular, as discussed in this section, the latter case is grounded on the fact we proved that any maximum error does not dependent on the particular ideal subset of edges to be deleted, but only on the sheer size of this subset (once the worst-case logarithmic contribution of the network size becomes fixed).
@@
 Instead of searching along all possible paths, which would demand exponential running time, each iteration corresponds to an one-step increase in the height of the tree from the root, and in each step the algorithm runs the InfoRank procedure in order to select the edge deletions to be made.
+4.1 How and Why MILS works
+The MILS process is a greedy algorithm that begins with patch decomposition, where the input data–an image, matrix, or graph–is divided into smaller regions or patches.
+These patches have their algorithmic complexity estimated using BDM, which quantifies how difficult these patches are to generate as the result of a short-length program (or many of them, according to Algorithmic Probability).
+During the deletion phase, MILS applies perturbations such as edge deletions in graphs or bit-flipping in images while maximizing preservation of the size of the shortest programs that generate the data.
+Instead of relying on conventional downsampling techniques, which often distort structural information, MILS strategically removes elements in a way that minimally affects BDM (its algorithmic complexity estimation), thereby attempting to preserve the original dataset’s causal, recursive or algorithmic structure.
+If after removing an element the complexity of the 2D array does not change, with high probability the removed element is not relevant to the generative process itself and can therefore be safely removed as compared to other more relevant.
+Then coarse-graining is performed by ensuring that the algorithmic complexity of each patch remains nearly constant even after deletion, meaning that the essential features of the data are retained despite the data dimensionality reduction.
+The remaining patches are then stitched together in a way that maintains their original relationships.
+If a region of the dataset is highly compressible and exhibits low BDM, it can tolerate more deletions without significantly affecting its algorithmic information content connected to its causal content.
+Conversely, highly complex regions with high BDM are preserved more carefully, as they contain intricate patterns that are difficult to reconstruct from limited data.
+Ultimately, MILS does not remove data arbitrarily but rather follows an algorithmic-complexity-preservation strategy, selecting deletions that minimise changes in the generative complexity of the data.
+By treating the dataset as a set of algorithmic programs and ensuring that deletions do not substantially alter the shortest program needed for reconstruction, MILS effectively reduces dimensionality while maintaining structural integrity.
+This makes it highly valuable for applications in data compression, robust machine learning, and high-dimensional data analysis, where retaining essential information is critical.
 5 Results
-4.1 A comparison between MILS and other data dimensionality reduction methods
+5.1 Experimental Validation and Benchmarking of MILS for Dimensionality Reduction in Machine Learning Pipelines
+As indicated, MILS method offers a novel approach by leveraging principles of algorithmic complexity to achieve superior compression with minimal degradation of accuracy.
+MILS is particularly relevant in quantisation scenarios where preserving information per bit is crucial, such as low-power edge computing, compressed neural networks, and large-scale retrieval-augmented systems (RAGs).
+By incorporating MILS into ML pipelines, models can maintain higher accuracy per bit of compressed data, making it an attractive alternative to conventional dimension reduction methods.
+To evaluate the efficiency of MILS, we conduct a benchmark study focusing on classification accuracy performance per bit of compressed image.
+This metric provides insight into how well different dimension reduction methods retain information relative to their storage and computational cost.
+The MILS method is compared against several established techniques across different quantisation levels to determine its effectiveness in preserving predictive performance while minimizing information loss.
+We utilize the MNIST dataset from the ‘sklearn‘ library as the benchmark dataset for this study.
+It consists of 1797 8x8 images of digits, encompassing 17 features in each image (17 grayscale tones).
+There are about 180 samples per digit class in this dataset and it is suitable for quick benchmarking due to its relatively small size.
+Given its widespread use in evaluating ML techniques, MNIST serves as an ideal benchmark to test the efficacy of MILS in a controlled setting.
+Three levels of quantisation are considered for our analysis and were obtained after using the ‘gfloat’ Python package:
+  • 1-bit binary representation (used by MILS)
+  • 4-bit floating point representation (OCP E2M1)
+  • 8-bit floating point representation (OCP E4M3)
+  • 16-bit floating point representation (Brain Floating Point bfloat16)
+Besides, MILS is compared against three classes of dimension reduction techniques:
+  • Matrix decomposition-based methods: Principal Component Analysis (PCA) and Non-Negative Matrix Factorisation (NMF).
+  • Random projection-based methods: Sparse Random Projection (SPJ) and Gaussian Random Projection (GRJ).
+  • Spectral methods: Kernel PCA with radial basis function (KPCAR) and polynomial kernels (KPCAP). The link of Kernel PCA with spectral methods was previously established in [40].
+To properly compare compression techniques of images, we need to consider their reconstructed counterparts to assess whether information loss has occurred.
+Thus, rather than using the compressed representations directly, we analyze the reconstructed images to determine if the compression workflow led to a degradation of relevant information.
+Thus, the following Machine Learning pipeline was implemented to perform the benchmarks:
+  • Create an image compression step, where different models were used to compress the original 8x8 images. The output of this step is also a 8x8 image (the one coming from reconstruction).
+  • From the compressed-reconstructed version of the 8x8 image, train a Support Vector Classifier to classify the dataset according to the MNIST classes (digits from 0 to 9).
+Since we will perform model selection, a 5-fold StratifiedShuffleSplit cross-validation (CV) is performed with test size equal to 20% of input samples, ensuring that class balance is retained during training and that samples are shuffled.
+For simplicity, no hyperparameter tuning will be performed, thus a validation dataset is not needed.
+For the CV results assessment, the mean balanced accuracy metric was chosen (the average of recall obtained on each class).
+This Pipeline was repeated for varying compression rates, hereby governed by the number of components used in traditional techniques or the number of edges kept in the MILS workflow.
+We chose the following values for these: 1,2,4,8,16,32 and 64, being the latter equivalent to using the full uncompressed image as input.
+Besides the pyMILS implementation described in the Supplementary Material, we used another implementation which considers Algorithm 2 (where each edge is visited to check for candidate deletion cases).
+This MILS implementation is shown in Listing 1.
+Listing 1: MILS core function
+bdm = BDM(ndim=2,partition=PartitionCorrelated)
+pe = PerturbationExperiment(bdm, metric=’bdm’)
+def MILS(X,N):
+X_red = X.copy()
+idx = np.argwhere(X_red)
+while idx.shape[0]>N:
+pe.set_data(X_red)
+delta_bdm = pe.run(idx)
+b = idx[np.argmax(-np.abs(delta_bdm))]
+X_red[*b] = 0
+idx = np.argwhere(X_red)
+return X_red
+Such implementation relies on the ‘pyBDM’ Python package and its implementation of a Perturbation Experiment (which changes elements in a 2D array and calculates the resulting change in BDM).
+By using this function, we can define the compressor-reconstruction step for the MILS algorithm as shown in Listing 2.
+Listing 2: MILS step
+def MILS_step(N,input_series):
+s1,s2 = input_series.shape
+mils_v = Parallel(n_jobs=32)(delayed(MILS)(*x) for x in [[z,N] for z in input_series.reshape(s1,8,8)])
+return np.array(mils_v).reshape(s1,s2)
+In short, MILS_step performs the deletion of edges for the whole input dataset passed to the function.
+Several input samples are processed in parallel to enhance the speed of the transformation.
+It is important to notice that we consider that the compressed and reconstructed versions of an image after MILS are the same, following the rationale that changes are carried out in place.
+This compression-reconstruction step is then incorporated to the full Pipeline as presented in Listing 3.
+Listing 3: MILS Pipeline
+start = time.monotonic()
+perf_MILS=[]
+for nel in [2**k for k in range(7)]:
+transformerMILS = FunctionTransformer(partial(MILS_step,nel))
+sss = StratifiedShuffleSplit(n_splits=5, test_size=0.2, random_state=0)
+pipelineMILS = Pipeline([(’transformer’, transformerMILS),
+(’classifier’, SVC(gamma=’auto’))])
+scr = cross_val_score(pipelineMILS, (X>7).astype(int), y, n_jobs=-1, cv = sss,scoring=’balanced_accuracy’)
+perf_MILS.append([nel,scr])
+end = time.monotonic()
+general_time.append(["MILS",end-start])
+It can be seen that prior to the application of MILS, the data is binarized by simply attributing 1 to feature values greater than 7 and 0 otherwise.
+By running the Pipeline in Listing 3, the performance of MILS can be assessed.
+To illustrate how other methods were implemented, Listing 4 presents the equivalent Pipeline for PCA reduction technique.
+Listing 4: PCA Pipeline
+def PCA_step(N,input_series):
+mdl = Pipeline([(’transformer’, StandardScaler()),(’pca’, PCA(n_components=N))])
+mdl = mdl.fit(input_series)
+mdl_outp = mdl.named_steps["pca"].inverse_transform(mdl.transform(input_series))
+return mdl_outp
+perf_PCAs_quant=[]
+for fmt,tm in zip([format_info_ocp_e2m1,format_info_ocp_e4m3,format_info_bfloat16],["PCA_e2m1","PCA_e4m3","PCA_bfloat16"]):
+start = time.monotonic()
+perf_PCA_q=[]
+for nel in [2**k for k in range(7)]:
+sss = StratifiedShuffleSplit(n_splits=5, test_size=0.2, random_state=0)
+transformerPCA = FunctionTransformer(partial(PCA_step,nel))
+pipelinePCA = Pipeline([(’transformer’, transformerPCA),
+(’quantizer’,FunctionTransformer( lambda x: round_ndarray(fmt, x,sat=True))),
+(’classifier’, SVC(gamma=’auto’))])
+scrPCA = cross_val_score(pipelinePCA, round_ndarray(fmt, X,sat=True), y, n_jobs=-1, cv = sss,scoring=’balanced_accuracy’)
+perf_PCA_q.append([nel,scrPCA])
+perf_PCAs_quant.append(perf_PCA_q)
+end = time.monotonic()
+general_time.append([tm,end-start])
+Listing 4 presents the implementation of a PCA_step, similar to MILS_step, where the compression and reconstruction is carried out.
+Differently from the case of MILS, where only a binary quantisation is considered, for all the other methods the 4-bit,8-bit and 16-bit quantisations are used.
+Before fitting each model, no additional preprocessing techniques were applied except for the PCA case, where a StandardScaler() (subtracts the mean and divides by the standard deviation of each feature) was applied.
+Also, all the simulations considered the ‘scikit-learn’ implementation of the algorithms.
+Figure 3 presents the performance assessment results, which demonstrate that MILS consistently delivers higher accuracy per bit across all quantisation scenarios in the MNIST classification task.
+The mean balanced accuracy per bit is calculated as the mean balanced accuracy of the 5-fold CV divided by $64\times q$, where $q$ is the quantization level of the compressed-reconstructed figure.
+Figure 3: Performance assessment (mean accuracy per bit of compressed image) of different compression algorithms, including MILS and other traditional approaches.
+MILS substantially outperforms all the considered algorithms due to its highly compressed binary setting.
+Notably, as the number of components/edges is reduced, the performance gap between MILS and competing methods narrows, but MILS maintains a lead in all tested cases and quantisation levels.
+We attribute the superiority of MILS to its ability to optimally preserve essential structural and algorithmic complexity features of the data, leading to better generalisation in classification tasks.
+This is done while keeping the memory footprint to a minimum, since only binary values are used.
+Figure 4, however, shows that the MILS implementation used leads to high computational cost (hereby expressed as compute time) when compared to other methods.
+The compute time for MILS is at least an order of magnitude higher than conventional methods, largely due to its reliance on pre-computed algorithmic complexity measures that are only available for small (4x4) patches.
+This limitation may be mitigated through the optimisation of MILS-specific implementation, including more efficient parallel computing strategies.
+Also, it is crucial to expand the precomputed Complexity-Theoretic Measure (CTM) database for image patches larger than 4×4, reducing lookup overhead.
+Figure 4: Compute time in seconds for all the methods considered in this particular ML pipeline. MILS is less memory intensive at the cost of increasing computing times.
+Overall, MILS presents a highly promising alternative to traditional statistical Machine Learning dimensionality reduction algorithms, particularly in quantized ML scenarios.
+Its ability to leverage algorithmic complexity for superior information preservation makes it especially valuable for applications in Large Language Models (LLMs) and retrieval-augmented systems (RAGs) because MILS relies exclusively on binary operations, with potential for hardware acceleration and similarity search optimisation.
+Future work should focus on refining its computational efficiency to facilitate broader adoption in practical ML applications.
+5.2 Comparisons with other data dimensionality reduction methods
 Figs.
-34567, and 8 illustrate the effectiveness of the Minimal Information Loss Selection (MILS) method in retaining crucial local and global characteristics of synthetic and natural networks, encompassing varied types and topologies.
+56789, and 10 illustrate the effectiveness of the Minimal Information Loss Selection (MILS) method in retaining crucial local and global characteristics of synthetic and natural networks, encompassing varied types and topologies.
 These figures demonstrate MILS’s performance, typically surpassing leading graph sparsification algorithms.
@@
 The former seeks to minimise edges while maintaining the same reachability within a directed graph, while the latter, thoroughly introduced in [2], aims to reduce network dimensions based on spectral similarity of graph Laplacians, thus preserving significant properties through the adjacency matrix Laplacian spectrum.
-Specifically, Figs.
-3 and 4 showcase the performance of MILS in preserving the degree distribution and edge betweenness distribution in a typical synthetically (recursively) generated Erdős-Rényi (ER) random graph with low edge density, comparing its results against both random edge deletion and spectral sparsification methods.
-While MILS is not significantly better at preserving the clustering coefficient of random networks, Fig.
-5 shows that MILS does significantly better at preserving the clustering coefficient of real-world (biological, social, and electric grid) networks taken from [49], outperforming both transitive and spectral reduction/sparsification methods.
-Figs. 6, 7 and 8 illustrate how MILS outperforms spectral sparsification at preserving edge betweenness, and degree and eigenvector centralities, respectively.
-4.2 An application of MILS to image lossy compression
-Fig.
-9(A) and Fig.
-9(B) illustrate the application of MILS in the reduction and coarse-graining of two simple cases of the space-time evolution of Elementary Cellular Automata (ECA) rules 22 and 158.
-The reduction is by minimisation of algorithmic information loss.
-MILS effectively extracts the salient elements that characterise each of these systems.
-The examples in Fig.
-9(A) and Fig.
-9(B) depicts the way in which regions with high or low algorithmic content can be ranked, selected, or preserved for dimensional reduction purposes.
-The method performs an unsupervised lossy compression able to preserve both cellular automata’s main features (not covered in grey), with no intervention and no parameter choice.
-MILS proceeds by deleting the regions with the lowest algorithmic information content contribution and maximizing the preservation of the features that contribute the most to the algorithmic description of the objects.
-The extracted features will not be as clear as in these examples as they may pick more complicated patterns, even not statistical ones based on algorithmic probability.
-Unlike statistical approaches, the algorithm can approximate (and thus preserve/extract) features that are of an algorithmic nature and which are not statistically apparent as it was in this case (see [68, 66]) and next examples.
-Fig.
-10 shows the MILS algorithm applied to images.
-Fig.
-10(A) shows how vertical and horizontal compression preserves features even if distorting the image hence showing different mechanisms and goals than those from popular image compression algorithms.
-The purpose of image compression algorithms such as JPEG is to maximise storage compressibility and require a decoder, here the application is directly to the image itself both at storage and visualisation stages.
-Is also important to mention that our purpose is not to maximise storage compression but to minimise the loss of algorithmic content in the reduction process.
-Fig.
-10(C) shows how the algorithm preserves the main features of the image leaving almost intact the formulae.
-MILS is highly context-sensitive and adapts to different kind of information such as images.
-For example, it handles in different ways images as seen in Fig.
-10.
-This is because information in Fig.
-10(A) is distributed more uniformly than Fig.
-10(C) but, according to the algorithm, some features in certain regions are more important than others and thus some distortions are allowed if such features have high algorithmic content according to the underlying computer programs that reconstruct the compressed images.
-However, in Fig.
-10(C), blank space between formulae can be sacrificed first as they do not contain any information.
-While this behaviour for handling different cases can be replicated in ad-hoc algorithms, MILS handles it naturally and determines when distorting can be allowed or proportion can be preserved, something that image compression algorithms would not do as their goal is to recover the same image dimensions and maximise storage compression and not to preserve algorithmic information content even sacrificing otherwise fundamental properties of the image for those algorithms such as image dimension.
-5 Conclusion
-In this paper, we have introduced Minimal Information Loss Selection (MILS), a novel agnostic approach for data summarisation that is particularly effective in reducing data dimensionality while preserving essential features such as degree distribution, clustering coefficient, edge betweenness, and centralities.
-As studied in Section 4.1, MILS demonstrates comparable or superior performance to existing network reduction algorithms on different classes and types of networks.
-Moreover, it extends to lossy compression of images and bi-dimensional data as shown in Section 4.2.
-It operates on the principle of local perturbation analysis and algorithmic information dynamics, offering robustness against varying methods of approximating algorithmic complexity.
-The method presented in this article optimally preserves the algorithmic information content (which might be affected by the local perturbations) whenever an optimal method is used to approximate algorithmic complexity.
-This guarantees that its performance can only increase as more computational resources are made available, which differs from other methods based on statistics for which theoretical limitations are known to hold.
-Nevertheless, the empirical application of MILS only scales linearly with the worst-case time complexity of the chosen method for approximating algorithmic complexity.
-Funding
-FSA acknowledges support from the São Paulo Research Foundation (FAPESP), grants $2023$/$05593$-$1$ and $2021$/$14501$-$8$. NAK is supported by Horizon $2020$ EU grant (MultipleMS, $733161$).
-Appendix A Figures
 Figure 5: MILS or neutral edge deletion (blue) outperforms random edge deletion (red) at preserving both edge degree distribution (top, showing removed edges) and edge betweenness distribution (bottom) on an Erdős-Rényi random graph of vertex size 100 and low edge density ($\sim 4\%$) after up to 60 edges were removed (degree distribution comparison) and 150 edges were removed (edge betweenness) out of a total of 200 edges (notice also the scale differences on the $x$-axis).
@@
 Green highlights the overlapping and the preserved area of the distributions after random deletion (top), MILS and spectral removal (bottom pairs).
-Figure 5: MILS mean clustering coefficient preservation against two other sophisticated graph sparsification methods based on graph spectral and transitive reduction techniques on biological, electric and social networks taken from [49].
+Specifically, Figs.
+5 and 6 showcase the performance of MILS in preserving the degree distribution and edge betweenness distribution in a typical synthetically (recursively) generated Erdős-Rényi (ER) random graph with low edge density, comparing its results against both random edge deletion and spectral sparsification methods.
+While MILS is not significantly better at preserving the clustering coefficient of random networks, Fig.
+7 shows that MILS does significantly better at preserving the clustering coefficient of real-world (biological, social, and electric grid) networks taken from [41], outperforming both transitive and spectral reduction/sparsification methods.
+Figure 7: MILS mean clustering coefficient preservation against two sophisticated graph sparsification methods based on graph spectral and transitive reduction techniques on biological, electric and social networks taken from [41].
 The transitive method does not allow selection of edges to be deleted, and in some cases it either fails to significantly reduce the network edge density if no cycles are present (such as, generally, in electric and genetic networks) and/or takes the clustering coefficient to 0 (e.g.
@@
 Comparisons with other methods are unnecessary because they destroy local or global properties by design, such as clustering coefficients for the spanning tree algorithm.
+Figs. 8, 9 and 10 illustrate how MILS outperforms spectral sparsification at preserving edge betweenness, and degree and eigenvector centralities, respectively.
 Figure 8: Stacked histograms showing edge betweenness preservation of MILS versus spectral sparsification across different families of networks.
@@
 On average MILS preserved the eigenvector centrality distribution of these representative networks 1.5 times better.
+5.3 An application of MILS to image lossy compression
+Fig.
+11(A) and Fig.
+11(B) illustrate the application of MILS in the reduction and coarse-graining of two simple cases of the space-time evolution of Elementary Cellular Automata (ECA) rules 22 and 158.
+The reduction is by minimisation of algorithmic information loss.
+MILS effectively extracts the salient elements that characterise each of these systems.
+The examples in Fig.
+11(A) and Fig.
+11(B) depicts the way in which regions with high or low algorithmic content can be ranked, selected, or preserved for dimensional reduction purposes.
+The method performs an unsupervised lossy compression able to preserve both cellular automata’s main features (not covered in grey), with no intervention and no parameter choice.
+MILS proceeds by deleting the regions with the lowest algorithmic information content contribution and maximizing the preservation of the features that contribute the most to the algorithmic description of the objects.
+The extracted features will not be as clear as in these examples as they may pick more complicated patterns, even not statistical ones based on algorithmic probability.
+Unlike statistical approaches, the algorithm can approximate (and thus preserve/extract) features that are of an algorithmic nature and which are not statistically apparent as it was in this case (see [16, 11]) and next examples.
 Figure 11: In (A) and (B) we show a feature selection and image reduction by application of MILS, starting from the original and second step, highlighting the regions that are earmarked to be omitted (in grey) versus the features that are kept along the way, thereby optimally preserving the main properties of these objects, properties whose persistence enables a ranking of such features.
@@
 The MILS algorithm selects features which are considered most important in the compressed image reconstruction.
+Fig.
+12 shows the MILS algorithm applied to images.
+Fig.
+12(A) shows how vertical and horizontal compression preserves features even if distorting the image hence showing different mechanisms and goals than those from popular image compression algorithms.
+The purpose of image compression algorithms such as JPEG is to maximise storage compressibility and require a decoder, here the application is directly to the image itself both at storage and visualisation stages.
+Is also important to mention that our purpose is not to maximise storage compression but to minimise the loss of algorithmic content in the reduction process.
+Fig.
+12(C) shows how the algorithm preserves the main features of the image leaving almost intact the formulae.
 Figure 12: Minimal Information Loss Reduction, a MILS-based lossy compression algorithm.
@@
 C: Compressed formula by MILS vs random row deletion (D) with resulting images of the same size but MILS preserving text proportion and minimizing information loss while random row deletion distorts the text.
-Appendix B Preliminary definitions and background results
-B.1 Cellular automata
-To illustrate how our algorithm works and can be applied we will use a very well studied dynamical system called a cellular automaton.
-A cellular automaton is a computer program that applies in parallel a global rule composed of local rules on a tape of cells with symbols (e.g.
-binary).
-Broadly speaking, it consists of a set of states, a neighbourhood template and a local rule $f$.
-Here we will use space-time diagrams of cellular automata to illustrate the way in which the MILS algorithm operates.
-Thoroughly studied in [1], Elementary Cellular Automata (or ECA) are defined as one-dimensional cellular automata that take into consideration in their local rules the cells next to the centre cell and the centre cell.
-In the case of 1-dimensional CA it is common to introduce the radius of the neighbourhood template, which can be written as $\langle-r,-r+1,\dots,r-1,r\rangle$ and has length $12r+$ 1 cells.
-Given a radius $r$, the local rule is a function $f:\mathbb{Z}_{|S|}^{{|S|}^{(2r+1)}}\rightarrow\mathbb{Z}_{|S|}$, where $S$ is the set of single-cell states.
-Elementary Cellular Automata (ECA) have a radius $r=1$ (closest neighbours), having the neighbourhood template $\langle-1,0,1\rangle$, meaning that the three-cell neighbourhood comprises a central cell.
-From this it follows that the rule space for ECA contains $2562^{2^{3}}=$ 256 rules.
-Enumeration of ECA rules: It is common to follow the lexicographic ordering scheme introduced by Wolfram [1]. According to this encoding, the 256 ECA rules can be encoded by 8-bits.
-A space-time diagram captures the evolution of a cellular automaton for a given initial condition and is read from the top starting from time step $t=0$ (the initial condition) and evolving towards the bottom of the diagram (see Figs.
-1.(A) and 1.(B) in the main article).
-B.2 Classical Information Theory and Shannon Entropy
-Central to information theory is the concept of Shannon’s information entropy, which quantifies the average number of bits needed to store or communicate the statistical description of an object.
-Definition B.1.
-Let $X=(R,p)$ be an ensemble, where $R$ is the set of all possible outcomes (the random variable), $n=|R|$ and $p(x_{i})$ is the probability of an outcome $x_{i}\in R$.
-The Shannon entropy of $X$ is given by
-$H(X)=-\sum_{i=1}^{n}p(x_{i})\log_{2}p(x_{i}).$ (2)
-It is clear from this definition that in order to calculate $H(X)$ one has to know (or assume the existence of) a probability mass function of the ensemble $X$.
-While entropy can be used as a measure of the information contained in a graph (or another object), we will restrict our analyses to the Block Decomposition Method, which we describe in Section B.6.
-B.3 Graph definitions and notation
-Graphs (or networks) are mathematical structures that allow a simple representation of systems consisting of a set of agents and a set of interactions between them.
-Formally, we may define a graph as follows.
-Definition B.2.
-A (labeled)111 In this article, we will only deal with labeled graphs.
-graph is an ordered pair $G=(V,E)$ comprising a finite set $V(G)=\{1,\dots,n\}$ of labeled vertices with $n\in\mathbb{N}$ and a set $E(G)\subseteq\mathbb{E}(G)=V\times V$ of edges.
-For simplicity of notation, we will often let $V=V(G)$ and $E=E(G)$.
-A graph $H$ is contained in $G$, and we write $H\subseteq G$, if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$.
-In this case, we say $H$ is a subgraph of $G$.
-If either $V(H)\neq V(G)$ or $E(H)\neq E(G)$, we say $H$ is a proper subgraph of $G$, and we normally use $H\subset G$ to denote this situation.
-If on the other hand $V(H)=V(G)$ and $E(H)=E(G)$, we say the graphs $H$ and $G$ are equal.
-A directed graph is a graph where the edges have a direction associated with them. In a directed graph (possibly with self-loops), we represent edges by ordered pairs
-$(v,w)\in\mathbb{E}(G)\coloneqq\{(x,y)\mid x,y\in V\}=V\times V$
-instead of unordered pairs as in Definition B.3, such that $v$ is the source vertex and $w$ is the target vertex of the edge.
-We call a traditional directed graph if $G$ is a directed graph without self-loops, so that
-$(v,w)\in\mathbb{E}_{d}(G)\coloneqq\mathbb{E}(G)\setminus\{(x,x)\mid x\in V\}$
-Definition B.3.
-A labeled undirected graph $G=(V,E)$ without self-loops is a labeled graph with a restriction $\mathbb{E}_{c}$ in the edge set $E$ such that each edge is an unordered pair with
-$E\subseteq\mathbb{E}_{c}\left(G\right)\coloneqq\{\{x,y\}\mid x,y\in V\}$
-We also refer to these graphs as simple labeled (or classical) graphs. In this paper, we will normally refer to simple graphs as just graphs.
-We say that two vertices $v,w\in V$ are adjacent if there is an edge $e=\{v,w\}$ in $E$ that connects them, and call $v$ and $w$ the endpoints of such undirected edge.
-The adjacency matrix $A=(a_{ij})$ of $G$ is a matrix that encodes all the adjacency relationships of the graph.
-The entries of this matrix are $a_{vw}=1$ if vertices $v$ and $w$ are adjacent and $a_{vw}=0$ otherwise.
-Note that while the adjacency matrix of an undirected graph is always symmetric, this is not necessarily true for a directed graph where adjacency can be asymmetric.
-The list of eigenvalues of the graph’s adjacency matrix sorted from largest to smallest is known as the spectrum of the graph.
-The degree $d(v)$ of $v$ is the number of vertices to which $v$ is connected, that is $d(v)=\sum_{w\in V}a_{vw}$.
-The degree matrix $D=(d_{ij})$ of $G$ is a diagonal matrix that contains information about the degree of each vertex, and is defined by $d_{ij}=d(i)$ if $i=j$ and $d_{ij}=0$ otherwise.
-We are now in the position to introduce the notion of graph Laplacian, which will be of interest for us in forthcoming sections.
-Consider the following definition.
-Definition B.4.
-Let $G$ be a simple graph on $n$ vertices. The Laplacian matrix (of order $n$) of the graph $G$ is defined as $L=D-A$, where $D$ is the degree matrix and $A$ is the adjacency matrix of $G$.
-Throughout the text, and trusting that it does not lead to confusion, we will use the terms nodes and vertices, and links and edges, interchangeably.
-B.4 Spectral sparsification
-Generally speaking, the goal of network sparsification is to approximate a given graph $G$ by a sparse graph $H$ on the same set of vertices.
-If $H$ is close to $G$ in some appropriate metric, then $H$ can be used as a signature preserving important properties of $G$ for faster computation after reducing the size of $G$ and without introducing too much error.
-Obvious trivial sparsification methods include edge deletion by some criterion, such as the outermost ones (called the $k$-shell method [22, 39], often used to identify the core and the periphery of the network), but most of these methods (such as this shell one) are rather arbitrary or ad-hoc, rather than general methods aimed at preserving important algebraic, topological or dynamical properties of the original graph, all of which constitute and contribute to the information content of the graph, that is, the information necessary to fully describe a network and reconstruct the network from that description.
-A popular sparsification algorithm is the spanning tree [36] designed to preserve node distance but clearly destroy all other local node properties, such as the clustering coefficient.
-Not many non-trivial methods for network sparsification exist today.
-Some clearly destroy local properties, such as the spanning tree algorithm, which destroys the clustering coefficient.
-It is acknowledged [58, 59, 19], however, that spectral graph sparsification is among the most efficient, both at preserving important algebraic and dynamical properties of a network and in terms of fast calculation.
-In part the dearth of methods is due to a lack of assessment tools to decide whether one method is better than another in general terms, rather than being designed to preserve one or another specific graph theoretic property (e.g.
-the transitive edge deletion method destroys the clustering coefficient of the original graph [14]).
-The spectral method considered in this paper is a high-quality algorithm described in [19, 58].
-Transitive reduction was introduced in [14].
-A graph $G$ is said to be transitive if, for every pair of vertices $u$ and $v$, not necessarily distinct, $(u,v)\in E(G)$ whenever there is a directed path in $G$ from $u$ to $v$.
-That is, if there is a path from a vertex $x$ to a vertex $y$ in graph $G$, there must also be a path from $x$ to $y$ in the transitive reduction of $G$, and vice versa.
-If a given graph is a finite directed acyclic graph, its transitive reduction is unique, and is a subgraph of the given graph.
-Graph sparsification is the approximation of an arbitrary graph by a sparse graph.
-We compare MILS against random, simple (e.g.
-spanning tree) and two powerful graph sparsification and reduction methods (spectral and transitive).
-Spectral graph sparsification is based on the spectral similarity of graph Laplacians.
-A spectral sparsifier is a subgraph of the original whose Laplacian quadratic form is approximately the same as that of the original graph on all real vector inputs.
-Spectral graph sparsification is a stronger notion than cut sparsifiers [59] and is considered one of the most, if not the most, sophisticated sparsification or network reduction method, as it is believed to preserve some of the most important algebraic, topological and potentially dynamical properties of a network.
-B.5 Graph-theoretic measures
-The global clustering coefficient of $G$ is the fraction of paths of length 2 in $G$ that are closed over all paths of length two in $G$.
-The mean or average clustering coefficient is the mean over all local clustering coefficients of vertices of $G$.
-The betweenness centrality for a vertex $i$ in a connected graph is given by $\sum_{s,t\in v\land s\neq i\land t\neq i}\frac{n_{s,t}^{i}}{n_{s,t}}$, where $n_{s,t}$ is the number of shortest paths from s to t and $n^{i}_{s,t}$ is the number of shortest paths from $s$ to $t$ passing through $i$.
-The ratio $\frac{n_{s,t}^{i}}{n_{s,t}}$ is taken to be zero when there is no path from $s$ to $t$.
-Degree centrality is a measure of the centrality of a node in a network and is defined as the number of edges (including self-loops) that lead into or out of the node.
-The degree centrality of $G$ is the list of nonnegative integers (“degree centralities") lying between 00 and $n-1$ inclusive, where $n$ is the number of vertices of $G$, and identifies nodes in the network by their influence on other nodes in their immediate neighbourhood.
-Eigenvector centrality is a list of normalized nonnegative numbers (“eigenvector centralities", also known as Gould indices) that are particular centrality measures of the vertices of a graph.
-Eigenvector centrality is a measure of the centrality of a node in a network based on the weighted sum of centralities of its neighbours.
-It therefore identifies nodes in the network that are connected to many other well-connected nodes.
-For undirected graphs, the vector of eigenvector centralities $c$ satisfies the equation $c=1/\lambda_{1}a.c$, where $\lambda_{1}$ is the largest eigenvalue of the graph’s adjacency matrix $a$.
-B.6 Algorithmic information theory and approximations to graph information content
-Here, we present some basics concepts regarding algorithmic information theory. For a complete introduction to these concepts and theorems, see [34, 24, 26, 43, 18].
-First, for some preliminary notation, let $\{0,1\}^{*}$ be the set of all finite binary strings.
-Let $|X|\in\mathbb{N}$ denote the number of elements (i.e., the cardinality) in a set, if $X$ is a finite set, or the number of elements in a data structure, if $X$ is computational object.
-In the particular case of strings, let $l(x)$ denote the length of a finite bit string $x\in\{0,1\}^{*}$.
-Let $(x)_{2}$ denote the binary representation of the number $x\in\mathbb{R}$.
-Let $x\upharpoonright_{n}$ denote the ordered sequence of the first $n$ bits of the fractional part in the binary representation of $x\in\mathbb{R}$.
-That is, $x\upharpoonright_{n}=x_{1}x_{2}\dots x_{n}\equiv(x_{1},x_{2},\dots,x_{n})$, where ${(x)}_{2}=y.x_{1}x_{2}\dots x_{n}x_{n+1}\dots$ with $y\in\{0,1\}^{*}$ and $x_{1},x_{2},\dots,x_{n}\,\in\{0,1\}$.
-With respect to weak asymptotic dominance of function $f$ by a function $g$, we employ the usual $\mathbf{O}(g(x))$ for the big O notation, $\mathbf{\Omega}(g(x))$ for the big Omega notation, and $\mathbf{\Theta}(g(x))$ for the big Theta notation; and with respect to strong asymptotic dominance by a function $g$, we employ the usual $\mathbf{o}(g(x))$ when $g$ dominates $f$, ${\bm{\omega}}(g(x))$ when $f$ dominates $g$, and $\sim g(x)$ when both $f$ and $g$ dominates each other.
-The reader should pay attention not to confuse the big Omega notation (in boldface) $\mathbf{\Omega}(g(x))$ with the halting probability $\Omega$ as in Definition B.10.
-A Turing machine $\mathbf{M}$ is a general abstraction of a computer program, similar to cellular automata but sequential, that given an input, may either produce an output (and halt) or never halt.
-As one of the most fundamental mathematical objects in theoretical computer science, the Turing machine is thus an algorithmically mechanistic causal explanation of an deterministic process’ outcome, and underpins the central ideas of the algorithms introduced here.
-With respect to the chosen computational model, let $\mathbf{L^{\prime}_{U}}$ denote a binary prefix-free [^2] The group of valid programs constitutes a prefix-free set (no element is a prefix of any other, which is a property necessary to keep $10<\mathbf{m}(G)<$ 1).
-(or self-delimiting) universal programming language for a prefix universal Turing machine $\mathbf{U}$.
-For example, a program of $\mathbf{U}$ may represent a 2-dimensional (prefix-free) Turing machine (a typical deterministic Turing machine whose single head can move up and down as well as left and right) starting from an empty 2-dimensional grid (instead of the typical 1-dimensional tape Turing machine) as defined in [61, 68].
-As in [43, 34, 18]:
-Notation B.5.
-[…]
-$l\left(\left<x,y\right>\right)\leq l(x)+l(y)+\mathbf{O}\left(\log_{2}\left(l(x)+l(y)\right)\right)$
-holds and the bit string $x,y\right>$ is computably recognizable as an ordered pair in polynomial time. For example, with an encoding in the form
-$x,y\right>\coloneqq$
-$\underbrace{0\cdots 0}_{\log_{2}\left(l\left(\left(2\right)_{2}\right)\right)\text{ times}}1\,\left(2\right)_{2}\underbrace{0\cdots 0}_{\log_{2}\left(\log_{2}\left(l(x)\right)\right)\text{ times}}1\,\left(l(x)\right)_{2}\,x\,\underbrace{0\cdots 0}_{\log_{2}\left(\log_{2}\left(l(y)\right)\right)\text{ times}}1\,\left(l(y)\right)_{2}\,y\text{ .}$
-[…]
-Definition B.6.
-The (unconditional) prefix333 In this paper, we will normally refer to prefix algorithmic complexity as just algorithmic complexity.
-algorithmic complexity (also known as K-complexity, prefix Kolmogorov complexity, prefix program-size complexity or prefix Solomonoff-Komogorov-Chaitin complexity) of a finite binary string $w$, denoted here by $C(w)$, is the length of the shortest program $w^{*}\in\mathbf{L^{\prime}_{U}}$ such that $\mathbf{U}(w^{*})=w$444 $w^{*}$ denotes the lexicographically first $\mathrm{p}\in\mathbf{L^{\prime}_{U}}$ such that $l(\mathrm{p})$ is minimum and $\mathbf{U}(p)=w$.
-In addition, when an input is informed to the Turing machine, we also have that the conditional prefix algorithmic complexity of a binary finite string $y$ given a binary finite string $x$, denoted by $C(y\,|x)$, is the length of the shortest program $w\in\mathbf{L^{\prime}_{U}}$ such that $\mathbf{U}(\left<x,w\right>)=y$.
-Note that $C(y)=C(y\,|\epsilon)$, where $\epsilon$ is the empty string.
-Moreover, the joint prefix algorithmic complexity of strings $x$ and $y$ is denoted by $C(x,y)\coloneqq C(\left<x,y\right>)$.
-From algorithmic information theory (AIT) [43, 26, 34, 24], one knows the following results:
-Lemma B.1.
-For every $x,y\in\{0,1\}^{*}$ and $n\in\mathbb{N}$,555 For the sake of simplicity, we choose to omit the ceiling operator on the logarithm when calculating the length of encoding natural numbers, so that in such cases $\log_{2}\left(x\right)$ actually stands for $\lceil\log_{2}\left(x\right)\rceil$ in the present paper.
-Note that in terms of asymptotic dominance and bounds it is indeed irrelevant.
-$C(x)$ $\leq l(x)+\mathbf{O}(\log_{2}(l(x)))$ (3)
-$C(y\,|x)$ $\leq C(y)+\mathbf{O}(1)$ (4)
-$C(y)$ $\leq C(x\,y)+C(l(x))+\mathbf{O}(1)$ (5)
-$C(x\,y)$ $\leq l(y)+C(l(y))+C(x)+\mathbf{O}(1)$ (6)
-$C(x)\leq C(x,y)+\mathbf{O}(1)$ $\leq\;C(y)+C(x\,|y)+\mathbf{O}(1)$ (7)
-$C(n)$ $=\mathbf{O}(\log_{2}(n))$ (8)
-Lemma B.2.
-[…]
-$C(f_{c}(n))\leq C(n)+\mathbf{O}(1)$
-For graphs (as those defined in Section B.3) we use the notation:
-Notation B.7.
-Let $\left(e_{1},\dots,e_{\left|\mathbb{E}(G)\right|}\right)$ be a previously fixed ordering (or indexing) of the set $\mathbb{E}(G)$.
-For an edge set $E(G)$, let $C(E(G))\coloneqq C(\left<E(G)\right>)$, where the edge set string $E(G)\right>$ denotes
-[…]
-such that666 See also Definition B.8.
-$z_{i}=1\iff e_{i}\in E(G)\text{ ,}$
-where $z_{i}\in\{0,1\}$ with $1\leq i\leq n=|\mathbb{E}(G)|$. The same applies analogously to the conditional and joint cases of the algorithmic complexity.
-From [9, 11, 12], we remember that the family of all possible simple graphs is always recursively labeled.
-That is, there is a unified recursive way to encode a sequence of all possible undirected edges given any unordered pair $\{x,y\}$ of natural numbers $x,y\in\mathbb{N}$.
-Thus, we know that, for any arbitrarily chosen method of recursively encoding a matrix into a string,
-$\left|C(G)-C(E(G))\right|=\mathbf{O}(1)$
-holds, where $C(G)$ denotes the algorithmic complexity of generating the adjacency matrix of any arbitrarily chosen $G$.
-Hence, one can interchangeably employ $C(G)$ or $C(E(G))$ to measure the algorithmic information content of a graph when dealing with the theoretical optimal form of the algorithmic complexity as in Definition B.6.
-Moreover, from [9, 11, 12], we will have that
-$C(G\,|\,x)$ $=\mathbf{O}(1)$ (9)
-$C(x\,|\,G)$ $=\mathbf{O}(1)$ (10)
-$C(x)$ $=C(G)\pm\mathbf{O}\left(1\right)\text{ ,}$ (11)
-where $x\in\{0,1\}^{*}$ is the characteristic string of $G$ [9]. The characteristic string of a graph is any bit string that determines the presence or absence of edges in $G$:
-Definition B.8.
-Let $\left(e_{1},\dots,e_{\left|\mathbb{E}(G)\right|}\right)$ be any arbitrarily fixed ordering of all possible edges of a graph $G$ as in Notation B.7.
-We say that a string $x\in\{0,1\}^{*}$ with $l(x)=\left|\mathbb{E}(G)\right|$ is a characteristic string of a graph $G$ iff, for every $e_{j}\in\mathbb{E}(G)$,
-[…]
-where $1\leq j\leq l(x)$.777 And the same definition applies analogously to the simple graph case.
-Thus, we have that the algorithmic information content of a graph is equivalently captured (except for a constant that only depends on the chosen universal programming language) by the algorithmic complexity of the adjacency matrix, the edge set string, or of the characteristic string.
-Note that the same cannot always be said in the case of multilayered ou multidimensional networks [11, 12].
-See also [9, 10] for more details on characteristic strings and algorithmically characteristic strings.
-We know from the algorithmic coding theorem [26, 24, 43, 34] that
-$C\left(x\right)=$ (12)
-$-\log\left(\sum\limits_{\mathbf{U}\left(p\right)=x}\frac{1}{2^{\left|p\right|}}\right)\pm\mathbf{O}(1)=$
-$-\log\left(\mathbf{m}\left(x\right)\right)\pm\mathbf{O}(1)\text{ ,}$
-holds, where: $\mathbf{m}\left(\cdot\right)$ is a maximal computably enumerable semimeasure; and
-$\mathbf{P}_{\mathbf{U}}\left[\text{``event }x\text{ occurs''}\right]=\sum\limits_{\mathbf{U}\left(p\right)=x}2^{-\left|p\right|}$
-denotes the universal a priori probability of an arbitrary event.
-$\mathbf{P}_{\mathbf{U}}$ can be understood as the probability of randomly generating (by an i.i.d.
-stochastic process) a prefix-free (or self-delimiting) program that outputs $x$.
-A computably enumerable semimeasure $\mathbf{m}\left(\cdot\right)$ is said to be maximal if, for any other computably enumerable semimeasure $\mu\left(\cdot\right)$ with domain defined for all possible encoded objects, where $\sum\limits_{x\in\left\{0,1\right\}^{*}}\mu\left(x\right)\leq 1$, there is a constant $C>0$ (which does not depend on $x$) such that, for every encoded object $x$, $\mathbf{m}\left(x\right)\geq C\,\mu\left(x\right)\text{ .}$ We also have it that $2^{-\mathbf{K}\left(x\right)}$ is called the algorithmic probability of $x$, which therefore converges to the universal a priori probability (except of an object-independent constant).
-Also note that the algorithmic coding theorem applies analogously to the conditional algorithmic complexity $\mathbf{K}(z\,|w)$.
-The concept of algorithmic probability (also known as Levin’s semi-measure) yields a method for approximating algorithmic complexity related to the frequency of patterns in the adjacency matrix of a network, including therefore the number of subgraphs in a network.
-Following from Equation (12), the algorithmic probability [57, 42, 25] of a graph $G$ is a measure that describes the probability that a random computer program $p$ will produce $G$ when run on a 2-dimensional tape universal (prefix-free) Turing machine $\mathbf{U}$.
-That is,
-$\mathbf{m}(G)=\sum_{\mathbf{U}(p)=A_{G}}1/2^{|p|}\text{ .}$
-An example of a popular 2-dimensional tape Turing machine is Langton’s ant [41], commonly referred to as a Turmite.
-The Algorithmic Coding Theorem [26, 43, 34, 24] therefore establishes the connection between $\mathbf{m}(G)$ and $C(G)$ as:
-$|-\log_{2}\mathbf{m}(G)-C(G)|<c\text{ ,}$ (13)
-where $c$ is some fixed constant, independent of $G$.
-This way, the size $C(G)$ of the algorithmic information content of a graph $G$ (i.e., its prefix algorithmic complexity) is given by the so-called algorithmic coding theorem [42, 30, 56] by simply rewriting Eq.
-13 as:
-$C(G)=-\log_{2}\mathbf{m}(G)\pm\mathbf{O}(1)\text{ .}$ (14)
-In [32] a technique called Coding Theorem Method (CTM) was advanced for approximating $\mathbf{m}(G)$ by means of a function that considers all Turing machines of increasing size (by number of states).
-Indeed, for Turing machines $\mathbf{M}$ with $n$ states, $k$ alphabet symbols (usually 2 alphabet symbols only), and empty input $\lambda$ in which $n$ is small enough, we have that
-$\textit{D}(n,k)(s)=\frac{\left|\left\{(\mathbf{M}\in\left(n,k\right)\colon\mathbf{M}\left(\lambda\right)=s\right\}\right|}{\left|\left\{(\mathbf{M}\in\left(n,k\right)\colon\exists w^{\prime}\left(\mathbf{M}\left(\lambda\right)=w^{\prime}\right)\right\}\right|}$ (15)
-is computable for values of the Busy Beaver problem [54] that are known, providing a means to numerically approximate the algorithmic complexity of small graphs, such as network motifs.
-In Equation (15), $\mathbf{M}\left(x\right)=y$ denotes machine $\mathbf{M}$ returning $y$ with input $x$ and, therefore, $\textit{D}(n,k)(s)$ gives the frequency of halting Turing machines with $n$ states and $k$ alphabet symbols that ended up returning $s$ as output.
-The algorithmic coding theorem then establishes that graphs produced with lower frequency by random computer programs have higher algorithmic complexity, and vice versa.
-In this article, we will use the Block Decomposition Method (BDM) [68, 61, 72] as an estimator of algorithmic complexity C(G) and, hence, of the information difference $I(G,e_{i})$, but the algorithm and methods introduced are independent of the particular method used to approximate algorithmic complexity.
-The BDM consists in decomposing the adjacency matrix of a graph into subgraphs of sizes for which complexity values have been estimated, then reconstructing an approximation of the algorithmic complexity of the graph by adding the complexity of the individual pieces according to the rules of information theory, as follows:
-$K_{log}(G)=\sum_{(r_{u},n_{u})\in Adj(G)_{d\times d}}\log_{2}(n_{u})+K_{m}(r_{u})\text{ ,}$ (16)
-where $Adj(G)_{d\times d}$ represents the set with elements $(r_{u},n_{u})$ obtained when decomposing the adjacency matrix of $G$ into all subgraphs of size $d$ contained in $G$ and $K_{m}(r_{u})$ is an approximation to $C(r_{u})$ using the CTM [32, 56, 55, 62] from Equations (14) and (15).
-In each $(r_{u},n_{u})$ pair, $r_{u}$ is one such submatrix of the adjacency matrix and $n_{u}$ its multiplicity (number of occurrences).
-As can be seen from the formula, repeated subgraphs only contribute to the complexity value with the subgraph BDM complexity value once plus a logarithmic term as a function of the number of occurrences.
-This is because the information content of subgraphs is only sub-additive, as one would expect from the growth of their description lengths.
-Applications of CTM and BDM have been explored in [32, 56, 55, 62], and include applications to graph theory and complex networks [61] and [62], where the technique was first introduced.
-The only parameters used in the application of BDM are the use of strings up to 12 bits for strings and 4 bits for arrays given the current best CTM approximations [56] and the suggestions in [68] based on an empirical distribution based on all Turing machines with up to 5 states, and with no string/array overlapping in the decomposition for maximum efficiency (as it runs in linear time), and for which the error (due to boundary conditions) has been shown to be bounded [68].
-Definition B.9.
-We say a real number $x\in\left[0,1\right]\subset\mathbb{R}$ is 1-random (K-incompressible up to a constant, K-random or prefix algorithmically random) if and only if it satisfies
-$C(x\upharpoonright_{n})\geq n-\mathbf{O}(1)\text{ ,}$
-where $n\in\mathbb{N}$ is arbitrary.
-In order to avoid ambiguities between plain and prefix algorithmic complexity and ambiguities in relation to randomness deficiencies, we choose to say that an algorithmically random real number with respect to prefix algorithmic complexity in Definition B.9 is $\mathbf{O}(1)$-K-random.
-One of the most important results in algorithmic information theory is the investigation and proper formalisation of a mathematical theory for randomness [43, 34, 24].
-We will employ one of these important mathematical objects: the halting probability (see Definition B.10).
-This is a widely known example of infinite binary sequence, or real number, that is algorithmically random with respect to prefix algorithmic complexity, i.e., $\mathbf{O}(1)$-K-random.
-Definition B.10.
-Let $\Omega\in\left[0,1\right]\subset\mathbb{R}$ denote the halting probability (also known as Chaitin’s constant or Omega number). The halting probability is defined by
-$\Omega=\sum\limits_{\begin{subarray}{c}\exists y\left(\mathbf{U}(\mathrm{p})=y\right)\\ \mathrm{p}\in\mathbf{L^{\prime}_{U}}\end{subarray}}\quad\frac{1}{2^{l(\mathrm{p})}}$
-Theorem B.3.
-Let $n\in\mathbb{N}$. Then,
-$C(\Omega\upharpoonright_{n})\geq n-\mathbf{O}(1)$
-That is, $\Omega$ is $\mathbf{O}(1)$-K-random.
-The same analogous notion of incompressibility can be assigned to classical graphs (see Definition B.3) as in [9, 10]:
-Definition B.11.
-We say a simple graph $G$ is $\mathbf{O}(1)$-K-random iff it satisfies
-$C(E(G))\geq\binom{\left|V(G)\right|}{2}-\mathbf{O}(1)$
-This way, we have from [9] that:
-Corollary B.4.
-There is an infinite number of classical graphs $G$ (as in Definition B.3) that are $\mathbf{O}(1)$-K-random.
-In particular, those in which their characteristic strings are long enough initial segments of a $\mathbf{O}(1)$-K-random real number.
-B.7 Information difference
-The information difference is based on a concept of algorithmic/causal perturbation analysis as introduced in [61, 64, 72].
-The procedure consists in the perturbation of all elements of a system by the removal of elements whose effects on its algorithmic information content are measured and ranked accordingly.
-Formally:
-Definition B.12.
-Let $G$ be a graph with edges $e_{1},\ldots,e_{|E(G)|}\in E(G)\subseteq\mathbb{E}(G)$.
-Now, let $G\backslash e_{i}$ denote the graph obtained by removing $e_{i}$ from $G$, and $I(G,e_{i})$ denote the information difference or information value/contribution of $e_{i}$ to $G$ given by
-$I(G,e_{i})\coloneqq C(G)-C(G\backslash e_{i})\text{ ,}$
-where $C(G)$ is the algorithmic complexity of graph $G$. If one allows multiple edge deletions of a subset $F\subseteq E(G)$, we analogously define
-$I(G,F)\coloneqq C(G)-C(G\setminus F)\text{ .}$
-Also, for the analogously inverse operation of edge insertion $G\cup e_{i}$, we denote the information difference by
-$I^{-1}(G,e_{i})\coloneqq C(G)-C(G\cup e_{i})$
-and
-$I^{-1}(G,F)\coloneqq C(G)-C(G\cup F)\text{ ,}$
-respectively.
-These differences can also be referred to the graph’s nodes, but here we will restrict our attention to the edges.
-In addition, we focus our results on destructive perturbations, i.e., the edge deletion problem.
-Nevertheless, we will investigate in the next Section C.1 the edge-insertion case (i.e., edge insertion) as an analogous inverse case of edge deletion.
-This will become most relevant in the proof of Theorem C.1.
-When taking the difference $C(G)-C(G\backslash e_{i})$ by itself we will refer to it as the graph (dis)similarity between graph $G$ and $G\backslash e_{i}$.
-$I(G,e_{i})$ applied to graphs suggests a similarity “distance” between graphs based on algorithmic information content (in [63], we show that this similarity measure can classify networks by the family they belong to, differentiating variant synthetic and natural network topologies similar to graph motifs, as shown in [49]).
-In the description of the algorithm to be presented in Section B.8, replacing the underlying methods to approximate the algorithmic information content (which is the theoretically optimal measure of information content) by, e.g., Shannon entropy or lossless compression algorithms represents special cases of the more general algorithm based on algorithmic complexity, and thus it covers all these less powerful cases.
-The idea of a dynamic study/calculus of the (possible) changes that can be wrought upon an object to evaluate the contribution of each of its components for different purposes was introduced in [72, 71], and here we extend these ideas to the area of data/network dimension reduction.
-Note that MILS can actually be applied to any general object and only to graphs or networks.
-Let $\mathbb{U}$ be an object (possibly multidimensional, like those in [11, 12]) and $I(\mathbb{U},u_{i})$ be the information difference of an element $u_{i}\in\mathbb{U}$ to $\mathbb{U}$, defined by $I(\mathbb{U},u_{i})=C(\mathbb{U})-C(\mathbb{U}\backslash u_{i})$ [60].
-If $u_{i}$ is a neutral information element, then by definition it does not contribute to the algorithmic information content of $\mathbb{U}$, and so $I(\mathbb{U},u_{i})\simeq\log|\mathbb{U}|$, which indicates that $u_{i}$ is part of the dynamical causal path of $\mathbb{U}$ and $\mathbb{U}$ can regenerate $u_{i}$ by its normal dynamical course.
-(See Theorem C.8 for an investigation of neutral elements in graphs in general, and the existence of maximally neutral elements).
-In general, however, if there is no element $u_{i}$ such that $I(\mathbb{U},u_{i})\simeq\log|\mathbb{U}|$—which may be the case found in practice due to the semi-computability of algorithmic complexity—, there is always the most neutral element (possibly not unique, in case which heuristics such as InfoRank in Algorithm 2 can deal with) in $\mathbb{U}$ to be removed in the application of MILS.
-Moreover, for the most neutral element $u_{i}$, we will have it that
-$\left|I(\mathbb{U},u_{i})\right|=\min\{|{C(\mathbb{U})-C(\mathbb{U}\backslash u_{i})}|\sim\log|\mathbb{U}|)\}\text{ ,}$
-that is, the element with an information contribution closest to $\log|\mathbb{U}|$. This result can be directly obtained by an analogous argument to those in Section C.1.
-B.8 An algorithm for minimal information loss selection (MILS) and its universal decision-probability tree diagram
-MILS is an unsupervised and mostly parameter-free algorithm, i.e., asymptotically independent of model or domain, as it does not need to be instructed or designed to preserve any particular property, and maximises the preservation of all computable elements that contribute to the algorithmic information content of the data.
-Let $G$ be a graph and $\left|E\left(G\right)\right|$ denote its number of edges.
-Ideally, MILS seeks to reduce $G$ to a graph on $N<\left|E\left(G\right)\right|$ edges so that the loss of information is minimized.
-It does this by calculating the information differences between $G$ and $G\backslash S$, where $\left|S\right|\geq 1$, and then finding the subset $F\subset E(G)$ such that $I(G,F)\leq I(G,S)$ for all non-empty proper subsets of edges $S\subset E(G)$, repeating this task with $G\coloneqq G\backslash F$ until the target size is reached.
-The algorithm’s time complexity is clearly in $O(exp)$ because of how MILS performs searches and deletions across all subsets of edges, but significant improvements to this bound are still possible, as we will show in this article.
-One can easily modify the following algorithms in order to deal with multiple edge deletions at once, as in $I(G,F)$ to the above, but possibly at the expense of much more computational resources.
-A more efficient but suboptimal version of such an exponential-time MILS iterates over single elements (in the case, edges) or singletons.
-Algorithm 3 performs such a sequential deletions of edges, and then removes the edge that contributes less to the information content of the graph before moving to the next step.
-When $e$ is such that $\left|I(G,e)\right|$ is minimal, we call it a neutral information edge [72] because it is the edge that contributes less to the information content of $G$ (in particular, it minimizes information loss or the introduction of spurious information into the network according to the information difference when removed from the original network).
-The pseudocode in Algorithm 3 assumes that there is a unique such $e$, which may not necessarily be the case.
-Algorithm 2 solves this problem by performing simultaneous perturbations on all edges with an information contribution given by those in minLoss, which contains the neutral information edges.
-To this end, Algorithm 2 also introduces InfoRank, a method that produces a ranking of $e_{1},\dots,e_{|E(G)|}$ from least informative to most informative edge, i.e., a list of edges sorted in increasing order by their information contribution to $G$.
-Therefore, note that unlike minLoss in Algorithm 3, the set minLoss in Algorithm 2 can have more than one edge.
-This ranking facilitates the search for the (most) neutral elements of the system (see Section C.3), which in turn helps MILS preserve the components that maximise the information content of the resulting object.
-(As described in [72], note that Algorithms 3 and 2 may be applied, mutatis mutandis, to nodes or to any data element of an object or dataset, e.g.
-a pixel of an image, or a row or column in a spreadsheet).
-Algorithm 2 Minimal Information Loss Selection (MILS)
-1:function InfoRank($G$)
-2:     $\textit{informationLoss}\leftarrow\varnothing$
-3:      // for each edge $e_{i}$
-4:     for $e_{i}\in E(G)$ do
-5:          // store information contribution into informationLoss
-6:         $\textit{informationLoss}\leftarrow\textit{informationLoss}\cup\{I(G,e_{i})\}$
-7:     end for
-8:     sort informationLoss in increasing order
-9:      // return information rank
-10:     return informationLoss
-11:end function
-12:function MILS($G,N$), $1\leq N\leq|E(G)|$
-13:     while $|E(G)|>N$ do
-14:          // calculate minimal loss across all edges
-[…]
-16:          // remove all candidate edges from $G$
-17:         $\textit{G}\leftarrow G\backslash\{e_{i}\in E(G):I(G,e_{i})=\textit{minLoss}\}$
-18:     end while
-19:     return $G$
-20:end function
-Algorithm 3 MILS with sequential perturbation by deletion on edges
-1:function Perturb($G$)
-2:     $\textit{informationLoss}\leftarrow\varnothing$
-3:      // for each edge $e$
-4:     for $e\in E(G)$ do
-5:          // store information contribution into informationLoss
-6:         $\textit{informationLoss}\leftarrow\textit{informationLoss}\cup\{I(G,e)\}$
-7:     end for
-8:      // calculate minimal information loss across all edges
-9:     $\textit{minLoss}\leftarrow\min({informationLoss})$
-10:     for $e\in E(G)$ do
-11:          // if $e$ is a neutral edge
-12:         if $I(G,e)=\textit{minLoss}$ then
-13:               // remove $e$
-14:              return $G\backslash e$
-15:         end if
-16:     end for
-17:end function
-18:function Sparsify($G,N$), $1\leq N\leq|E(G)|$
-19:      // until reaching target size
-20:     while $|E(G)|>N$ do
-21:          // perturb edges in $G$
-22:         $G\leftarrow\textsc{Perturb}(G)$
-23:     end while
-24:     return $G$
-25:end function
-Fig.
-11 presents three illustrative examples of Algorithm 2 returning the same graph (with $N\leq 2$ present edges) as output.
-The topmost matrices in columns A, B, and C display the adjacency matrices of distinct directed graphs (without self-loops), all of them with four vertices.
-These graphs are subjected to the edge deletions that Algorithm 2 performs so as to generate its final output.
-The edge deletions occur in each iteration $j$ of [Algorithm 2, Step 13] which is indicated by the downward blue arrows (except for the bottommost blue arrows), while the subsequent red squares (i.e., the matrix entries) indicate where the deletions occurred in the previous step.
-The bottommost blue arrows in each Fig.
-11A, 11B, and 11C refer to [Algorithm 2, Step 19] after which the final output of the algorithm is returned, given an adjacency matrix of a graph $G$, such as those three topmost $4\times 4$ matrices, as input.
-The input adjacency matrix in Fig.
-11A corresponds exactly (by concatenating its rows) to the $16$-bit-length substring that starts at the $17$-th bit of the first $64$ bits of the halting probability calculated for a particular Turing machine in [23].
-Notice that this substring was chosen for illustrative purposes so that: its string is proven to be algorithmically random (i.e., incompressible); it starts at the next bit after the first $4^{2}$ bits of its original string, where $16$ is equal to the length of the substring itself and, therefore, its starting position at the original string neither adds nor subtracts information from the algorithmic complexity of the substring; the resulting diagonal is empty, which corresponds to an adjacency matrix of a graph without self-loops; and the number of present edges is almost the same as absent edges, as it is expected to occur in sufficiently long segments of an incompressible number.
-Since the input graph in Fig.
-11A is (algorithmically) random, we can assume for illustrative purposes that a $16$-bit-length random string is sufficiently long in comparison to a reference universal machine.
-In this scenario, every edge becomes non-redundant.
-Then, no edge in the (hierarchical) list given by InfoRank will be as neutral as the other ones.
-Therefore, minLoss will contain only one edge at the time in each iteration $j$ of [Algorithm 2, Step 13], resulting in $3$ iterations of deletions necessary to reach the desired target of $N\leq 2$ edges presented in Fig.
-11A.
-As the adjacency matrix given as input becomes less random, one should expect that the number of iterations can be reduced.
-Indeed, this is what happens in Fig.
-11B and 11C.
-Fig.
-11B presents a slight variation of the input matrix in Fig.
-11A.
-Unlike in the latter case, the second row is a repetition of the first row, that is, the matrix corresponds to a sequence in which the first $4$ bits are repeated and the remaining $8$ bits equals to those of the input matrix in Fig.
-11A.
-For illustrative purposes, we can assume that the algorithmic complexity of the program that repeats a given $4$-bit-length string is sufficiently small in comparison to the algorithmic complexity of the $16$-bit-length string that corresponds to the random input matrix of Fig.
-11A.
-In this case, we will have that InfoRank will consider the information difference resulting from the deletion of the edge (i.e., the deletion of the second black square) in the second row within the same ballpark of the information difference resulting from the deletion of the edge in the first row.
-In other words, the first edge is a neutral information edge as much as the second edge is.
-Therefore, both edges will be deleted in only one iteration step, namely $j=1$.
-Notice that the subsequent remaining $8$ bits are exactly those of Fig.
-11A, subsequence which is random.
-Then, one iteration will be required for each edge deletion.
-Since $N\leq 2$ was the input parameter of Algorithm 2, just one extra black square will have to be deleted in step $j=2$ before the algorithm returns the output.
-Fig.
-11C presents one of the most redundant cases in which all oriented edges are present, that is, the input is an adjacency matrix of a complete (directed) graph.
-In accordance with the illustrative assumptions for Fig.
-11A and Fig.
-11B, we can also assume that the algorithmic complexity of the program that repeats every subsequent bit other than the ones in the diagonal of the matrix is sufficiently small.
-Notice that the last $3$ bits in the last row corresponds to a subsequence of a random sequence of Fig.
-11A.
-Therefore, except for the last two edges in the last row—two edges which belong to the last $3$ bits in the last row corresponding to a subsequence of a random sequence used in cases A and B—, every edge will be considered neutral in InfoRank; and at the first iteration step ($j=1$), all of these will belong to minLoss.
-Figure 11: Three examples of edge deletions occurring in each iteration $j$ of [Algorithm 2, Step 13], indicated by the downward blue arrows (except for the bottommost blue arrows), while the subsequent red squares indicate where the deletions occurred in the previous step.
-The bottommost blue arrows in each Fig.
-11A, 11B, and 11C refer to [Algorithm 2, Step 19] after which the final output of the algorithm is returned, given an adjacency matrix of a graph $G$ (such as anyone of the three topmost $4\times 4$ matrices) as input.
-The input adjacency matrix in Fig.
-11A corresponds exactly to a substring of the halting probability as in [23].
-Mixing both redundant and incompressible components, the second row of the input matrix in Fig.
-11B is a repetition of the first row, and the remaining $8$ bits equals to those of the input matrix in Fig.
-11A.
-Fig.
-11C presents one of the most redundant cases in which all oriented edges are present, that is, the input is an adjacency matrix of a complete (directed) graph.
-In Section C.3 we will show that Algorithm 2 is deterministic, polynomially time bounded, and describes a criterion to select and remove the neutral elements of an object.
-We employ this more efficient version in all our experiments, and even in this limited form the procedure excels at preserving important characteristics of the networks under study.
-See Theorem C.11.
-B.8.1 A decision-probability tree diagram for MILS
-The tree diagram in Fig. 12 illustrates the decision process involved in each iteration of Algorithm 2 that results in the next size-reduced graph.
-MILS Tree Diagram
-Figure 12: Tree Diagram for Algorithm 2 with the sequence of graphs with their respective subsets of edges reduced (as selected in the iterations $j$ of [Algorithm 2, Step 13]) highlighted in the color blue.
-So, one has that $F_{i_{1}}=\textit{minLoss}=\min\left(\textsc{InfoRank}\left(G\right)\right)$, $F_{i_{2}}=\textit{minLoss}=\min\left(\textsc{InfoRank}\left(G\setminus F_{i_{1}}\right)\right)$, and so on, which continues up until the final graph $G\setminus F_{i_{h}}$ is found by the algorithm.
-This process is represented by the blue path $\left(G,G\setminus F_{i_{1}},G\setminus F_{i_{2}},\dots,G\setminus F_{i_{h}}\right)$ from the root $G$ to the leaf $G\setminus F_{i_{h}}$ in $h$ iterations.
-Notice that from each parent vertex/graph, there are as many children as non-empty subsets of (present) edges in the parent.
-Fig.
-12 illustrates a speed-up both globally (since the number of iterations $j\leq h$ is always upper bounded by $\mathbf{O}\left(\left|E\left(G\right)\right|-N\right)$) and locally (since from each parent, Algorithm 2 does not have to search for the best candidate across all possible subsets of edges of the parent) achieved by Algorithm 2 from the decision process that this tree diagram represents.
-Each child vertex represents a graph resulting from the deletion of one or more edges in the parent vertex (i.e., a graph before the destructive perturbations on the edges).
-In each descending step starting from the root at the top, as the height of the tree downward increases, a subset of (present) edges is deleted from the graph at the lower level, that is, a deletion from the parent vertex/graph closer to the root vertex.
-Notice that from each parent vertex, there can be as many child vertices/graphs as the size of the power set $\mathcal{P}\left(E(G)\right)$ of the edges in the graph represented by the parent vertex.
-The child vertex chosen by Algorithm 2 is the one whose neutral information edges (according to InfoRank) were deleted.
-In this manner, the iteration process in [Algorithm 2, Step 13] is represented by the rooted path highlighted in the color blue in the Fig.
+MILS is highly context-sensitive and adapts to different kind of information such as images.
+For example, it handles in different ways images as seen in Fig.
 12.
-A subset of edges that has neutral information with respect to the parent graph is deleted from the parent graph in order to generate the child graph in this path.
-InfoRank allows Algorithm 2 to speed up the selection of the child vertex that (in general, suboptimally) minimises the information difference with respect to the parent vertex, not having to search over all the possible subsets of present edges in the parent vertex.
-This constitutes a local speed-up between each parent vertex and its children in Fig.
-12, in particular replacing a search over an exponential-size set of subsets with a search over a list (of edges), therefore quadratic with the number of vertices/nodes in the root graph.
-For the illustrative purposes of Fig.
-12, one can take the arrow labelings in the form $i_{j}$ so as to correspond to the exact sequence of subsets of edges that Algorithm 2 chose to delete from the output of InfoRank.
-(Notice that in practice one would in general start with a fixed labeling, and thereafter track the rooted path in the tree that corresponds to the exact sequence of subsets of edges that Algorithm 2 ended up choosing to delete).
-Thus, since Algorithm 2 starts by deleting the first element in the list given by InfoRank, and possibly continuing the deletions until subsequently exhausting every neutral information edge in the ordered list, one will have only $\left|E\left(G\right)\right|$ possible choices of subsets of edges to delete from the output of InfoRank.
-This is displayed in Fig.
-12 by the search interval comprising the set $\left\{x\middle|i_{j}-k\leq x\leq i_{j}+k\right\}$, for each iteration $j$ with $2k+1\coloneqq\left|E\left(G^{\prime}\right)\right|$,888 Or $12k+1\coloneqq\left|E\left(G^{\prime}\right)\right|+$ 1, if $\left|E\left(G^{\prime}\right)\right|$ is an even number in which case one can assume $i_{j}-k$ to correspond to the empty subset just to keep the symmetry of the interval around $i_{j}$ for illustrative purposes.
-where $G^{\prime}$ is any graph.
-By illustrating the speed-up aforementioned in this paragraph, no matter how many subsets of edges a parent graph has, the width of the interval between $i_{j}-k$ and $i_{j}+k$ remains999 Except for including or not the empty set to account for the eventual even value of $\left|E\left(G\right)\right|$.
-$\left|E\left(G^{\prime}\right)\right|\leq\left|E\left(G\right)\right|$ across all the iterations $j$ for every child graph $G^{\prime}$ until the total height $h$ of the tree rooted in graph $G$ is reached by Algorithm 2.
-The single-thread process (between $j=0$ and $j=h$) highlighted in blue in Fig.
-12 continues until the leaf vertex, i.e., the graph with $\leq N$ edges is achieved by Algorithm 2.
-This constitutes a global speed-up between the root vertex and the final output of Algorithm 2, particularly replacing a search over all possible rooted paths with a recursive iteration in [Algorithm 2, Step 13] whose maximum number of iterations is always upper bounded by $h\leq\left|E\left(G\right)\right|-N+1$, therefore quadratic with the number of vertices in the root graph.
-This makes the height of the tree always much shorter in comparison to its width.
-It is also important to notice that the decision tree diagram (i.e., the one for the minimisation of the information difference between the parent and the child vertex) in Fig.
-12 can be subsumed into a probability tree diagram.
-Each path from the root to a leaf has an associate joint probability of a sequence of independent events as defined in Equation (23), and each event corresponds to a destructive perturbation on a subset of edges.
-The rooted path highlighted in the color blue in the Fig.
-12 illustrates the path to the final leaf vertex (i.e., the vertex representing the outcome of Algorithm 2) that also maximises the (algorithmic) probability of edge deletions, while each edge in this path corresponds to an iteration in [Algorithm 2, Step 13].
-Now, one can demonstrate that the MILS tree diagram in Fig.
-12 shows that a decision process of minimisation of information differences is subsumed into a maximization of (universal algorithmic) probabilities.
-In turn, it evinces the generality of our methods (based on the foundational contributions of algorithmic information theory).
-This follows from the algorithmic coding theorem [26, 24, 43, 34] as in Equation (12).
-To this end, from Definition B.12, first note that from each parent graph $G$ (which is the graph before the destructive perturbations take place) and each child graph $G\setminus F$ (the one after a destructive perturbation on the subset $F$ of present edges occurs) there is a unique value
-$\left|I(G,F)\right|=\left|C(G)-C(G\setminus F)\right|\text{ .}$ (17)
-Let $\mathfrak{T}$ be a tree in which each arrow (labeled $i_{j}$) from $G$ to $G\setminus F$ has a probability measured by $\mathbf{P}$ as defined by
-$\mathbf{P}\left[G\to G\setminus F\right]=c_{G}\,\left(\frac{1}{2^{\left|C(G)-C(G\setminus F)\right|}}\right)\text{ ,}$ (18)
-where $c_{G}$ is a normalising constant that only depends on $G$ so that
-$c_{G}\coloneqq\sum\limits_{F\in\mathcal{P}\left(E(G)\right)}\frac{1}{2^{\left|C(G)-C(G\setminus F)\right|}}\text{ .}$ (19)
-Let $\mathfrak{p}_{F}^{-}$ denote the destructive algorithmic perturbation [13] that generates $G\setminus F$ from $G$, and $\mathfrak{p}_{F}^{+}$ the constructive perturbation that realises the inverse operation.
-Therefore, from the results in algorithmic information theory (AIT) discussed in Section B.6, one can straightforwardly demonstrate the inequality
-$\frac{c_{G}}{2^{\max\left\{C(\mathfrak{p}_{F}^{-}),C(\mathfrak{p}_{F}^{+})\right\}}}\leq\mathbf{P}\left[G\to G\setminus F\right]\leq 1\text{ .}$ (20)
-Hence, one also has that
-$c_{G}\,\left(\sum\limits_{F\in\mathcal{P}\left(E(G)\right)}\frac{1}{2^{\max\left\{C(\mathfrak{p}_{F}^{-}),C(\mathfrak{p}_{F}^{+})\right\}}}\right)\leq\sum\limits_{F\in\mathcal{P}\left(E(G)\right)}\mathbf{P}\left[G\to G\setminus F\right]=1$ (21)
-holds. Thus, each rooted path $G\to^{*}G_{\left(h,i_{h}\right)}$ (from the root $G$ to a leaf $G_{\left(h,i_{h}\right)}$) is a sequence
-$\left(G,G_{\left(1,i_{1}\right)},G_{\left(2,i_{2}\right)},\cdots,G_{\left(h,i_{h}\right)}\right)$ (22)
-of graphs with an associate joint probability (of independent events, i.e., independent perturbations $\mathfrak{p}_{F_{i_{j}}}^{-}$) given by
-$\mathbf{P}\left[G\to^{*}G_{\left(h,i_{h}\right)}\right]=\prod\limits_{\scriptsize\begin{array}[]{c}1\leq j\leq h\text{,}\\ \left|\left\{i_{k}\middle|k=j\right\}\right|=\sum\limits_{i_{j-1}}\left|\mathcal{P}\left(E\left(G_{\left(j-1,i_{j-1}\right)}\right)\right)\right|-1\end{array}}\mathbf{P}\left[G_{\left(j-1,i_{j-1}\right)}\to G_{\left(j,i_{j}\right)}\right]\text{ ,}$ (23)
-where $G_{\left(0,0\right)}\coloneqq G$ , $G_{\left(j,i_{j}\right)}\coloneqq G\setminus F_{i_{j}}$, |$\left|\left\{i_{j}\middle|1\leq j\leq h\right\}\right|=\sum\limits_{1\leq k\leq h}\left|\left\{i_{k}\middle|k=j\right\}\right|$, and $\left|E\left(G_{\left(h,i_{h}\right)}\right)\right|=\left|E\left(G\setminus F_{i_{h}}\right)\right|\leq N$.
-Notice that if $G_{\left(j-1,i_{j-1}\right)}\to G_{\left(j,i_{j}\right)}$ is a arrow in the tree, then $F_{i_{j}}\subseteq F_{i_{j-1}}$.
-Now, one can employ the algorithmic coding theorem (as in Equation (12)) to demonstrate in Theorem B.1 that Equation (23) is an upper bound (except for an independent constant) for any other arbitrarily chosen (computable) probability (semi-)measure $\mathbf{P}^{\prime}$.
-In other words, for any arbitrary $\mathbf{P}^{\prime}$ capable of calculating the probability of any path in the MILS tree diagram—even if such a probability is not defined upon the assumption of independence of the subsequent perturbations like $\mathbf{P}$ is—, if a certain path maximises $\mathbf{P}^{\prime}$, then $\mathbf{P}$ in Equation (23) is also maximised.
-This results show that a decision tree diagram becomes subsumed into a probability tree diagram in the theoretical optimal limit no matter which probability one chooses to be maximised, which is a result that highlights the robustness of the methods employed in the present paper.
-In addition, Theorem B.1 also shows that the independence of events assumed in $\mathbf{P}$ does not make one loose any generality in comparison to other probability measures that one tries to maximise.
-Notice that the algorithmic coding theorem promptly demonstrates a universal/maximal probability semimeasure for minimising either $C(G)$ or $C\left(G\setminus F\,\middle|\,G\right)$.
-Theorem B.1 then contributes by showing that this does not only holds for the unconditional and conditional classical cases directly, but also for the minimization of Equation (17) across the whole decision tree of subsequent independent events, which in turn implies the maximisation of Equation (23) across the probability tree diagram.
-Thus, we will prove that the algorithmic-probability-based method can be extended to Equation (17); additionally, we shall prove that this applies both globally to
-$\mathbf{P}\left[G\to^{*}G_{\left(h,i_{h}\right)}\right]$
-(for a path $G\to^{*}G_{\left(h,i_{h}\right)}$ between the root and a leaf) and locally to every child vertex and tree arrow along such a path; and the assumption of independent events in Equation (23) does not imply any loss of generality in the theoretical asymptotic limit.
-This result shows the generality and theoretical robustness of the methods based on algorithmic information theory: in particular, a decision tree based on minimisation of an index value is proven to be subsumed into a maximisation of path probabilities of (subsequent independent) destructive perturbations; in addition, this holds agnostically with respect to the choice of the computation model, programming language, or another arbitrary (computable) probability (semi-)measure that one aims at maximising (whether or not one assumes independence between parent vertices and subsequent child events in such a measure).
-Theorem B.1 (Decision-Probability Tree Diagram).
-[…]
-$\mathbf{P}^{\prime}_{j}\left[G\to^{*}G\setminus F_{i_{j}}\right]\leq\epsilon\,\mathbf{P}\left[G\to^{*}G\setminus F_{i_{j}}\right]$ (24)
-holds for every graph $G\setminus F_{i_{j}}$ in the path from $G$ to $G\setminus F_{i_{h}}$, where $j\leq h$.
-Proof.
-By assumption, we have that $\mathbf{P}^{\prime}_{j}\left[G\to^{*}G\setminus F_{i_{j}}\right]$ is defined for every $i_{j}\in\left\{i_{j}\middle|1\leq j\leq h\right\}$.
-In this way, one can employ an argument by induction on $j$.
-First, we have that $\mathbf{P}^{\prime}_{1}\left[G\to G\setminus F_{i_{1}}\right]$ is well defined for every $i_{1}\in\left\{i_{k}\middle|k=1\right\}$.
-Then, from Equation (12), we know that there is a constant $\epsilon_{1}>0$ such that
-$\mathbf{P}^{\prime}_{1}\left[G\to G\setminus F_{i_{1}}\right]\leq\epsilon_{1}\mathbf{P}\left[G\to G\setminus F_{i_{1}}\right]$ (25)
-holds. Suppose now that for every $m\leq n$ and every $i_{m}\in\left\{i_{k}\middle|k=m\right\}$, there is a constant $\epsilon_{m}$ such that
-$\mathbf{P}^{\prime}_{m}\left[G\to^{*}G\setminus F_{i_{m}}\right]\leq\epsilon_{m}\mathbf{P}\left[G\to^{*}G\setminus F_{i_{m}}\right]$ (26)
-holds.
-Since $\left\{\mathbf{P}^{\prime}_{j}\right\}$ is a computable list by hypothesis and $\mathbf{P}^{\prime}_{n+1}$ is well-defined, there would be another computable measure $\mathbf{P}^{\prime\prime}_{n+1}\left[G\setminus F_{i_{n}}\to G\setminus F_{i_{n+1}}\right]$ that is well-defined for every $i_{n}$ and $i_{n+1}$ such that
-$\mathbf{P}^{\prime\prime}_{n+1}\left[G\setminus F_{i_{n}}\to G\setminus F_{i_{n+1}}\right]=c_{n}\,\frac{\mathbf{P}^{\prime}_{n+1}\left[G\to^{*}G\setminus F_{i_{n+1}}\right]}{\mathbf{P}^{\prime}_{n}\left[G\to^{*}G\setminus F_{i_{n}}\right]}$ (27)
-holds, where $c_{n}$ is a normalising constant. Therefore, from Equation (12), we would also have that there is a constant $\epsilon_{n+1}$ such that
-$\mathbf{P}^{\prime\prime}_{n+1}\left[G\setminus F_{i_{n}}\to G\setminus F_{i_{n+1}}\right]\leq\epsilon_{n+1}\mathbf{P}\left[G\setminus F_{i_{n}}\to G\setminus F_{i_{n+1}}\right]$ (28)
-holds. Finally, the desired proof then follows directly by making
-$\epsilon\coloneqq\max\left\{\epsilon_{k}\middle|1\leq k\leq h\right\}\text{ .}$
-Thus, as proven in this Section B.8.1 (see also the forthcoming Section C.3), MILS is, by design, optimal in the computability and algorithmic-information theoretic sense, and only dependent on the method for approximating algorithmic complexity in the preservation of any possible feature of interest that contributes to the (algorithmic) information content of a network such as its degree distribution and other graph-theoretic, algebraic or topological features—even including those not necessarily captured by any graph theoretic measure or classical information approach [66, 67].
-B.8.2 PyMILS and variations of MILS
-Algorithm 4 Python implementation of the Minimal Information Loss Selection algorithm (PyMILS)
-1:function pymils($image$)
-2:     while $size(image)>size(image)_{min}$ do
-3:         $bdm_{initial}=BDM(image)$
-4:         Randomly select one of the following: $(rows,columns)$
-5:         For some chunk size $c$, randomly select clustered groups of size $c$ of either rows or columns
-6:         for each clustered group, do
-7:              remove that group to get $image_{step}$
-8:              measure $bdm_{step}=BDM(image_{step})$
-9:              replace that group
-10:         end for
-11:         Remove the clustered group from $image$ that would result in a $bdm_{step}$ closest to $bdm_{initial}$.
-12:     end whilereturn final image
-13:end function
-We also present a variation of Algorithm 2 that instead resorts in heuristically allowing a random selection of elements when they have the same information value.
-Notice that unlike Algorithms 3 and 2 that always tries to minimise information loss in a finer-grained fashion, the random-sampling variation in Algorithm 4 blindly chooses the elements to be removed at each iteration so as to pick the one that minimises the information loss at the end.
-Therefore, its time complexity overhead falls under the percentage of samplings which is a parameter specified by the user.
-Given a single input image, the pseudocode for pymils is described in Algorithm 4.
-A user can tune pymils (Algorithm 4) with the following parameters:
-  1. 1.
-    $size(image)_{min}$: Float $0-1$. Minimum final image size ratio (cannot be smaller than 4 rows or columns).
-  2. 2.
-    Sampling: Float $0-1$. Percent of rows/columns to randomly sample at each step.
-  3. 3.
-    Chunk size $c$: Integer. Size of neighbouring row/column group to remove at each step.
-Algorithm 4 was implemented in a Python package called pymils.
-This package is freely available and can be downloaded through GitHub.
-To test the validity of pymils, we used several different images that were converted into binary numpy arrays101010numpy is a Python package that efficiently stores and handles matrices and arrays..
-To show the effectiveness of MILS as implemented through pymils, we first test its ability to compress images from the DataSaurus Dozen [48].
-Figure 13 shows the 13 images as binary arrays.
-Each of these images have the same summary statistics, including the x/y mean, the x/y standard deviation, and linear correlation coefficient to two decimal points.
-This set of different images has been used as a cautionary tale against trusting summary statistics alone without looking directly at the underlying dataset.
-Figure 13: The Datasaurus Dozen datasets [48] visualised as binary arrays.
-Figure 14 shows the initial complexity as approximated by BDM and the subsequent change in complexity after pymils was implemented on each image.
-From Figure 14, it is clear that the different images have different algorithmic complexity values, in contrast with summary statistics.
-In addition, pymils is able to compress these images such that the resulting image is close to the same complexity as the original image.
-Figure 14: Scatterplots showing the initial algorithmic complexity as approximated by BDM (top) and the change in complexity after pymils compressed each image to a 50% the original image size.
-We tried various other binary array images, including shapes and 1-dimensional elementary cellular automata (ECA) patterns with random initial states.
-Figure 15 shows an image of a face before and after image compression via pymils.
-In this example, pymils removed most of the redundancy from the lower half of the original image while preserving complex features in the top half of the original image.
-Figure 15: Example image of a face pasted above blank space before (top) and after (bottom) pymils, with a 50% size reduction.
-To test the ability of pymils to compress an image while preserving algorithmic complexity, we compared it to two other methods: randomly removing rows/columns and PCA.
-We randomly removed chunks of rows or columns of the same size from images without measuring the resulting complexity values.
-The PCA analysis was done with two components (for binary values) on each image and then mapping the resulting values onto an image of the same resulting size of pymils (50% the original size).
-Results from this comparison are shown in Figure 16.
-As a result, pymils performs much better on average at preserving algorithmic complexity ($\Delta BDM$) during image compression.
-Both random deletion and PCA can introduce additional noise in the image, which results in a negative $\Delta BDM$ value.
-Figure 16: Boxplots comparing pymils with random row/column deletion and PCA over various test images (top) and Elementary Cellular Automata (ECA) images (bottom).
-The fact that for BDM the average is closer to zero and the variance (box size) is shorter and more compact, means that the procedure is succeeding at preserving the (algorithmic) information content of the object after dimension reduction, while other methods destroy the information content and, therefore, the underlying program complexity (i.e., its causal structure).
-In addition to testing whole images, the initial and resulting BDM values were tracked for the four image quadrants in each image.
-Results from Figure 17 suggests pymils performs relatively similar across each image quadrant.
-Figure 17: Boxplots comparing pymils with random row/column deletion and PCA over various test image quadrants (top) and ECA image quadrants (bottom).
-In the case of the general images, quadrant averages after reduction had little effect, but for ECAs and their linear evolution, quadrants have different impact depending on the complexity of the region hence also sensitive to different features.
-In future research, other variations of Algorithm 2, like Algorithm 4 may be proposed for example in order to improve the sheer cubic dependence on the network size.
-In this direction, such variations can stem from improvements on the embedded sorting algorithm and its usage within the loop in [Algorithm 2, Step 13].
-At the expense of increasing the running time, another future variation may employ a breadth-first search or depth-first search across the edges listed in increasing order of information contribution to the parent graph.
-This would allow the consideration over a larger quantity of $h$-hop-length rooted paths.
-B.9 MILS and other dimensionality reduction methods
-Other dimension reduction methods such as UMAP, AE, VAE, t-SNE, PHATE, and others (for surveys see [21, 35, 31]) are, in a deep fundamental sense, special cases of MILS.
-This is because MILS implements a universal algorithm that is ultimately a generalisation of all other computable methods based on probability distributions, algebra or topology.
-For example, t-SNE and UMAP, are heavily based on building probability distributions thus they will suffer a similar fate to those reported in [66].
-In any event, the ergodicity of the algorithmic complexity approximation implemented by the coding theorem method (CTM) over the software space implies that the BDM will eventually encompass all this space if enough computational resources are expended.
-In other words, even the most sophisticated algorithms implementing algebraic or topological feature selection beyond those implementing classical probability approaches will be eventually instantiated by CTM (if the algebraic or topological features are recursively enumerable) and therefore MILS can be seen as an asymptotic generalisation of all other algorithms.
-The question is simply which one may extract certain features of interest faster and which one will be less biased and more universal in the sense that features are not pre-selected (e.g.
-algebraic or topological) like when picking a method that will maximise but also bias towards some properties of the data in an embedding.
-MILS, in that sense, is the most universal and agnostic but will not necessarily overtake all others in all applications at the same rate or efficiency, which would only be the case with unbounded computational resources.
-Appendix C Theoretical results
-C.1 Algorithmic information oscillations from local perturbations
-The main idea of the proof of Theorem C.1 is to construct an algorithm for naming the exact edge subset $F\subseteq\mathbb{E}(G)$ that are being deleted (or inserted).
-Then, we show that, once this enumeration is calculated, this general algorithm employs (or reversibly employs) the perturbation $\mathfrak{p}_{F}^{-}$ (see [13] for a formalisation of algorithmic perturbations) in order to generate $G\setminus F$ from $G$ (or a perturbation $\mathfrak{p}_{F}^{+}$ for $G$ from $G\setminus F$).
-Note that the proof is made for any possible subset $F\subseteq\left|\mathbb{E}_{c}(G)\right|$ of perturbed edges, where $\left|F\right|=r\leq\left|\mathbb{E}_{c}(G)\right|$.
-Therefore, if one takes $\left|F\right|=1$, Theorem C.1 immediately applies to single edge deletions (or insertions) as in Definiton B.12.
-In addition, the reader is invited to note that the proof of Theorem C.1 holds for the classical case of graphs, but it can be extended to directed graphs.
-Theorem C.1.
-Let $G=(V,E)$ be an arbitrary simple graph with $E\subseteq\mathbb{E}_{c}(G)$ and $\left|V\right|=N\in\mathbb{N}$.
-Let $F\subseteq\mathbb{E}_{c}(G)$ be an arbitrary subset of edges, where $\left|F\right|=r\geq 1$.
-Then,
-$\left|I(G,F)\right|\leq 2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(N)\right))\text{ ,}$
-if edges in $F\neq\emptyset$, where $F\subseteq E(G)$, are being deleted, or
-$\left|I^{-1}(G,F)\right|\leq 2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(N)\right))\text{ ,}$
-if edges in $F\neq\emptyset$, where $F\subseteq\mathbb{E}_{c}(G)\setminus E(G)$, are being inserted.
-Proof.
-We will split the proof in two cases. First, we will prove
-$C(G)\leq C(G\setminus F)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(N)\right))$ (29)
-and then
-$C(G\setminus F)\leq C(G)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(N)\right))\text{ .}$ (30)
-First, let $F\subseteq\mathbb{E}_{c}(G)$ be an arbitrary subset of edges.
-Let $p_{1}\in\{0,1\}^{*}$ be a string that represents the algorithm running on a prefix universal Turing machine $\mathbf{U}$ such that $p_{1}$ receives $a\in\mathbf{L^{\prime}_{U}}$, $b\in\{0,1\}$, and $k,r\in\mathbb{N}$ as inputs.
-Then, $p_{1}$ calculates $\mathbf{U}\left(a\right)$ and check whether or not $\mathbf{U}\left(a\right)=A_{G^{\prime}}$ for some graph $G^{\prime}$, according to an arbitrarily chosen recursively bijective encoding of an arbitrary adjacency matrix into a bit string.
-Then, $p_{1}$ execute the following procedure:
-  • If $\mathbf{U}\left(a\right)=A_{G^{\prime}}$, then:
-    • if $b=0$ (i.e., edges were deleted), then:
-      $p_{1}$ enumerates the subsets of size $r$ in the power set $\mathcal{P}(\overline{E}(G^{\prime}))$, which comprises all the subsets of the set
-      $\overline{E}(G^{\prime})=\mathbb{E}_{c}(G^{\prime})\setminus E(G^{\prime})$
-      of absent undirected edges (i.e., the symmetric zeros) in $A_{G^{\prime}}$.
-Then, $p_{1}$ replaces in $A_{G^{\prime}}$ every 00 with $1$ that corresponds to the elements in the $k$-th subset $X_{k}\subseteq\overline{E}(G^{\prime})$ in this enumeration of subsets, where $\left|X_{k}\right|=r$.
-Finally, $p_{1}$ returns this resulting adjacency matrix of graph $G^{\prime}\cup X_{k}$.
-    • if $b=1$ (i.e., edges were inserted), then:
-      $p_{1}$ enumerates the subsets of size $r$ in the power set $\mathcal{P}(E(G^{\prime}))$, which comprises all the subsets of the set
-      $E(G^{\prime})\subseteq\mathbb{E}_{c}(G^{\prime})$
-      of present undirected edges (i.e., the symmetric $1$’s) in $A_{G^{\prime}}$.
-Then, $p_{1}$ replaces in $A_{G^{\prime}}$ every $1$ with 00 that corresponds to the elements in the $k$-th subset $X_{k}\subseteq E(G^{\prime})$ in this enumeration of subsets.
-Finally, $p_{1}$ returns this resulting adjacency matrix of graph $G^{\prime}\setminus X_{k}$.
-  • Otherwise, $p_{1}$ starts any arbitrarily chosen non-halting computation.
-Note that, since $\left(G\setminus F\right)\cup F=G$ and $\left(G\cup F\right)\setminus F=G$ and $C(p_{1})$ only varies by $\mathbf{O}(1)$ on choosing between deletions and insertions, we can study only the edge-deletion case without loss of generality.
-Also note that one can always define a recursively bijective encoding of an arbitrary adjacency matrix into a bit string.
-Therefore, if $a=(A_{G^{\prime}})^{*}$ is the shortest program that generates the adjacency matrix $A_{G^{\prime}}$ of an arbitrary graph $G^{\prime}$, then program $p_{1}$ always halts in the cases:
-  • $b=0$, $k\leq\binom{\left|\overline{E}(G^{\prime})\right|}{r}$, and $1\leq r\leq\left|\overline{E}(G^{\prime})\right|$ or;
-  • $b=1$, $k\leq\binom{\left|E(G^{\prime})\right|}{r}$, and $1\leq r\leq\left|E(G^{\prime})\right|$.
-Proof of the upper bound (29):.
-We have that there are only
-$\left|\overline{E}(G\setminus F)\right|$ $=\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G\setminus F)\right|=$ (31)
-$=\left|\mathbb{E}_{c}(G)\right|-\left|E(G\setminus F)\right|=$
-$=\frac{N^{2}-N}{2}-\left|E(G)\right|+\left|F\right|=$
-$\leq 2\left(\frac{N^{2}-N}{2}\right)=\mathbf{O}\left(N^{2}\right)$
-possible absent edges in $G\setminus F$. We know that the exact collection $F$ of edges that were deleted in $G$ can be univocally encoded in less than or equal to
-[…]
-bits. Note that
-$\binom{\left|\overline{E}(G\setminus F)\right|}{r}=\frac{\left|\overline{E}(G\setminus F)\right|!}{r!(\left|\overline{E}(G\setminus F)\right|-r)!}<2^{\left|\overline{E}(G\setminus F)\right|}\text{ .}$ (33)
-In addition, we have from Equation (31) that, if $r=\left|F\right|$, then
-$\frac{\left|\overline{E}(G\setminus F)\right|!}{r!(\left|\overline{E}(G\setminus F)\right|-r)!}=\\ \frac{\left(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|+\left|F\right|\right)!}{\left|F\right|!(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|+\left|F\right|-\left|F\right|)!}=\\ =\frac{\left(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|+\left|F\right|\right)!}{\left|F\right|!(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|)!}=\\ =\frac{\left(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|+1\right)\cdots\left(\left|\mathbb{E}_{c}(G\setminus F)\right|-\left|E(G)\right|+\left|F\right|\right)}{\left|F\right|\cdots 1}\leq\\ \leq\left(\left|\mathbb{E}_{c}(G)\right|-\left|E(G)\right|+\left|F\right|\right)^{\left|F\right|}\leq\\ \leq\left(2\left(\frac{N^{2}-N}{2}\right)\right)^{\left|F\right|}\leq\left(N^{2}\right)^{\left|F\right|}\text{ .}$ (34)
-Now, let $X_{k}$ corresponds to the exact subset $F$ of edges, $a=(A_{G\setminus F})^{*}$, $b=0$, and $r=\left|F\right|$.
-Note that, by supposition and construction, $\left|F\right|\leq\left|\overline{E}(G\setminus F)\right|$ and $\left|F\right|\geq 1$.
-Therefore, by the minimality of the prefix algorithmic complexity [43, 34, 26], the construction of program $p_{1}$, and Equations (32), (33), and (34) we will have that
-$C(G)\leq l\left(\left<a,b,k,r,p_{1}\right>\right)\leq$ (35)
-$\leq C(G\setminus F)+\log_{2}\left(\binom{\left|\overline{E}(G\setminus F)\right|}{r}\right)+\\ \mathbf{O}\left(\log_{2}\left(\log_{2}\left(\binom{\left|\overline{E}(G\setminus F)\right|}{r}\right)\right)\right)+\\ \log_{2}\left(\left|F\right|\right)+\mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))+\mathbf{O}(1)\leq$ (36)
-$\leq C(G\setminus F)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(2\left|F\right|\log_{2}\left(N\right)\right)\right)+\log_{2}\left(\left|F\right|\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))+\mathbf{O}(1)\leq$ (37)
-$\leq C(G\setminus F)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\mathbf{O}(\log_{2}\left(\log_{2}(N)\right))$ (38)
-Proof of the lower bound (30):.
-We have that there are only
-$\left|E(G)\right|\leq\frac{N^{2}-N}{2}=\mathbf{O}\left(N^{2}\right)$ (39)
-possible present edges in $G$. We know that the exact collection $F$ of edges that were deleted in $G$ (or inserted in $G\setminus F$) can be univocally encoded in
-$\log_{2}\left(\binom{\left|E(G)\right|}{r}\right)+\mathbf{O}\left(\log_{2}\left(\log_{2}\left(\binom{\left|E(G)\right|}{r}\right)\right)\right)$ (40)
-bits. From Equation (39), we also analogously have as in Equation (31) that, if $r=\left|F\right|$, then
-$\binom{\left|E(G)\right|}{r}$ $=\frac{\left|E(G)\right|!}{\left|F\right|!(\left|E(G)\right|-\left|F\right|)!}$ $=$ (41)
-$=\frac{\left(\left|E(G)\right|-\left|F\right|+1\right)\cdots\left(\left|E(G)\right|\right)}{\left|F\right|\cdots 1}\leq$
-$\leq\left(\left|E(G)\right|\right)^{\left|F\right|}\leq$
-$\leq\left(\frac{N^{2}-N}{2}\right)^{\left|F\right|}\leq\left(N^{2}\right)^{\left|F\right|}\text{ .}$
-Now, let $X_{k}$ corresponds to the exact subset $F$ of edges, $a=(A_{G})^{*}$, $b=1$, and $r=\left|F\right|$.
-Note that, by supposition, $\left|F\right|\leq\left|E(G)\right|$ and $\left|F\right|\geq 1$.
-Therefore, by the minimality of the prefix algorithmic complexity [43, 34, 26], the construction of program $p_{1}$, and Equations (40) and (41), we will have that
-$C(G\setminus F)\leq l\left(\left<a,b,k,r,p_{1}\right>\right)\leq$ (42)
-$\leq C(G)+\log_{2}\left(\binom{\left|E(G)\right|}{r}\right)+\mathbf{O}\left(\log_{2}\left(\log_{2}\left(\binom{\left|E(G)\right|}{r}\right)\right)\right)+\\ \log_{2}\left(\left|F\right|\right)+\mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))+\mathbf{O}(1)\leq$ (43)
-$\leq C(G)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(2\left|F\right|\log_{2}\left(N\right)\right)\right)+\log_{2}\left(\left|F\right|\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))+\mathbf{O}(1)\leq$
-$\leq C(G)+2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\mathbf{O}(\log_{2}\left(\log_{2}(N)\right))$
-If one wants tighter lower and upper bounds for multiple edge deletions as a function of the edge density (i.e., the size of $\left|E(G)\right|$ in comparison to $\frac{N^{2}-N}{2}$), then one can re-write the statement of Theorem C.1 and achieve:
-$-\left|F\right|\log_{2}\left(\left|E(G)\right|\right)-\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)-\\ \mathbf{O}\left(\log_{2}\left(\log_{2}\left(\left|E(G)\right|\right)\right)\right)-\mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))-\mathbf{O}(1)\\ \leq C(G)-C(G\setminus F)\leq\\ \left|F\right|\log_{2}\left(\frac{N^{2}-N}{2}-\left|E(G)\right|+\left|F\right|\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\\ +\mathbf{O}\left(\log_{2}\left(\log_{2}\left(\frac{N^{2}-N}{2}-\left|E(G)\right|+\left|F\right|\right)\right)\right)+\\ \mathbf{O}(\log_{2}\left(\log_{2}(\left|F\right|)\right))+\mathbf{O}(1)$ (44)
-Note that, if the number of perturbations are about the same order of the present edges in $G$, then Equation (44) would tell us that the algorithmic information loss is allowed to be logarithmically stronger than the algorithmic information gain, should $\left|E(G)\right|\ll\frac{N^{2}-N}{2}$ holds.
-As one might intuitively expect, when the frequency of edge occurrences is nearly equidistributed (in particular, when the graph is incompressible), we can also retrieve from algorithmic information theory [26, 43, 34, 24] some of the particular cases in which the loss and gain of algorithmic information are both dominated by a slightly more-than-linear factor of the number of destructive perturbations:
-Corollary C.2.
-Let $G=(V,E)$ be an arbitrary $\mathbf{O}(1)$-K-random simple graph (as in Definition B.11) with $E\subseteq\mathbb{E}_{c}(G)$ and $\left|V\right|=N$, where $N\in\mathbb{N}$ is large enough such that
-$\left|\left|E(G)\right|-\frac{N^{2}-N}{4}\right|\leq\mathbf{O}\left(\sqrt{\frac{N^{2}-N}{2}\log_{2}\left(\frac{N^{2}-N}{2}\right)}\right)$ (45)
-holds. Let $F\subseteq\mathbb{E}_{c}(G)$ be an arbitrary subset of edges, where $\left|F\right|=r\geq 1$. Then,
-[…]
-$C(G)-C(G\setminus F)\leq$
-[…]
-if edges in $F\neq\emptyset$, where $F\subseteq E(G)$, are being deleted.
-Proof.
-From [9], as stated in Section B.6, we have that
-$C(E(G))=C(G)\pm\mathbf{O}(1)=C(x)\pm\mathbf{O}(1)\text{ ,}$ (47)
-where $x$ is the characteristic string of $G$. Therefore, in accordance with Definitions B.8 and B.11, we will have that
-$C(x)\geq l(x)-\mathbf{O}(1)$
-In this case, from the Borel normality of the string $x$ [43, 24], we know there is a large enough characteristic string of length $\frac{{N_{0}}^{2}-N_{0}}{2}$ such that the condition in Equation (45) is always satisfied for every $N\geq N_{0}\in\mathbb{N}$.
-Now, remember Equation (31).
-Therefore, to achieve the desired proof, just replace $\left|E(G)\right|$ with
-$\frac{N^{2}-N}{4}\pm\mathbf{O}\left(\sqrt{\frac{N^{2}-N}{2}\log_{2}\left(\frac{N^{2}-N}{2}\right)}\right)$
-into Equation (44). ∎
-One can modify Theorem C.1 (or Equation 44) and find an alternative form that does not depend on $|F|$ and edge density (which in fact corresponds to the worst-case information difference in a straightforward encoding of the set $F$ in $\sim\log_{2}\left(2^{\frac{N^{2}-N}{2}}\right)$ bits, thereby simpler than the encoding applied in Theorem C.1):
-Corollary C.3.
-Let $G=(V,E)$ be an arbitrary simple graph (as in Definition B.3) with $E\subseteq\mathbb{E}_{c}(G)$ and $\left|V\right|=N\in\mathbb{N}$.
-Let $F\subseteq\mathbb{E}_{c}(G)$ be an arbitrary subset of edges, where $\left|F\right|=r\geq 1$.
-Then,
-$\left|I(G,F)\right|=\\ \mathbf{O}(N^{2})\text{ ,}$
-if edges in $F\neq\emptyset$, where $F\subseteq E(G)$, are being deleted, or
-$\left|I^{-1}(G,F)\right|=\mathbf{O}(N^{2})\text{ ,}$
-if edges in $F\neq\emptyset$, where $F\subseteq\mathbb{E}_{c}(G)\setminus E(G)$, are being inserted.
-Proof.
-First, remember that
-$\binom{n}{k}\leq 2^{n}$ (48)
-for very $n\in\mathbb{N}$ and $0\leq k\leq n$.
-As described in the beginnig of the proof of Theorem C.1, we can focus on the edge deletion case without loss of generality, so that the edge insertion case derives in an analogous inverse manner.
-This way, we have by definition that
-$\left|F\right|\leq\left|E(G)\right|\leq\frac{N^{2}-N}{2}$ (49)
-and
-$\left|\overline{E}(G\setminus F)\right|\leq\frac{N^{2}-N}{2}\text{ .}$ (50)
-Therefore, the proof follows analogously from Theorem C.1, except for subsequentially replacing Equations (32) and (40) with
-$\left|\overline{E}(G\setminus F)\right|+\mathbf{O}\left(\log_{2}\left(\left|\overline{E}(G\setminus F)\right|\right)\right)$ (51)
-and
-$\left|E(G)\right|+\mathbf{O}\left(\log_{2}\left(\left|E(G)\right|\right)\right)\text{ ,}$ (52)
-respectively in the proof of Theorem C.1. ∎
-Moreover, for single edge perturbations (i.e., $\left|F\right|=1$), these upper and lower bounds become really tight in general, as expected:
-Corollary C.4.
-Let $G=(V,E)$ be an arbitrary simple graph with $E\subseteq\mathbb{E}_{c}(G)$ and $\left|V\right|=N\in\mathbb{N}$. Then,
-$\left|I(G,e_{i})\right|=\left|C(G)-C(G\setminus e_{i})\right|=\mathbf{O}(\log_{2}(N))\text{ ,}$
-if an edge is being deleted, or
-$\left|I^{-1}(G,e_{i})\right|=\left|C(G)-C(G\cup e_{i})\right|=\mathbf{O}(\log_{2}(N))\text{ ,}$
-if an edge is being inserted.
-Proof.
-First, note that in this particular case we have $r=|F|=1$. Therefore, the proof follows directly from Theorem C.1, where
-$2\left|F\right|\log_{2}\left(N\right)+\mathbf{O}\left(\log_{2}\left(\left|F\right|\right)\right)+\mathbf{O}(\log_{2}\left(\log_{2}(N)\right))\leq\\ 2\log_{2}\left(N\right)+\mathbf{O}\left(1\right)+\mathbf{O}(\log_{2}\left(\log_{2}(N)\right))\text{ .}$
-C.2 Eventual optimal values for graph complexity from edge deletions
-The semi-computability of the algorithmic complexity is a well-known phenomenon in algorithmic information theory and theoretical computer science [24, 26, 30, 43]: one can approximate the value of $C(x)$ from above with computable enumerations, but there is no computable function that return the value of $C(x)$ in general for every $x$.
-In other words, one equivalently cannot have an algorithm that in general can calculate how much the error between the approximate value to algorithmic complexity and its actual value is decreasing over time (as the enumerations of halting programs unfolds in the approximation from above).
-This contrasts with the result we will present in Theorem C.8, where we prove the existence of a single edge that carries maximum information about the whole network (i.e., it is maximally redundant with respect to the information content of the network) whose actual size of algorithmic information content (i.e., the exact algorithmic complexity value) can be realistically approximated.
-Furthermore, in this section we will explore the possibility of taking advantage of a single-edge perturbation on $G$ with purpose of calculating the value of $C(G)$.
-In other words, although no algorithm can determine in general the exact edge deletions that produce the desired phenomenon, we will present some particular cases in which an edge deletion works like an oracle that carries the sufficient amount of information for calculating the value of $C(G)$.
-Thus, algorithms for approximations of some arbitrary measures defined upon algorithmic information content, such as MILS algorithms, can eventually harvest this information in order to achieve these optimal values, should enough computational resources be expended.
-See Theorem C.9.
-Contributing to theoretical foundations of the methods employed in the present article, these eventually optimal calculations, although unpredictable in general, highlight the empirical significance of general methods based on algorithmic complexity, such as BDM, when compared with other lossless compression methods that do not (at least in a computable enumeration manner) asymptotically cover the space of partial recursive functions for compressing data.
-We will study a case in which the characteristic string (i.e., a bit string that univocally determines the presence or absence of the edges [11, 12]) of a compressible and dense enough graph already contains useful information for identifying where the edge deletion occurred, so that the deleted edge’s positioning becomes information itself.
-So, the aim is to prove a particular case and, thus, show that the phenomenon exists and may occur more often.
-To achieve Theorem C.9 with this purpose, the main idea of its proof is to construct some particular cases of simple graphs $G$ whose each characteristic string’s prefix—in the form $x\upharpoonright_{k}$ (i.e., a $k$-bit-length initial segment of $x$)—, respective to each $G$, carries information sufficient for computing the entire characteristic string $x$, so that this prefix is logarithmically much smaller than the characteristic string (i.e., $k=\mathbf{o}\left(\log_{2}\left(l(x)\right)\right)$), where $l\left(x\right)$ denotes the length of the string $x$.
-Moreover, these graphs should be sufficiently dense.
-Thus, there will be a program that, given the perturbed graph as input, can always decide where a perturbation landed onto the characteristic string, should the edge deletion occur in the characteristic string’s suffix.
-Hence, one just need to choose the actual index $i$ of the edge $e_{i}$ destroyed (i.e., the bit in the characteristic string in which $1$ was replaced with 00) that has
-$C\left(i\right)\leq\mathbf{o}\left(\log_{2}\left(\left|V(G)\right|\right)\right)$
-bits of algorithmic complexity. In addition, the information in the binary encoding of $i$ carries all the necessary information of the minimum generating program of $G$, so that
-$C\left(G^{*}|i\right)=\mathbf{O}\left(1\right)$
-and
-$C\left(i|\left(G\setminus e_{i}\right)\right)=\mathbf{O}\left(1\right)\text{ .}$
-In this manner, we will get that one becomes able to compute $G^{*}$ from $G\setminus e_{i}$, and vice versa.
-Before reaching the desired theorem, we introduce the concept of prefix self-determined graphs.
-Let $F_{G}$ be an infinite family of simple graphs.
-Let $p\in\{0,1\}^{*}$ only depend on the choice of $F_{G}$.
-By prefix self-determined, we mean that there is a program $p$ that can return the string $x$ given only a proper prefix of the characteristic string $x$ of $G$ as input.
-Formally, as in Definition C.1, we say a graph $G\in F_{G}$ is prefix self-determined iff there is $y=x\upharpoonright_{n}$ with $n<l(x)$ such that $\mathbf{U}\left(\left<y,p\right>\right)=x$, where $x$ is the characteristic string of $G$.
-In this article, we are applying the notion of prefix self-determination to graphs through its characteristic strings, but prefix self-determination is a property that can be clearly assigned to strings (or any other type of encodable object) in general.
-Definition C.1.
-Let $F_{G}$ be an infinite family of simple graphs.
-Let $p\in\{0,1\}^{*}$ only depend on the choice of $F_{G}$.
-We say a graph $G\in F_{G}$ is prefix self-determined iff there is $y=x\upharpoonright_{n}$ with $n<l(x)$ such that $\mathbf{U}\left(\left<y,p\right>\right)=x$, where $x$ is the characteristic string of $G$.
-Thus, prefix self-determination is a stronger notion in comparison to the widely employed self-delimitation [24, 26] from proper prefixes in algorithmic information theory.
-From the prefix of a prefix self-determined string, one might not only get the length of the string, but also promptly compute the very string.
-In other words, a prefix self-determined string contains a compression of itself.
-It is immediate to see that, for an infinite family of prefix self-determined strings, if $y$ is the respective proper prefix of string $s$, then
-$C(s)\leq C(y)+\mathbf{O}(1)\leq l(y)+\mathbf{O}(1)\leq l(s)+\mathbf{O}(1)\text{ .}$ (53)
-As a natural extension, one may want to limit a maximum length for the prefix with respect to the string.
-This is the case of prefix self-determined graphs whose initial segment $y=x\upharpoonright_{n}$ of its characteristic string has length $n=\mathbf{O}\left(\log_{2}\left(l(x)\right)\right)$.
-Let $F_{G}$ be an infinite family of simple graphs.
-As in Definition C.2, we say a graph $G\in F_{G}$ is logarithmically prefix self-determined iff $G$ is prefix self-determined and $n=\mathbf{O}\left(\log_{2}\left(l(x)\right)\right)$.
-Thus, a logarithmically prefix self-determined graph has a very small subset of edges (which are encoded into the initial segment of the characteristic string) that algorithmically determine the whole graph.
-This property grasps the idea of a small subgraph carrying all the information sufficient for computably determining the graph, should the initial segment of the characteristic string be in turn a characteristic string of a vertex-induced subgraph.
-Definition C.2.
-Let $F_{G}$ be an infinite family of simple graphs.
-We say a graph $G\in F_{G}$ is logarithmically prefix self-determined iff $G$ is prefix self-determined and $n=\mathbf{O}\left(\log_{2}\left(l(x)\right)\right)$.
-Remember the usual notion of a family of dense graphs being defined as an infinite family $F_{G}^{{}^{\prime}}$ of graphs in which every member $G\in F_{G}^{{}^{\prime}}$ has $\left|E(G)\right|=\mathbf{\Theta}\left(\left|V(G)\right|^{2}\right)$.
-In order to prove that Definition C.2 and, as a direct consequence, Definition C.1 are not vacuous (i.e, that they are satisfiable), we can choose a particular type of self-delimiting program that returns itself concatenated with repetitions of itself (in addition, we choose a prefix that is also nearly incompressible):
-Lemma C.5.
-There is an infinite family $F_{G}$ of logarithmically prefix self-determined dense graphs whose respective proper prefixes $v\in\{0,1\}^{*}$ follows
-$\left|C(v)-l(v)\right|=\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)$
-and
-$l(v)=\mathbf{o}\left(\log_{2}\left(l(x)\right)\right)\text{ ,}$
-where $x$ is the characteristic string of $G$.
-Proof.
-This proof is based on constructing characteristic strings of the form $w\;r\upharpoonright_{k}w^{\prime}$, where $r\in\left[0,1\right]\subseteq\mathbb{R}$ is an arbitrary $\mathbf{O}(1)$-K-random (i.e., prefix algorithmically random as in Definition B.9) real number111111 For example, one may choose $r$ to be the halting probability $\Omega$ without loss of generality in this proof., and $w,r\upharpoonright_{k},w^{\prime}\in\{0,1\}^{*}$.
-Moreover, we want that
-$l(w)+k=\mathbf{o}\left(\log_{2}\left(l(w\;r\upharpoonright_{k}w^{\prime})\right)\right)=\mathbf{O}\left(\log_{2}\left(l(w\;r\upharpoonright_{k}w^{\prime})\right)\right)\text{ ,}$
-$\left|E(G)\right|=\mathbf{\Theta}\left(\left|V(G)\right|^{2}\right)\text{ ,}$
-and that there is a program $p_{2}$ such that
-$\mathbf{U}\left(\left<w\,r\upharpoonright_{k},p_{2}\right>\right)=w\;r\upharpoonright_{k}w^{\prime}\text{ .}$
-To this end, first let $p_{2}\in\{0,1\}^{*}$ be a string that represents the algorithm running on a prefix universal Turing machine $\mathbf{U}$ that receives $x\in\{0,1\}^{*}$ as input and:
-  1. (I)
-    first, decides whether or not there are $w\in\mathbf{L^{\prime}_{U}}$ and $y\in\{0,1\}^{*}$ such that $x=wy$, where $l(w)<l(x)$ ;
-  2. (II)
-    then,
-    1. (a)
-      if $x=wy$, and this holds where $l(w)$ is the minimum length of a string $w$ for which the above Step (I) holds:
-      1. i.
-        build the string
-        $z=y\,y\,y\cdots\,\left(y\upharpoonright_{h}\right)$
-        by concatenating the string $y$ until finding $h\in\mathbb{N}$, where $h\leq l(y)$, for which
-        $l(wz)=\frac{n^{2}-n}{2}\text{ ,}$
-        where
-        $n=\min\left\{t\middle|\begin{array}[]{l}t\in\mathbb{N}\\ \frac{t^{2}-t}{2}\geq 2^{{l(y)}^{2}}+l(w)\end{array}\right\}\text{ ;}$
-      2. ii.
-        return $wz$.
-    2. (b)
-      Otherwise, return “$error$”.
-Note that $p_{2}$ always halts for every input and we immediately have that either $\mathbf{U}\left(\left<x,p_{2}\right>\right)=wz$ or $\mathbf{U}\left(\left<x,p_{2}\right>\right)=\text{ ``$error$'' }$ ”.
-Since $w$ and $y$ were arbitrary in the construction of $p_{2}$, let $w$ encode the length of $y$ in language $\mathbf{L^{\prime}_{U}}$.
-Hence, we will have that
-$l(wy)$ $=l(w)+l(y)\leq$ (54)
-$\leq\mathbf{O}\left(\log_{2}\left(l(y)\right)\right)+\,l(y)\leq$
-$\leq\mathbf{o}\left(l(y)\right)+\,l(y)=$
-$=\mathbf{o}\left({l(y)}^{2}\right)$
-$=\mathbf{o}\left(\log_{2}\left(l(w)+2^{{l(y)}^{2}}\right)\right)=$
-$=\mathbf{o}\left(\log_{2}\left(l(wz)\right)\right)\text{ .}$
-From [9, 11, 12], we know the family of all possible simple graphs can be recursively labeled with a unique way of ordering the sequence of all possible edges.
-Thus, since by construction there is $n\in\mathbb{N}$ such that $l(wz)=\frac{n^{2}-n}{2}$, let $G$ be a simple graph whose characteristic string is $wz=\mathbf{U}\left(\left<wy,p_{2}\right>\right)$.
-This way, we have that
-$\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}=l(wz)\geq 2^{{l(y)}^{2}}+l(w)$ (55)
-and, by the construction of $p_{2}$ and the minimality of $|V(G)|$,
-$\mathbf{O}\left(2^{{l(y)}^{2}}+l(w)\right)+2^{{l(y)}^{2}}+l(w)\geq\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}=l(wz)$ (56)
-Until here, we proved the logarithmically prefix self-determination part and proved that
-$l(wy)=\mathbf{o}\left(\log_{2}\left(l(wz)\right)\right)\text{ .}$
-Now, let $r\in\left[0,1\right]\subseteq\mathbb{R}$ be an arbitrary $\mathbf{O}(1)$-K-random real number.
-From the Borel normality of $r$ (as in the proof of Corollary C.2), we know there is a large enough $k\in\mathbb{N}$ such that the number #$\#_{1}(r\upharpoonright_{k})$ of occurrences of $1$’s in $r\upharpoonright_{k}$ follows
-|#$\left|\#_{1}(r\upharpoonright_{k})-\frac{k}{2}\right|=\mathbf{O}\left(\sqrt{k\log_{2}\left(k\right)}\right)\text{ .}$ (57)
-Since $w$ and $y$ were arbitrary in the construction of $p_{2}$, let $y\coloneqq r\upharpoonright_{k}$.121212 Note that, for any large enough length $l(wz)$, the string $wz$ cannot ever be $\mathbf{O}(1)$-K-random, since it is based on multiple repetitions of the same string.
-Therefore, since $k>k_{0}$ can be arbitrarily large given a large enough $k_{0}$, we will obtain an infinite family $F_{G}$ of distinct graphs with characteristic strings of the form $wz$.
-Remember that, by construction, $w$ encodes the integer value $k$ and that $k>k_{0}$ can be arbitrarily large given that large enough $k_{0}$.
-Therefore, from Equations (55), (56), and (57), by construction of $p_{2}$ and since $l(w)=\mathbf{O}\left(\log_{2}\left(k\right)\right)$, we will have that
-$\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}&\geq\left|E(G)\right|\\ &\geq\frac{2^{k^{2}}-k}{2}-\left(\frac{2^{k^{2}}-k}{k}\right)\,\mathbf{O}\left(\sqrt{k\log_{2}\left(k\right)}\right)\\ &=\mathbf{\Omega}\left(2^{k^{2}}+l(w)\right)\\ &=\mathbf{\Omega}\left(\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}\right)\text{ .}\end{aligned}$ (58)
-Notice that Equation (58) holds because
-$\frac{2^{k^{2}}-k}{2}-\left(\frac{2^{k^{2}}-k}{k}\right)\,\mathbf{O}\left(\sqrt{k\log_{2}\left(k\right)}\right)\gg\frac{2^{k^{2}}-k}{2}-\frac{2^{k^{2}}-k}{4}$ (59)
-and
-$\frac{2^{k^{2}}-k}{2}>\frac{2^{k^{2}}+l(w)}{4}$ (60)
-hold for sufficiently large $k$.
-Therefore, Equation (58) proves $\left|E(G)\right|=\mathbf{\Theta}\left(\left|V(G)\right|^{2}\right)$.
-Now, the fact that $C(r\upharpoonright_{k})=C(y)\geq k-\mathbf{O}(1)$ holds follows directly from Definition B.9.
-In addition, from Lemma B.1, we have that
-$C\left(wy\right)\leq l\left(r\upharpoonright_{k}\right)+\mathbf{O}\left(\log_{2}\left(k\right)\right)+\mathbf{O}\left(\log_{2}\left(\log_{2}\left(k\right)\right)\right)\\ \leq k+\mathbf{O}\left(\log_{2}\left(k\right)\right)$ (61)
-and
-$C(y)\leq C(wy)+\mathbf{O}\left(\log_{2}\left(\log_{2}\left(k\right)\right)\right)$ (62)
-Therefore, by making $v\coloneqq wy$, where $l(y)=k$ and $l(w)=\mathbf{O}\left(\log_{2}\left(k\right)\right)$, and achieving
-$k-\mathbf{O}\left(\log_{2}\left(\log_{2}\left(k\right)\right)\right)\leq C(v)\leq k+\mathbf{O}\left(\log_{2}\left(k\right)\right)$ (63)
-from the incompressibility of $r$ and Equations (61) and (62), we will have it that
-$l(v)-\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)-\mathbf{O}\left(\log_{2}\left(\log_{2}\left(l(v)\right)\right)\right)\leq C(v)\leq\\ l(v)+\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)$ (64)
-Now we show in Lemma C.6 that the compression of the characteristic string and its density are so strong that enables one to pick and delete the right edge in a way such that this very deletion encodes the optimal compression of $G$:
-Lemma C.6.
-There is an infinite family $F_{G}$ of logarithmically prefix self-determined dense simple graphs (in particular, any family satisfying Lemma C.5) such that, for every large enough member $G\in F_{G}$, there is at least one edge deletion so that
-$C\left(G^{*}|\left(G\setminus e_{i}\right)\right)=\mathbf{O}(1)\text{ .}$
-Proof.
-Let $F_{G}$ be an arbitrary infinite family of graphs satisfying Lemma C.5.
-Thus, there is $v=x\upharpoonright_{n}$ with $n<l(x)$ such that $\mathbf{U}\left(\left<v,p_{2}\right>\right)=x$, $n=\mathbf{o}\left(\log_{2}\left(l(x)\right)\right)$, and $\left|C(v)-l(v)\right|=\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)$, where $x$ is the characteristic string of $G$.
-Let $p_{3}\in\{0,1\}^{*}$ be a string that represents the algorithm running on a prefix universal Turing machine $\mathbf{U}$ that receives the adjacency matrix $A_{G^{\prime}}$ and $p_{2}$ as inputs and:
-  1. 1.
-    applies any previous arbitrarily chosen method of recursively encoding a matrix into a string and check whether or not $A_{G^{\prime}}$ is a matrix;
-  2. 2.
-    if $A_{G^{\prime}}$ is a matrix, then:
-    1. (a)
-      reversively applies the previous arbitrarily chosen method of recursively encoding a matrix into a string to produce the correspondent characteristic string $x^{\prime}$;
-    2. (b)
-      decides whether or not there are $w^{\prime}\in\mathbf{L^{\prime}_{U}}$ and $z^{\prime}\in\{0,1\}^{*}$ such that $x^{\prime}=w^{\prime}\,z^{\prime}$, where $l(w^{\prime})<l(x^{\prime})$ and $w^{\prime}$ encodes the binary representation of a natural number $h\in\mathbb{N}$ in language $\mathbf{L^{\prime}_{U}}$;
-    3. (c)
-      if $x^{\prime}=w^{\prime}\,z^{\prime}$ holds such that $l(w^{\prime})$ is the minimum prefix for which this holds, then:
-      1. i.
-        makes $v\coloneqq w^{\prime}\,y^{\prime}$, where $y^{\prime}$ is the string of the next $\mathbf{U}\left(w^{\prime}\right)$ bits in $x^{\prime}$ right after the end of $w^{\prime}$;
-      2. ii.
-        makes $x\coloneqq\mathbf{U}\left(\left<v,p_{2}\right>\right)$  ;
-      3. iii.
-        decides whether or not $x^{\prime}=x$ ;
-      4. iv.
-        if $x^{\prime}<x$ and $l(x)=l(x^{\prime})$, then:
-        1. A.
-          enumerates all the $1$’s in $x$;
-        2. B.
-          search for the first bit of $x$ that displays value $1$ while the same bit of $x^{\prime}$ displays value 00;131313 Notice that this step evinces one of the advantages of using the characteristic string formalism from [9, 11, 12].
-        3. C.
-          calculates de index $i\in\mathbb{N}$, where the $i$-th $1$ in $x$ corresponds to this first bit in which the values differ from each other;
-        4. D.
-          returns the binary representation $\left(i\right)_{2}\in\{0,1\}^{*}$ of the integer $i\in\mathbb{N}$.
-      5. v.
-        otherwise, returns “$error$”.
-    4. (d)
-      otherwise, returns “$error$”.
-  3. 3.
-    otherwise, returns “$error$”.
-Note that, if $\mathbf{U}$ halts on $v,p_{2}\right>$, then $p_{3}$ always halts. We will have it that by hypothesis, the minimality of algorithmic complexity, and Lemmas B.1 and B.2,
-$C(x)$ $\leq C(v)+\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)+\mathbf{O}(1)\leq$ (65)
-$\leq l(v)+\mathbf{O}\left(\log_{2}\left(l(v)\right)\right)+\mathbf{O}(1)=$
-$=\mathbf{o}\left(\log_{2}\left(l(x)\right)\right)$
-holds.
-This occurs because $\mathbf{U}\left(\left<v,p_{2}\right>\right)=x$ and $l(v)=\mathbf{o}\left(\log_{2}\left(l(x)\right)\right)$ hold.
-Therefore, once a method of recursively encoding a matrix into a string to produce the correspondent string $x^{\prime}$ is fixed, we will have that, from Lemmas B.1 and B.2 and the fact that $x$ is the characteristic string of $G$,
-$l(G^{*})=C(G)\leq\mathbf{o}\left(\log_{2}\left(\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}\right)\right)+\mathbf{O}(1)$ (66)
-holds. Therefore, since there are only up to
-$2^{\mathbf{o}\left(\log_{2}\left(\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}\right)\right)+\mathbf{O}(1)}$ (67)
-possible bit strings of length $\leq l(G^{*})$ and $G$ is dense, then there is at least one present edge that is the $i$-th bit of $x$, where $i>l(v)$, and $i\in\mathbb{N}$ corresponds to the program $G^{*}\in\mathbf{L^{\prime}_{U}}$.
-Then, $\mathbf{U}\left(\left<A_{(G\setminus e_{i})},p_{2},p_{3}\right>\right)=G^{*}$ and, since $p_{2}$ and $p_{3}$ are fixed, $C\left(G^{*}|(G\setminus e_{i})\right)=\mathbf{O}(1)$.
-∎
-Notice that the procedure in the proof of Lemma C.6 was designed in such a way that it can be recursively inverted so as to make the perturbed graph be computably retrievable from $G^{*}$ too:
-Lemma C.7.
-There is an infinite family $F_{G}$ of logarithmically prefix self-determined dense simple graphs (in particular, any family satisfying Lemma C.5) such that, for every large enough member $G\in F_{G}$, there is at least one edge $e_{i}$ that satisfies Lemma C.6 such that
-$C\left(\left(G\setminus e_{i}\right)|G^{*}\right)=\mathbf{O}(1)$
-also holds.
-Proof.
-The idea of the proof is to construct a program that, with entry $G^{*}$ as input, performs the inverted operation that program $p_{3}$ from the proof of Lemma C.6 does.
-Then, it builds the correspondent adjacency matrix of the perturbed graph, that is, the adjacency matrix of $A_{(G\setminus e_{i})}$ that satisfies Lemma C.6.
-To this end, let $p^{\prime}_{3}\in\{0,1\}^{*}$ be a string that represents the algorithm running on a prefix universal Turing machine $\mathbf{U}$ that receives the $\left(i\right)_{2}$ as input and:
-  1. 1.
-    makes $A_{G}\coloneqq\mathbf{U}\left(\left(i\right)_{2}\right)$  ;
-  2. 2.
-    builds the characteristic string $x$ of $A_{G}$;
-  3. 3.
-    replace only the $i$-th $1$ in the characteristic string $x$ with 00 and builds the resulting string $x^{\prime}$, where $x^{\prime}<x$;
-  4. 4.
-    returns $A_{G^{\prime}}$ for which $x^{\prime}$ is its characteristic string.
-If $i$ corresponds to the program $G^{*}$ in the proof of Lemma C.6, then we will have that $p^{\prime}_{3}$ halts and $A_{G^{\prime}}=A_{(G\setminus e_{i})}$.
-Therefore, we will have that, from the minimality of algorithmic complexity,
-$C\left(\left(G\setminus e_{i}\right)|G^{*}\right)=\mathbf{O}(1)$ (68)
-holds. ∎
-Now remember that a neutral information edge $e$ is one such that the information difference is minimal so that it contributes less to the whole information content of $G$.
-By combining the previous two Lemmas C.6 and C.7 one can easily obtain the following result:
-Theorem C.8 (Maximally neutral information edge).
-There is an infinite family $F_{G}$ of logarithmically prefix self-determined dense simple graphs (in particular, any family satisfying Lemma C.5) such that, for every large enough member $G\in F_{G}$, there is at least one (maximally) neutral information edge $e_{i}\in E(G)$ such that
-$\left|I(G,e_{i})\right|=\left|C(G)-C(G\backslash e_{i})\right|=\mathbf{O}(1)$ (69)
-holds.
-Proof.
-The proof follows directly from Lemmas C.6 and C.7 and Definition B.12 by employing the basic properties of AIT in Lemmas B.1 and B.2. ∎
-Identifying complex systems that displays such a phenomenon, as the graphs in Theorem C.8, may reveal those systems that are capable of carrying the sufficient information about where the perturbation occur and, hence, it enables future investigation on how this correlates with robustness or self-repair, two concepts which may be also related to the preservation of network topological properties studied in the present article.
-Moreover, one can take advantage of Lemma C.5 in order to obtain one of the particular cases in which a computer can already find nearly optimal values of algorithmic complexity in polynomial computation time (i.e., running time) even with brute force, as shown in Theorem C.9.
-The proof of Theorem C.9 is based on the fact that, although program $p_{2}$ demands an exponential computation time with input $wy$, the computation time of emulating program $wy,p_{2}\right>$ is still polynomial with respect to $\left|E(G)\right|$, since $G$ is dense from Lemma C.5.
-This way, from the “almost” incompressibility of the prefix $wy$ also demonstrated in Lemma C.5, we will achieve the desired upper bound for the error with respect to the optimal value $C(G)$.
-Theorem C.9.
-Let $G$ be a large enough simple graph that belongs to an infinite family $F_{G}$ of logarithmically prefix self-determined dense simple graphs.
-Then, there is a polynomially time-bounded exhaustive search algorithm that, with input $G\in F_{G}$, can approximate the integer value $C(G)$ with an error bounded by $\mathbf{o}\left(\log_{2}\left(\log_{2}\left(\left|V(G)\right|\right)\right)\right)$ bits.
-Proof.
-First, we choose without loss of generality a self-delimiting universal programming language $\mathbf{L^{\prime}_{U}}$ that is recognizable in polynomial time as in Section B.6.
-Let $\mathbf{P}$ be the usual time complexity class of polynomially time-bounded deterministic Turing machines [18, 52].
-As a direct consequence, then we will have that Steps (I) or (II)b in the proof of Lemma C.5 can be solved in $\mathbf{P}$.
-Note that Step (II)(a)ii is basically the concatenation of the string $w$ obtained from Step (I) with the string resulting from Step (II)(a)i.
-Thus, since we know the concatenation operation [50] and Step (I) are in $\mathbf{P}$, we can just focus on analyzing the time complexity of Step (II)(a)i, which is indeed where the computation is most demanding.
-In addition, from the minimality of $n$, construction of $p_{2}$, and Equation (56), we have that at most
-$\frac{\mathbf{O}\left(2^{{l(y)}^{2}}+l(w)\right)}{l(y)}$
-concatenations will be needed to find the minimum $n$ with
-$l(wz)=\frac{n^{2}-n}{2}\geq 2^{{l(y)}^{2}}+l(w)\text{ .}$
-Now, given $k\geq 4$ tapes with one of the tapes filled with the binary representation of this maximum number of concatenations, one of the tapes filled with $w$, and another tape filled with $y$ (which leaves at least one of the remaining tapes empty or blank), we will have that there is a $k$-tape Turing machine that returns $wz$ in at most
-$\mathbf{O}\left(\left(\frac{2^{{l(y)}^{2}}+l(w)}{l(y)}\right)\,\left(\max\left\{l(y),l(w)\right\}\right)\right)$ (70)
-computation steps.
-This holds because the tape in which one informs the number of concatenations (needed to calculate $n$) can be filled exactly with $\frac{\mathbf{O}\left(2^{{l(y)}^{2}}+l(w)\right)}{l(y)}$ as many contiguously repeated $1$’s; and this tape works as a counter to the number of concatenations the $k$-tape Turing machine must perform in the remaining empty tape in order to halt when this tape is filled with $wz$.
-Now, note that the worst-case time complexity for another machine to construct the counter string (i.e., the one constituted by $1$’s with length $\frac{\mathbf{O}\left(2^{{l(y)}^{2}}+l(w)\right)}{l(y)}$) from input $wy$ is upper bounded by an exponential term $2^{\mathbf{O}\left({l(y)}^{2}\right)}$.
-Therefore, since the worst-case time complexity of simulating any $k$-tape Turing machine is also in $\mathbf{P}$ [18, 52], then the time complexity of $p_{2}$ is upper bounded by an exponential term as a function of $l(wy)$ (i.e., $p_{2}$ represents an algorithm in $\mathbf{EXPTIME}$).
-In fact, since $wz$ is the output of $p_{2}$, we will necessarily also have a best-case time complexity (for one-way single-tape Turing machines) of
-$2^{{l(y)}^{2}}+l(w)-l(wy)$
-by just comparing the lengths of the output and input.
-Thus, since $l(wy)=\mathbf{o}\left(\log_{2}\left(l(wz)\right)\right)$, $p_{2}$ is also exponential in the best case scenario, which thereby implies that the worst-case time complexity of $p_{2}$ is indeed exponential.
-Remember by hypothesis that $G$ is dense, i.e., $\left|E(G)\right|=\mathbf{\Theta}\left(\left|V(G)\right|^{2}\right)$.
-Then, we will have that $\left|E(G)\right|=\mathbf{\Theta}\left(l(wz)\right)$.
-Now, note that there are only up to $2^{l(\left<wy,p_{2}\right>)}$ possible bit strings with length at most $l(\left<wy,p_{2}\right>)$.
-Since the worst-case time complexity of simulating any Turing machine in a universal Turing machine is polynomially bounded with respect to the original computation time of the machine being simulated [18, 52], then we have that exhaustively searching all possible outputs of exponentially time-bounded programs with length $\leq l(\left<wy,p_{2}\right>)$ has a worst-case computation time bounded by
-[…]
-for some large enough constant $C\in\mathbb{N}$. Remember our choice of the pairing function in Notation B.5, so that, since $p_{2}$ is fixed, we will have that
-$l\left(\left<wy,p_{2}\right>\right)\leq l(wy)+l(p_{2})+\mathbf{O}\left(\log_{2}\left(l(wy)+l(p_{2})\right)\right)=\\ \mathbf{O}\left(l(wy)\right)\text{ .}$
-Therefore, since $l(wy)=\mathbf{o}\left(\log_{2}\left(wz\right)\right)$ by hypothesis, Equation (71) holds, and as we have shown, $p_{2}$ represents an algorithm in $\mathbf{EXPTIME}$ given $wy$ as input, we will have that there is an exhaustive search algorithm with $G$ as input that will eventually test $wy,p_{2}\right>$ as the generating program of $G$ in worst-case computation time upper bounded by
-$\mathbf{o}\left(\left(l\left(wz\right)\right)^{c^{\prime}}\right)\leq\mathbf{O}\left(\left|E(G)\right|^{c^{\prime\prime}}\right)\text{ ,}$ (72)
-for some large enough constants $c^{\prime},c^{\prime\prime}\in\mathbb{N}$.
-Therefore, such an algorithm returns a value $C^{\prime}(G)\leq l\left(\left<wy,p_{2}\right>\right)$ in polynomial computation time with input $G$.
-Notice that, from Lemma C.5, we have that $\left|C(wy)-l(wy)\right|=\mathbf{O}\left(\log_{2}\left(l(wy)\right)\right)$, $G$ is dense, and $l(wy)=\mathbf{o}\left(\log_{2}\left(l(wz)\right)\right)$ hold, and $w$ encodes the length of $y$.
-Then, from Lemma C.5 and the basic properties of algorithmic information theory in Section  B.6, we will have it that
-$C(G)$ $\leq C^{\prime}(G)+\mathbf{O}(1)$ (73)
-$\leq l(wy)+\mathbf{O}(1)$
-$\leq l(y)+\mathbf{O}\left(\log_{2}\left(l(y)\right)\right)$
-$\leq l(y)+\mathbf{o}\left(\log_{2}\left(\log_{2}\left(l(wz)\right)\right)\right)\text{ .}$
-We also obtain that
-$C^{\prime}(G)+\mathbf{O}(1)$ $\geq C(y)$ (74)
-$\geq l(y)-\mathbf{O}(1)$
-$\geq C(y)-\mathbf{O}\left(\log_{2}\left(l(y)\right)\right)-\mathbf{O}(1)$
-$\geq C(y)-\mathbf{o}\left(\log_{2}\left(\log_{2}\left(l(wz)\right)\right)\right)-\mathbf{O}(1)\text{ .}$
-Finally, from Equations (73) and (74) and the fact that $l(wz)=\frac{\left|V(G)\right|^{2}-\left|V(G)\right|}{2}$, we will have it that
-$\left|C^{\prime}(G)-C(G)\right|=\mathbf{o}\left(\log_{2}\left(\log_{2}\left(\left|V(G)\right|\right)\right)\right)$
-holds, which finishes the proof of Theorem C.9. ∎
-Thus, Theorem C.9 shows that there is an algorithm that, with $G$ as input, calculates a value $C^{\prime}(G)$ that could be considered as a good approximation of the algorithmic complexity of $G$ in terms of compression rate.
-Although the actual $C(G)$ and the approximation $C^{\prime}(G)$ may differ by a number of bits of about a logarithmic term of its own value $C(G)$ (or of the order of a double logarithmic term of the input size), they are both asymptotically comparable with respect to $\left|V(G)\right|$.
-This phenomenon is expected to have an even faster speedup for algorithms that approximate algorithmic complexity values, such as BDM that is based on the decomposition into blocks of objects whose complexity is approximated using the Coding Theorem Method (CTM) [32, 56, 55, 62], which in turn precomputes lookup tables so as to exchange computation time for space/memory [68].
-In this sense, BDM can reach optimality for even larger families of graphs beyond that of Theorem C.9.
-This is because one can easily extend the latter in order to demonstrate that BDM not only can find the nearly optimal approximations for logarithmically prefix self-determined graphs, but also can find those for graphs that are self-determined by logarithmically size-bounded $2$D- or $3$D-partitions of most (if not all) of its constituent subgraphs.
-Additionally, the precomputation of the lookup table enables BDM to take more advantage of larger quantities of constituent blocks in comparison to other brute-force methods of approximating algorithmic complexity from scratch every time.
-We suggest that this theoretical finding might be one of the reasons MILS is performing better than other size reduction methods.
-Future research should be pursued with the purpose of revealing other intrinsic structural characteristics underlying MILS and other data size reduction techniques, particularly with respect to modularity [72, 38, 67].
-The following corollary follows directly from the combination of Theorems C.8 and C.9.
-Corollary C.10.
-Let $G$ be a large enough simple graph that belongs to an infinite family $F_{G}$ of logarithmically prefix self-determined dense simple graphs.
-Then, there is a polynomially time-bounded exhaustive search algorithm such that for every large enough member $G\in F_{G}$, there is at least one edge $e_{i}\in E(G)$ such that this algorithm can approximate both the integer values $C(G)$ and $C(G\setminus e_{i})$ with an error bounded by $\mathbf{o}\left(\log_{2}\left(\log_{2}\left(\left|V(G)\right|\right)\right)\right)$ bits.
-Notice that Theorem C.9 and Corollary C.10 contribute to the theoretical foundations of the methods employed in the main article by showing that these eventually optimal calculations, although unpredictable in general, highlight the empirical significance of general methods based on algorithmic complexity, such as BDM, when compared with other lossless compression methods that do not (at least in a computable enumeration manner) asymptotically cover the space of partial recursive functions for compressing data.
-C.3 Time Complexity of Minimal Information Loss Selection (MILS)
-For MILS to be well-defined, we need to guarantee that the deletion of an element (e.g., a vertex or edge in a graph) produces a unique object in a deterministic fashion.
-In the case of edges in networks, the problem is when there are elements $e_{i},e_{j}\in E$ such that $I(G,e_{i})=I(G,e_{j})$ and the algorithm cannot uniquely decide whether to remove $e_{i}$ or $e_{j}$ first, which potentially (and likely) leads to the production of different objects.
-Algorithm 2 avoids this problem and shows that the procedure is robust.
-Note that this algorithm and Theorem C.11 are stated for graphs, but they can be generalized for multidimensional objects $\mathbb{U}$ in general.
-Theorem C.11.
-Let $T_{ID}(G)$ be the worst-case computation time of calculating the value $I(G,e_{i})$ for any $e_{i}\in E\left(G^{\prime}\right)$ by the chosen algorithmic-complexity approximating algorithm, where $G^{\prime}$ is any edge-induced subgraph of $G$141414 For example, by choosing the BDM, $T_{ID}(G)$ is the worst-case computation time of calculating the value of $K_{log}(G^{\prime})-K_{log}(G^{\prime}\setminus e_{i})$ for every $e_{i}\in E\left(G^{\prime}\right)$, where $G^{\prime}$ is any edge-induced subgraph of $G$.
-See Definition B.12 and Equation (16).
-Then, MILS (Algorithm 2) is a deterministic algorithm that has a cubic worst-case time complexity with input $E(G)$ and linear worst-case time complexity in $T_{ID}(G)$.
-Proof.
-Let $e_{i}$ and $e_{j}$ be any two edges of a graph $G$ (whose size is to be reduced by the application of MILS) with $I(G,e_{i})=I(G,e_{j})$, then MILS will remove $e_{i}$ and $e_{j}$ from $G$ simultaneously.
-This element deletion condition helps MILS to always produce the same output for a given input, which makes it a deterministic algorithm.
-To prove this claim, note that $I(G,e_{i})$ calculated by any algorithmic-complexity approximating algorithm (e.g., by the BDM) always produce an integer, ordering any set of integers is decidable and always produce unique outputs for the same set of integers.
-Therefore, for every iteration of the loop in [Algorithm 2, Step 13], there will be a unique size-reduced graph $G_{n-i-1}$ from a previous graph $G_{n-i}$ by deleting the edges in $\{e_{i}\in E(G_{n-i}):I(G_{n-i},e_{i})=\textit{minLoss}\}$.
-In fact, MILS can only vary the output for the same input if more computation time is expended on the CTM’s precomputation before the BDM.
-In order to study the time complexity of MILS, note that the worst-case of function $\text{MILS}(G,N)$ is given when $N=1$ and InfoRank(G) outputs a list of $\left|E(G)\right|$ distinct entries.
-This way, since one has $\left|E(G)\right|$ iterations in the loop in [Algorithm 2, Step 4] by insertion (thus, an operation in $\mathbf{O}(1)$) of the result of the approximation of $I(G,e_{i})$ in the array informationLoss, which has size $\left|E(G)\right|$, and sorting algorithms have quadratic worst-case time complexity, we will have a worst-case time complexity of function InfoRank(G) in
-$\mathbf{O}\left(\left|E(G)\right|\left(T_{ID}(G)\right)\right)+\mathbf{O}\left(\left|E(G)\right|^{2}\right)\text{ .}$ (75)
-In addition, since one has up to $\left|E(G)\right|$ iterations in the loop, finding minimums in the first dimension of arrays takes a linear worst-case computation time, and deleting elements in an array also takes a linear worst-case computation time, we will have a worst-case time complexity of function $\text{MILS}(G,N)$ in
-$\mathbf{O}\left(\left|E(G)\right|\,\left(T_{IR}(G)+\left|E(G)\right|\right)\right)\text{ ,}$ (76)
-where $T_{IR}(G)$ is the worst-case computation time of function InfoRank(G) for any $G$ (which was already studied in Equation (75) so that
-$T_{IR}(G)=\mathbf{O}\left(\left|E(G)\right|\left(T_{ID}(G)\right)+\left|E(G)\right|^{2}\right)$ (77)
-holds).
-Now, suppose without loss of generality that the inputs of Algorithm 2 are $E(G)$ and $N$.
-Therefore, since composing functions is closed under polynomial computation time as in [Algorithm 2, Step 15] and function InfoRank(G) is called recursively, we will have that the worst-case time complexity of Algorithm 2 is
-$\mathbf{O}\left(\left|E(G)\right|\,\left(\left|E(G)\right|\left(T_{ID}(G)\right)+\left|E(G)\right|^{2}+\left|E(G)\right|\right)\right)=$ (78)
-$\mathbf{O}\left(\left|E(G)\right|^{3}+\left|E(G)\right|^{2}\left(T_{ID}(G)\right)\right)$
-Algorithm 2 also produces a speedup with respect to Algorithm 3 when all elements have the same information value, and thus are to be deleted at the same time: while in Algorithm 3 there would be a time complexity overhead to perform this task, in Algorithm 2, given a previously calculated InfoRank(G), it will be required just one step in the loop and, therefore, just a linear computation time in $\mathbf{O}\left(\left|\textit{informationLoss}\right|\right)$, where informationLoss is the output of InfoRank(G).
-For example, any attempt to reduce the dimension of the complete graph (either by e.g.
-single-vertex or single-edge deletion) will produce an empty graph.
-Thus, MILS’ Algorithm 2 scales in a tractable manner with respect to the complexity approximating method for very large networks.
-This is because the worst-case running time is basically dominated by the amount of computational resources expended in the sole approximation to the values of algorithmic complexity as this amount only contributes linearly to the overall time complexity when the network size increases.
+This is because information in Fig.
+12(A) is distributed more uniformly than Fig.
+12(C) but, according to the algorithm, some features in certain regions are more important than others and thus some distortions are allowed if such features have high algorithmic content according to the underlying computer programs that reconstruct the compressed images.
+However, in Fig.
+12(C), blank space between formulae can be sacrificed first as they do not contain any information.
+While this behaviour for handling different cases can be replicated in ad-hoc algorithms, MILS handles it naturally and determines when distorting can be allowed or proportion can be preserved, something that image compression algorithms would not do as their goal is to recover the same image dimensions and maximise storage compression and not to preserve algorithmic information content even sacrificing otherwise fundamental properties of the image for those algorithms such as image dimension.
+6 Conclusion
+We have introduced a set of dimensionality reduction techniques based on a Minimal Information Loss Selection (MILS) algorithm, a recursive unsupervised Machine Learning approach particularly effective in reducing data dimensionality preserving essential features in data.
+In the context of networks, we have shown it preserves features such as degree distribution, clustering coefficient, edge betweenness, and centralities better than traditional and state-of-the-art algorithms.
+As studied in Section 5.2, MILS demonstrates comparable or superior performance to existing network reduction algorithms on different classes and types of networks.
+Moreover, it extends to lossy compression of images and bi-dimensional data as shown in Section 5.3.
+It operates on the principle of local perturbation analysis and Algorithmic Information Dynamics, offering robustness against varying methods of approximating algorithmic complexity.
+The method presented in this article optimally preserves the algorithmic information content (which might be affected by the local perturbations) whenever an optimal method is used to approximate algorithmic complexity, effectively introducing a lossy technique based on principles of lossless recursive compression.
+Its reliance on first principles guarantees that its performance can only increase as more computational resources are made available, which differs from other methods based on traditional statistical Machine Learning for which theoretical limitations are known to hold [11].
+Nevertheless, the empirical application of MILS only scales linearly with the worst-case time complexity of the chosen method for approximating algorithmic complexity.
+Funding
+FSA acknowledges support from the São Paulo Research Foundation (FAPESP), grants $2023$/$05593$-$1$ and $2021$/$14501$-$8$. NAK is supported by a Horizon $2020$ EU grant (MultipleMS, $733161$).